### v1

# Sequential Randomization Tests Using E-values: A Betting Approach for Clinical Trials

Abstract
Sequential monitoring of randomized trials traditionally relies on parametric assumptions or asymptotic approximations. We present a nonparametric sequential test – the randomization e-process (e-RT) – that derives validity solely from the randomization mechanism. Using a betting framework, e-RT constructs a test martingale by sequentially wagering on treatment assignments given observed outcomes. Under the null hypothesis of no treatment effect, the expected wealth cannot grow, guaranteeing anytime-valid Type I error control regardless of stopping rule. We prove validity and present simulation studies demonstrating calibration and power. The e-RT provides a conservative, assumption-free complement to model-based sequential analyzes.
Keywords: e-values, e-process, randomization test, sequential analysis, clinical trials
1 Introduction
Sequential monitoring of randomized controlled trials requires methods that control Type I errors regardless of when or why the monitoring stops. Traditional group-sequential designs rely on parametric assumptions and predetermined stopping boundaries. When these assumptions fail, or when trials adapt in ways not fully prespecified, validation guaranties may erode.
E-values and e-processes offer an alternative framework (shafer2021; vovk2021; ramdas2025). An e-value is a measure of evidence against a null hypothesis with a specific property: its expected value under the null is at most 1. This simple constraint yields anytime-valid inference: the Type I error guaranty holds at any stopping time, regardless of the stopping rule.
We present the randomization e-process (e-RT), a sequential test that derives validity solely from the trial’s randomization mechanism. The construction uses a betting framework: at each enrollment, we wager on the treatment assignment given the observed outcome. Under the null hypothesis of no treatment effect, treatment assignment is independent of outcome, and no betting strategy can systematically grow wealth. This independence, guaranteed by randomization, is the only assumption required.
The contributions of this note are as follows.
  1. 1.
    An explicit construction of e-RT with proof of validity
  2. 2.
    Simulation studies demonstrating Type I error calibration and power
2 The e-RT Construction
2.1 Setup
Consider a sequential randomized trial. At each enrollment $i=1,2,\ldots$, we observe $T_{i}$: assignment of treatment (control or intervention) and $Y_{i}$: binary outcome (event or no event). Treatment is assigned with known probability $p$, typically $p=0.5$ for 1:1 randomization.
The null hypothesis is that treatment has no effect on any patient. This is the sharp null: every patient would have had the same outcome regardless of assignment. Under this hypothesis, knowing the outcome of a patient tells you nothing about which arm they were assigned to – outcome and treatment are independent ($Y_{i}\perp T_{i}$).
2.2 Betting Construction
e-RT constructs a wealth process $W_{n}$ by betting sequentially on treatment assignments.
For each patient, after observing their outcome but before learning their treatment assignment, we place a bet. The bet is a fraction $\lambda_{i}\in[0,1]$ representing how much we wager on the patient being in the treatment arm.
Under the null hypothesis, outcome tells us nothing about assignment; hence we can’t systematically win. Under the alternative, outcomes differ by arm, so we can bet profitably.
The wealth updates as:
$W_{i}=W_{i-1}\times\begin{cases}\lambda_{i}/p&\text{if }T_{i}=\text{intervention}\\ (1-\lambda_{i})/(1-p)&\text{if }T_{i}=\text{control}\end{cases}$ (1)
starting from $W_{0}=1$. Note that the intervention and control here are simply labels, they could be called whatever (A versus B, etc).
2.3 Adaptive Betting Strategy
The betting strategy uses accumulated data to inform each bet. Here is how it works:
  1. 1.
    A patient is randomized (assignment blinded)
  2. 2.
    Outcome is observed
  3. 3.
    You check the historical records and see in which arm this outcome is more likely
  4. 4.
    You bet
  5. 5.
    Assignment is revealed
  6. 6.
    Wealth updated. If guess is correct, it grows; if wrong, it shrinks
Under the null hypothesis, any historical imbalance between arms is noise. Your bets are sometimes right, sometimes wrong, and wealth fluctuates randomly around 1. Under the alternative, the imbalance is real signal. Your bets are systematically correct, and wealth grows over time.
Formally, let:
$\hat{\delta}_{i-1}=\text{(event rate in treatment)}-\text{(event rate in control)}$ (2)
estimated from all patients before $i$. The betting fraction is:
$\lambda_{i}=\begin{cases}0.5+0.5\cdot c_{i}\cdot\hat{\delta}_{i-1}&\text{if }Y_{i}=\text{event}\\ 0.5-0.5\cdot c_{i}\cdot\hat{\delta}_{i-1}&\text{if }Y_{i}=\text{no event}\end{cases}$ (3)
where $c_{i}\in[0,1]$ is a confidence parameter that ramps up from 0 to 1 over a burn-in period, preventing wild bets when data are sparse:
$c_{i}=\min\left(1,\frac{i-\text{burn-in}}{\text{ramp}}\right)$ (4)
The above can be read as: At any specific point, the betting fraction $\lambda$ is defined as $0.5$ plus the wager if there has been an event and $0.5$ minus the wager if there hasn’t been an event. The wager is determined by the observed effect until now ($\hat{\delta}_{i-1}$) multiplied by a factor (that is optional) called $c_{1}$. This factor is simply a factor that avoids us from making full bets early on. It is determined by a burn-in period (where we don’t bet, just collect data), and a ramp, which is also arbitrarily defined, so that wager is "muffled" early on.
2.4 Worked Example
Consider a trial comparing intervention versus control, with a binary outcome (event or no event). Event is mortality, which is expected to be lower with intervention. Allocation is 1:1 ($p=0.5$). Assume burn-in is complete ($c=1$).
We look back at patients 1–199. Intervention arm has 100 patients, 35 events, so rate = 35.0%. Control arm has 99 patients, 40 events, so rate = 40.4%. $\hat{\delta}_{199}=0.350-0.404=-0.054$ (intervention looks protective).
Patient 200 has an event (dies). Where is this patient likely from? Events are more common in control (40.4% vs 35.0%), so probably control. We bet $\lambda=0.5+0.5\times(-0.054)=0.473$ on intervention and $1-\lambda=0.527$ on control. Assignment revealed and patient indeed is in control group. We guessed right! Our multiplier will be $0.527/0.5=1.054$, hence wealth grows 5.4%.
Now patient 201 has been enrolled. First, we update the counts with patient 200’s data. Intervention arm now has 100 patients, 35 events (rate = 35.0%), and control arm has 100 patients, 41 events (rate = 41.0%). The $\hat{\delta}_{200}=0.350-0.410=-0.060$. Patient 201 does not have an event. Non-events are more common in intervention (65.0% vs 59.0%), so probably intervention. We will bet once again. This time we bet $\lambda=0.5-0.5\times(-0.060)=0.530$ on intervention and $1-\lambda=0.470$ on control. Finally, assignment is revealed and we guessed right again: patient is in intervention. Our multiplier =0.530/0.5=1.060=0.530/0.5=1.060. Wealth grows another 6.0%.
Finally, patient 202 has been enrolled. We update counts once more with intervention having 101 patients and 35 events (rate = 34.7%) and control having 100 patients with 41 events (rate = 41.0%). Our $\hat{\delta}_{201}=0.347-0.410=-0.063$. Patient 202 has an event (dies). Events cluster in control, so we bet toward control: $\lambda=0.5+0.5\times(-0.063)=0.469$ on intervention and $1-\lambda=0.531$ on control. Assignment is revealed and we guessed wrong—patient is in intervention group! Our multiplier will be $0.469/0.5=0.938$. Wealth shrinks 6.2%.
Cumulative wealth is the product of all multipliers:
$W_{202}=W_{199}\times 1.054\times 1.060\times 0.938=W_{199}\times 1.048$ (5)
Despite one wrong guess, wealth grew 4.8% over these three patients.
Even when intervention works, some intervention patients die. We guess wrong sometimes. But if intervention truly reduces mortality, events cluster in control and non-events cluster in intervention — we are right more often than wrong, and wealth grows on average.
Under the null hypothesis (no effect), right and wrong guesses balance out. Wealth fluctuates randomly around 1.
2.5 Validity
Theorem 1.
Under the null hypothesis, the wealth process $(W_{n})$ is a test martingale — its expected value cannot grow.
Proof.
Under the null, outcome and treatment are independent. So after seeing the outcome, treatment assignment is still just a coin flip with probability $p$. The expected wealth multiplier is:
$\mathbb{E}[\text{multiplier}]$ $=p\times\frac{\lambda}{p}+(1-p)\times\frac{1-\lambda}{1-p}$ (6)
$=\lambda+(1-\lambda)$ (7)
$=1$ (8)
No matter how we choose our bet $\lambda$, we can’t systematically grow wealth when the null is true. ∎
Corollary 1.
For any stopping time $\tau$, $\mathbb{E}[W_{\tau}]\leq 1$ under the null. By Markov’s inequality, $P(W_{\tau}\geq 1/\alpha)\leq\alpha$. Thus rejecting when $W_{\tau}\geq 1/\alpha$ controls Type I error at level $\alpha$, regardless of the stopping rule.
3 Simulation Studies
We evaluated e-RT operating characteristics by simulation. For each scenario, we calculated the sample size required for a chi-square test to achieve the target power at $\alpha=0.05$, then ran 5,000 simulated trials at that sample size. We used burn-in = 50 patients and ramp = 100 patients. Control arm event rate was 40% in all scenarios.
3.1 Results
Table 1 presents Type I error and power for trials designed to detect 5% or 10% absolute risk reductions (ARR) with 80% or 90% power.
Table 1: e-RT Operating Characteristics
ARR Target Power $n$ Type I Error e-RT Power Median Crossing
5% 80% 2942 0.032 48.6% 1392 (47%)
10% 80% 712 0.021 50.4% 401 (56%)
5% 90% 3938 0.035 62.8% 1842 (47%)
10% 90% 954 0.025 65.9% 478 (50%)
Type I error was well controlled across all scenarios (0.021 – 0.035), below the nominal $\alpha=0.05$. This confirms the theoretical guarantee from the martingale property.
Power was approximately 50% for trials designed with 80% power, and 63–66% for trials designed with 90% power. When the e-RT rejected the null, it did so at approximately half the planned sample size (median crossing 47–56% of total enrollment).
3.2 Interpretation
The e-RT is not a replacement for the primary analysis—it is a continuous monitoring tool. A trial designed for 80% power with a traditional test gains a 50% chance of stopping early via e-RT, at roughly the halfway point. If the threshold is not crossed, the trial proceeds to completion and the planned analysis is conducted.
This represents a free option: anytime-valid monitoring with no alpha spending and no pre-specified interim looks. The “cost” is that the e-RT alone has lower power than a fixed-sample test. But when layered on top of a properly powered trial, it provides early stopping when effects are larger than anticipated.
3.3 Trajectory Examples
Figure 1 shows example e-RT trajectories from 30 simulated trials under the null hypothesis (both arms 40% event rate). Sample sizes correspond to trials designed for 80% power (n = 712, left) and 90% power (n = 954, right) to detect a 10% ARR. Under the null, wealth fluctuates randomly around 1. Some trajectories temporarily rise toward the threshold but eventually drift back down. The downward drift over time reflects the accumulating “cost” of betting on noise.
Figure 1: e-RT trajectories under the null hypothesis. Panel A: n = 712 (80% power design). Panel B: n = 954 (90% power design). Dashed red line: rejection threshold ($1/\alpha=20$). Dotted gray line: neutral (wealth = 1). Under the null, no trajectory crosses the threshold.
Figure 2 shows trajectories under the alternative hypothesis (40% vs 30% event rates, true ARR = 10%). With a real treatment effect, wealth grows systematically. Most trajectories cross the rejection threshold before enrollment completes, and many reach values far exceeding 20 — providing strong evidence against the null.
Figure 2: e-RT trajectories under the alternative hypothesis (true ARR = 10%). Panel A: n = 712 (80% power design). Panel B: n = 954 (90% power design). Dashed red line: rejection threshold ($1/\alpha=20$). Under the alternative, approximately half of trajectories cross the threshold, typically around the midpoint of enrollment.
4 Discussion
We have presented the e-RT, a nonparametric sequential test for randomized trials based on the betting framework for e-values. The method requires only that treatment assignment is randomized — no distributional assumptions about outcomes are needed. This makes it a robust complement to model-based analyses.
4.1 Operating characteristics
Our simulations demonstrate that e-RT controls Type I error conservatively (2–3% vs nominal 5%) while providing approximately 50% power for early stopping in trials designed with 80% power, and 63–66% power in trials designed with 90% power. When early stopping occurs, it happens at roughly half the planned sample size.
These results should be interpreted carefully. The e-RT is not designed to replace the primary analysis of a trial; at least not yet. Rather, it provides a continuous monitoring option that requires no alpha spending and no pre-specified interim analysis schedule. If the e-RT crosses its threshold, one could consider stopping the trial. If it does not cross, the trial may proceeds to its planned conclusion and primary analysis.
4.2 Relationship to existing work
The betting framework for hypothesis testing was developed by shafer2021. E-values and e-processes have been extensively studied (vovk2021; ramdas2022; ramdas2025). koning2025 develops e-values for group invariance, including permutation tests, using batch-based likelihood ratio statistics normalized by permutation expectations. Our construction differs in using patient-by-patient betting with adaptive learning of the effect size. These approaches share theoretical foundations but yield different procedures.
4.3 Limitations
Several limitations should be noted. This is an experimental method under development. So far, this has only been tested for binary outcomes. It is reasonable to assume that extending to continuous endpoints would be feasible; this is planned. It is uncertain how this method could be applied for time-to-event endpoints. Hypothetically, we could hold the bet value for all patients randomized in a period until outcomes arrive, and approaches such as restricted mean survival time could be used. This is also under development.
There are other limitations. The e-RT tests only whether event rates differ between arms. It does not directly estimate treatment effects or provide confidence intervals. The adaptive learning of $\hat{\delta}$ requires a burn-in period during which little evidence accumulates. Third, for trials where parametric assumptions are plausible, model-based sequential methods will generally have better power. Fourth, our simulations used a specific betting strategy (burn-in = 50, ramp = 100); other choices may yield different operating characteristics. Hypothetically, it could be possible to increase betting aggressiveness over time as evidence accumulates; this may result in better power. Finally, it is unclear how this method will behave in situations where heterogeneity in treatment effects exist or there are temporal instabilities in effect size. The method could be extended to bet according to relative effect size approaches, such as the odds ratio. This is too under development.
4.4 Conclusion
The e-RT provides anytime-valid sequential inference for randomized trials using only the guarantee of randomization. Its validity is unconditional on the data-generating process, making it a robust tool for trial monitoring. While it trades power for this robustness, it offers a valuable complement to traditional analysis methods.
4.5 Disclaimer
This is an experimental method under development. Application to real patients should only be considered under surveillance from an experienced statistician. The author is not responsible for consequences of use of this method.
Appendix A R Code
# e-RT Simulations: Type I Error and Power (Unified)
library(tidyverse)
# Core e-RT function
compute_eRT <- function(treatment, outcome, p = 0.5, burn_in = 50, ramp = 100) {
n <- length(treatment)
if (is.factor(treatment)) treatment <- as.numeric(treatment) - 1
if (is.factor(outcome)) outcome <- as.numeric(outcome) - 1
wealth <- numeric(n)
wealth[1] <- 1
for (i in 2:n) {
trt_prev <- treatment[1:(i-1)]
out_prev <- outcome[1:(i-1)]
rate_trt <- mean(out_prev[trt_prev == 1])
rate_ctrl <- mean(out_prev[trt_prev == 0])
if (is.nan(rate_trt)) rate_trt <- 0.5
if (is.nan(rate_ctrl)) rate_ctrl <- 0.5
delta_hat <- rate_trt - rate_ctrl
c_i <- max(0, min(1, (i - burn_in) / ramp))
if (outcome[i] == 1) {
lambda <- 0.5 + 0.5 * c_i * delta_hat
} else {
lambda <- 0.5 - 0.5 * c_i * delta_hat
lambda <- max(0.001, min(0.999, lambda))
if (treatment[i] == 1) {
multiplier <- lambda / p
} else {
multiplier <- (1 - lambda) / (1 - p)
wealth[i] <- wealth[i-1] * multiplier
return(wealth)
# Simulate a single trial
simulate_trial <- function(n, p_trt = 0.5, rate_trt, rate_ctrl) {
treatment <- rbinom(n, 1, p_trt)
outcome <- numeric(n)
outcome[treatment == 1] <- rbinom(sum(treatment == 1), 1, rate_trt)
outcome[treatment == 0] <- rbinom(sum(treatment == 0), 1, rate_ctrl)
return(data.frame(treatment = treatment, outcome = outcome))
# Unified simulation function
simulate_eRT <- function(n_sims = 5000,
p_ctrl,
p_trt = NULL,
hypothesized_ARR = NULL,
target_power = 0.80,
alpha = 0.05,
burn_in = 50,
ramp = 100) {
# Determine hypothesized ARR for sample size calculation
if (is.null(hypothesized_ARR) && is.null(p_trt)) {
stop("Mustspecifyeitherp_trtorhypothesized_ARR")
if (is.null(hypothesized_ARR)) {
hypothesized_ARR <- p_ctrl - p_trt
# Calculate sample size based on hypothesized effect
$s$ <- power.prop.test(p1 = p_ctrl, p2 = p_ctrl - hypothesized_ARR,
power = target_power, sig.level = alpha)
n_per_arm <- ceiling(ss\$n)
n_patients <- 2 * n_per_arm
# Determine true p_trt (null if not specified)
if (is.null(p_trt)) {
p_trt <- p_ctrl # Null is true
true_ARR <- 0
sim_type <- "TypeIError"
} else {
true_ARR <- p_ctrl - p_trt
sim_type <- "Power"
[…]
p_ctrl, hypothesized_ARR, target_power * 100))
[…]
p_ctrl, p_trt, true_ARR))
[…]
rejections <- 0
first_crossing <- numeric(n_sims)
final_evalues <- numeric(n_sims)
pb <- txtProgressBar(min = 0, max = n_sims, style = 3)
for (sim in 1:n_sims) {
trial <- simulate_trial(n_patients, rate_trt = p_trt, rate_ctrl = p_ctrl)
wealth <- compute_eRT(trial\$treatment, trial\$outcome, burn_in = burn_in, ramp = ramp)
final_evalues[sim] <- wealth[n_patients]
crossing <- which(wealth >= 1/alpha)
if (length(crossing) > 0) {
rejections <- rejections + 1
first_crossing[sim] <- crossing[1]
} else {
first_crossing[sim] <- NA
setTxtProgressBar(pb, sim)
close(pb)
rejection_rate <- rejections / n_sims
se <- sqrt(rejection_rate * (1 - rejection_rate) / n_sims)
median_stop <- median(first_crossing, na.rm = TRUE)
[…]
if (sim_type == "Power") {
[…]
} else {
[…]
return(list(
sim_type = sim_type,
rejection_rate = rejection_rate,
se = se,
n_per_arm = n_per_arm,
n_patients = n_patients,
p_ctrl = p_ctrl,
p_trt = p_trt,
hypothesized_ARR = hypothesized_ARR,
true_ARR = true_ARR,
target_power = target_power,
alpha = alpha,
median_crossing = median_stop,
first_crossing = first_crossing,
final_evalues = final_evalues
# Plot trajectories
plot_trajectories <- function(n_trials = 30,
p_ctrl,
p_trt = NULL,
hypothesized_ARR = NULL,
target_power = 0.80,
alpha = 0.05,
burn_in = 50,
ramp = 100,
title = NULL) {
if (is.null(hypothesized_ARR) && is.null(p_trt)) {
stop("Mustspecifyeitherp_trtorhypothesized_ARR")
if (is.null(hypothesized_ARR)) {
hypothesized_ARR <- p_ctrl - p_trt
if (is.null(p_trt)) {
p_trt <- p_ctrl # Null
# Calculate sample size
$s$ <- power.prop.test(p1 = p_ctrl, p2 = p_ctrl - hypothesized_ARR,
power = target_power, sig.level = alpha)
n_patients <- 2 * ceiling(ss\$n)
true_ARR <- p_ctrl - p_trt
if (is.null(title)) {
if (true_ARR == 0) {
title <- sprintf("e-RTUnderNull(n=%d,designedfor%.0f%%ARR)",
n_patients, hypothesized_ARR * 100)
} else {
title <- sprintf("e-RTUnderAlternative(n=%d,%.0f%%power)",
n_patients, target_power * 100)
trajectories <- list()
for (i in 1:n_trials) {
trial <- simulate_trial(n_patients, rate_trt = p_trt, rate_ctrl = p_ctrl)
wealth <- compute_eRT(trial\$treatment, trial\$outcome, burn_in = burn_in, ramp = ramp)
trajectories[[i]] <- data.frame(
patient = 1:n_patients,
wealth = wealth,
trial = i
df <- bind_rows(trajectories)
p <- ggplot(df, aes(x = patient, y = wealth, group = trial)) +
geom_line(alpha = 0.4, color = "steelblue") +
geom_hline(yintercept = 1/alpha, linetype = "dashed", color = "red") +
geom_hline(yintercept = 1, linetype = "dotted", color = "gray50") +
scale_y_log10() +
labs(
title = title,
subtitle = sprintf("Control=%.0f%%,Treatment=%.0f%%,TrueARR=%.0f%%",
p_ctrl * 100, p_trt * 100, true_ARR * 100),
x = "Patient",
y = "e-value(logscale)"
annotate("text", x = n_patients * 0.95, y = 1/alpha * 1.5,
label = sprintf("1/alpha=%.0f", 1/alpha), color = "red", hjust = 1) +
theme_minimal() +
theme(
plot.title = element_text(face = "bold"),
panel.grid.minor = element_blank()
return(p)
# Run all simulations
run_all <- function(n_sims = 5000,
p_ctrl = 0.40,
ARRs = c(0.05, 0.10),
target_powers = c(0.80, 0.90),
alpha = 0.05,
seed = 42) {
set.seed(seed)
results <- list()
# Create grid of scenarios
scenarios <- expand.grid(
hypothesized_ARR = ARRs,
target_power = target_powers,
stringsAsFactors = FALSE
# For each scenario, run Type I error and Power
all_results <- data.frame()
for (i in 1:nrow(scenarios)) {
hyp_ARR <- scenarios\$hypothesized_ARR[i]
tgt_pow <- scenarios\$target_power[i]
[…]
hyp_ARR * 100, tgt_pow * 100))
[…]
# Type I error (null true, same sample size)
[…]
t1 <- simulate_eRT(
n_sims = n_sims,
p_ctrl = p_ctrl,
p_trt = NULL, # Null is true
hypothesized_ARR = hyp_ARR,
target_power = tgt_pow,
alpha = alpha
# Power (alternative true)
[…]
pow <- simulate_eRT(
n_sims = n_sims,
p_ctrl = p_ctrl,
p_trt = p_ctrl - hyp_ARR, # True effect = hypothesized
target_power = tgt_pow,
alpha = alpha
all_results <- rbind(all_results, data.frame(
hypothesized_ARR = hyp_ARR,
target_power = tgt_pow,
n_patients = t1\$n_patients,
type1_error = t1\$rejection_rate,
type1_se = t1\$se,
eRT_power = pow\$rejection_rate,
power_se = pow\$se,
median_crossing = pow\$median_crossing
results\$summary <- all_results
# Print summary
[…]
print(all_results %>%
mutate(
hypothesized_ARR = sprintf("%.0f%%", hypothesized_ARR * 100),
target_power = sprintf("%.0f%%", target_power * 100),
type1_error = sprintf("%.3f", type1_error),
eRT_power = sprintf("%.1f%%", eRT_power * 100)
select(hypothesized_ARR, target_power, n_patients,
type1_error, eRT_power, median_crossing))
return(results)
# Run
if (interactive()) {
results <- run_all(
n_sims = 5000,
p_ctrl = 0.40,
ARRs = c(0.05, 0.10),
target_powers = c(0.80, 0.90),
alpha = 0.05
# Trajectory plots - Alternative
p1 <- plot_trajectories(p_ctrl = 0.40, p_trt = 0.30, target_power = 0.80)
ggsave("traj_alt_10pct_80pow.pdf", p1, width = 8, height = 5)
p2 <- plot_trajectories(p_ctrl = 0.40, p_trt = 0.30, target_power = 0.90)
ggsave("traj_alt_10pct_90pow.pdf", p2, width = 8, height = 5)
# Trajectory plots - Null
p3 <- plot_trajectories(p_ctrl = 0.40, p_trt = NULL, hypothesized_ARR = 0.10, target_power = 0.80)
ggsave("traj_null_10pct_80pow.pdf", p3, width = 8, height = 5)
p4 <- plot_trajectories(p_ctrl = 0.40, p_trt = NULL, hypothesized_ARR = 0.10, target_power = 0.90)
ggsave("traj_null_10pct_90pow.pdf", p4, width = 8, height = 5)

### v2

@@
 We prove validity and present simulation studies demonstrating calibration and power.
-The e-RT provides a conservative, assumption-free complement to model-based sequential analyzes.
+The e-RT provides a conservative, assumption-free complement to model-based sequential analyses.
 Keywords: e-values, e-process, randomization test, sequential analysis, clinical trials
@@
 Traditional group-sequential designs rely on parametric assumptions and predetermined stopping boundaries.
-When these assumptions fail, or when trials adapt in ways not fully prespecified, validation guaranties may erode.
+When these assumptions fail, or when trials adapt in ways not fully prespecified, validation guarantees may erode.
 E-values and e-processes offer an alternative framework (shafer2021; vovk2021; ramdas2025).
 An e-value is a measure of evidence against a null hypothesis with a specific property: its expected value under the null is at most 1.
-This simple constraint yields anytime-valid inference: the Type I error guaranty holds at any stopping time, regardless of the stopping rule.
+This simple constraint yields anytime-valid inference: the Type I error guarantee holds at any stopping time, regardless of the stopping rule.
 We present the randomization e-process (e-RT), a sequential test that derives validity solely from the trial’s randomization mechanism.
@@
 Treatment is assigned with known probability $p$, typically $p=0.5$ for 1:1 randomization.
-The null hypothesis is that treatment has no effect on any patient.
-This is the sharp null: every patient would have had the same outcome regardless of assignment.
+The null hypothesis is that treatment has no effect.
 Under this hypothesis, knowing the outcome of a patient tells you nothing about which arm they were assigned to – outcome and treatment are independent ($Y_{i}\perp T_{i}$).
@@
   1. 1.
-    A patient is randomized (assignment blinded)
+    A patient is randomized (assignment blinded).
   2. 2.
-    Outcome is observed
+    Outcome is observed.
   3. 3.
-    You check the historical records and see in which arm this outcome is more likely
+    Historical records are checked. The most likely arm for this outcome is defined.
   4. 4.
-    You bet
+    Bet is placed according to previous data.
   5. 5.
-    Assignment is revealed
+    Assignment is revealed.
   6. 6.
-    Wealth updated. If guess is correct, it grows; if wrong, it shrinks
+    Wealth updated. If guess is correct, it grows; if wrong, it shrinks.
 Under the null hypothesis, any historical imbalance between arms is noise.
-Your bets are sometimes right, sometimes wrong, and wealth fluctuates randomly around 1.
+Bets will be sometimes right, sometimes wrong, and wealth fluctuates randomly around 1.
 Under the alternative, the imbalance is real signal.
-Your bets are systematically correct, and wealth grows over time.
+Bets will be systematically correct, and wealth grows over time.
 Formally, let:
@@
 The above can be read as: At any specific point, the betting fraction $\lambda$ is defined as $0.5$ plus the wager if there has been an event and $0.5$ minus the wager if there hasn’t been an event.
-The wager is determined by the observed effect until now ($\hat{\delta}_{i-1}$) multiplied by a factor (that is optional) called $c_{1}$.
+The wager is determined by the observed effect until now ($\hat{\delta}_{i-1}$) multiplied by a factor (that is optional) called $c_{i}$.
 This factor is simply a factor that avoids us from making full bets early on.
 It is determined by a burn-in period (where we don’t bet, just collect data), and a ramp, which is also arbitrarily defined, so that wager is "muffled" early on.
+Finally, the $0.5$ multiplier before $c_{i}$ and the wager ($\hat{\delta}_{i-1}$) simply keeps $\lambda$ between zero and one.
 2.4 Worked Example
@@
 Under the null hypothesis (no effect), right and wrong guesses balance out. Wealth fluctuates randomly around 1.
+In summary: the outcome determines $\lambda$ (which direction to bet); the treatment arm determines which multiplier formula to apply (whether the bet paid off).
+The machinery is agnostic to whether the effect size is following the expected direction; it simply keeps betting according to what is more likely.
 2.5 Validity
@@
 Figure 1: e-RT trajectories under the null hypothesis.
-Panel A: n = 712 (80% power design).
-Panel B: n = 954 (90% power design).
+Left: n = 712 (80% power design).
+Right: n = 954 (90% power design).
 Dashed red line: rejection threshold ($1/\alpha=20$).
@@
 4 Discussion
-We have presented the e-RT, a nonparametric sequential test for randomized trials based on the betting framework for e-values.
+The e-RT is a nonparametric sequential test for randomized trials based on the betting framework for e-values.
 The method requires only that treatment assignment is randomized — no distributional assumptions about outcomes are needed.
@@
 4.1 Operating characteristics
-Our simulations demonstrate that e-RT controls Type I error conservatively (2–3% vs nominal 5%) while providing approximately 50% power for early stopping in trials designed with 80% power, and 63–66% power in trials designed with 90% power.
+Simulations demonstrate that e-RT controls Type I error conservatively (2–3% vs nominal 5%) while providing approximately 50% power for early stopping in trials designed with 80% power, and 63–66% power in trials designed with 90% power.
 When early stopping occurs, it happens at roughly half the planned sample size.
@@
 The e-RT is not designed to replace the primary analysis of a trial; at least not yet.
-Rather, it provides a continuous monitoring option that requires no alpha spending and no pre-specified interim analysis schedule.
+Rather, it provides a continuous monitoring option that requires no alpha spending and no prespecified interim analysis schedule.
 If the e-RT crosses its threshold, one could consider stopping the trial.
-If it does not cross, the trial may proceeds to its planned conclusion and primary analysis.
+If it does not cross, the trial may proceed to its planned conclusion and primary analysis.
 4.2 Relationship to existing work
@@
 The method could be extended to bet according to relative effect size approaches, such as the odds ratio.
-This is too under development.
+This is also under development.
 4.4 Conclusion
@@
 The author is not responsible for consequences of use of this method.
+4.6 Version Control
+  1. 1.
+    First Version (Dec 04, 2025)
+  2. 2.
+    Second Version (this version): Minor text adjustments; removed claim on sharp null.
 Appendix A R Code

### v3

@@
-Sequential Randomization Tests Using E-values: A Betting Approach for Clinical Trials
+Sequential Randomization Tests Using e-values: Applications for trial monitoring
 Abstract
 Sequential monitoring of randomized trials traditionally relies on parametric assumptions or asymptotic approximations.
-We present a nonparametric sequential test – the randomization e-process (e-RT) – that derives validity solely from the randomization mechanism.
-Using a betting framework, e-RT constructs a test martingale by sequentially wagering on treatment assignments given observed outcomes.
+We discuss a nonparametric sequential test and its application to continuous and time-to-event endpoints that derives validity solely from the randomization mechanism.
+Using a betting framework, these tests constructs a test martingale by sequentially wagering on treatment assignments given observed outcomes.
 Under the null hypothesis of no treatment effect, the expected wealth cannot grow, guaranteeing anytime-valid Type I error control regardless of stopping rule.
 We prove validity and present simulation studies demonstrating calibration and power.
-The e-RT provides a conservative, assumption-free complement to model-based sequential analyses.
+These methods provide a conservative, assumption-free complement to model-based sequential analyses.
 Keywords: e-values, e-process, randomization test, sequential analysis, clinical trials
@@
 Traditional group-sequential designs rely on parametric assumptions and predetermined stopping boundaries.
-When these assumptions fail, or when trials adapt in ways not fully prespecified, validation guarantees may erode.
+When these assumptions fail, or when trials adapt in ways not fully prespecified, validity guarantees may erode.
+Monitoring clinical trials, specially in the context of acutely ill patients, is of paramount importance.
+Traditional methods include $\alpha-spending$ functions.
+When interim analyses are spaced, evidence grows unnoticed between the interim analyses.
+This may both delay the implementation of potentially useful strategies or prolong trials when safety signals arise early.
 E-values and e-processes offer an alternative framework (shafer2021; vovk2021; ramdas2025).
@@
 This simple constraint yields anytime-valid inference: the Type I error guarantee holds at any stopping time, regardless of the stopping rule.
-We present the randomization e-process (e-RT), a sequential test that derives validity solely from the trial’s randomization mechanism.
-The construction uses a betting framework: at each enrollment, we wager on the treatment assignment given the observed outcome.
-Under the null hypothesis of no treatment effect, treatment assignment is independent of outcome, and no betting strategy can systematically grow wealth.
-This independence, guaranteed by randomization, is the only assumption required.
-The contributions of this note are as follows.
-  1. 1.
-    An explicit construction of e-RT with proof of validity
-  2. 2.
-    Simulation studies demonstrating Type I error calibration and power
-2 The e-RT Construction
+pmlr-v177-duan22a introduced interactive rank testing by betting (i-bet), which tests treatment effects by wagering on treatment assignments given observed outcomes.
+The intuition is discussed by ramdas2021game.
+Under the null hypothesis, randomization ensures that assignments are independent of outcomes, so no betting strategy can systematically accumulate wealth.
+We discuss potential methods for trial monitoring based on previous efforts.
+For binary outcomes, our construction implements the i-bet framework with a specific adaptive betting strategy.
+We present methods with unified notation, validity proofs, and simulation studies demonstrating Type I error control and power characteristics.
+2 Overall construction
 2.1 Setup
-Consider a sequential randomized trial.
-At each enrollment $i=1,2,\ldots$, we observe $T_{i}$: assignment of treatment (control or intervention) and $Y_{i}$: binary outcome (event or no event).
-Treatment is assigned with known probability $p$, typically $p=0.5$ for 1:1 randomization.
-The null hypothesis is that treatment has no effect.
-Under this hypothesis, knowing the outcome of a patient tells you nothing about which arm they were assigned to – outcome and treatment are independent ($Y_{i}\perp T_{i}$).
-2.2 Betting Construction
-e-RT constructs a wealth process $W_{n}$ by betting sequentially on treatment assignments.
-For each patient, after observing their outcome but before learning their treatment assignment, we place a bet.
-The bet is a fraction $\lambda_{i}\in[0,1]$ representing how much we wager on the patient being in the treatment arm.
-Under the null hypothesis, outcome tells us nothing about assignment; hence we can’t systematically win. Under the alternative, outcomes differ by arm, so we can bet profitably.
+Consider a sequential randomized trial with 1:1 allocation. At each enrollment $i=1,2,\ldots$, we observe:
+  • $T_{i}\in\{0,1\}$: treatment assignment (0 = control, 1 = intervention)
+  • $Y_{i}\in\{0,1\}$: binary outcome (0 = no event, 1 = event)
+Treatment is assigned with known probability $p=P(T_{i}=1)$, typically $p=0.5$.
+The null hypothesis is that treatment assignment has no effect on outcome:
+$H_{0}:Y_{i}\perp T_{i}\text{ for all }i$ (1)
+Under this hypothesis, observing the outcome provides no information about which arm the patient was assigned to.
+2.2 Wealth Process
+Following pmlr-v177-duan22a, we construct a wealth process by wagering on treatment assignments.
+After observing outcome $Y_{i}$ but before learning treatment assignment $T_{i}$, we choose $\lambda_{i}\in[0,1]$: the fraction wagered on intervention.
 The wealth updates as:
-$W_{i}=W_{i-1}\times\begin{cases}\lambda_{i}/p&\text{if }T_{i}=\text{intervention}\\ (1-\lambda_{i})/(1-p)&\text{if }T_{i}=\text{control}\end{cases}$ (1)
-starting from $W_{0}=1$. Note that the intervention and control here are simply labels, they could be called whatever (A versus B, etc).
-2.3 Adaptive Betting Strategy
-The betting strategy uses accumulated data to inform each bet. Here is how it works:
-  1. 1.
-    A patient is randomized (assignment blinded).
-  2. 2.
-    Outcome is observed.
-  3. 3.
-    Historical records are checked. The most likely arm for this outcome is defined.
-  4. 4.
-    Bet is placed according to previous data.
-  5. 5.
-    Assignment is revealed.
-  6. 6.
-    Wealth updated. If guess is correct, it grows; if wrong, it shrinks.
-Under the null hypothesis, any historical imbalance between arms is noise.
-Bets will be sometimes right, sometimes wrong, and wealth fluctuates randomly around 1.
-Under the alternative, the imbalance is real signal.
-Bets will be systematically correct, and wealth grows over time.
-Formally, let:
-$\hat{\delta}_{i-1}=\text{(event rate in treatment)}-\text{(event rate in control)}$ (2)
-estimated from all patients before $i$. The betting fraction is:
-$\lambda_{i}=\begin{cases}0.5+0.5\cdot c_{i}\cdot\hat{\delta}_{i-1}&\text{if }Y_{i}=\text{event}\\ 0.5-0.5\cdot c_{i}\cdot\hat{\delta}_{i-1}&\text{if }Y_{i}=\text{no event}\end{cases}$ (3)
-where $c_{i}\in[0,1]$ is a confidence parameter that ramps up from 0 to 1 over a burn-in period, preventing wild bets when data are sparse:
-$c_{i}=\min\left(1,\frac{i-\text{burn-in}}{\text{ramp}}\right)$ (4)
-The above can be read as: At any specific point, the betting fraction $\lambda$ is defined as $0.5$ plus the wager if there has been an event and $0.5$ minus the wager if there hasn’t been an event.
-The wager is determined by the observed effect until now ($\hat{\delta}_{i-1}$) multiplied by a factor (that is optional) called $c_{i}$.
-This factor is simply a factor that avoids us from making full bets early on.
-It is determined by a burn-in period (where we don’t bet, just collect data), and a ramp, which is also arbitrarily defined, so that wager is "muffled" early on.
-Finally, the $0.5$ multiplier before $c_{i}$ and the wager ($\hat{\delta}_{i-1}$) simply keeps $\lambda$ between zero and one.
+$W_{i}=W_{i-1}\times\begin{cases}\lambda_{i}/p&\text{if }T_{i}=1\\ (1-\lambda_{i})/(1-p)&\text{if }T_{i}=0\end{cases}$ (2)
+starting from $W_{0}=1$. When we bet toward the correct arm, wealth grows; when wrong, it shrinks.
+2.3 Betting Strategy
+The validity guarantee holds for any betting strategy where $\lambda_{i}$ depends only on $\mathcal{F}_{i-1}$.
+Power depends on choosing bets that grow wealth under the alternative.
+We use a strategy that learns the treatment effect from accumulating data.
+Let:
+$\hat{\delta}_{i-1}=\text{(event rate in intervention)}-\text{(event rate in control)}$ (3)
+estimated from patients $1,\ldots,i-1$. The betting fraction is:
+$\lambda_{i}=\begin{cases}0.5+0.5\cdot c_{i}\cdot\hat{\delta}_{i-1}&\text{if }Y_{i}=1\\ 0.5-0.5\cdot c_{i}\cdot\hat{\delta}_{i-1}&\text{if }Y_{i}=0\end{cases}$ (4)
+where $c_{i}\in[0,1]$ ramps from 0 to 1 over a burn-in period:
+$c_{i}=\min\left(1,\max\left(0,\frac{i-n_{0}}{n_{r}}\right)\right)$ (5)
+with $n_{0}$ the burn-in period and $n_{r}$ the ramp period. This prevents large bets when $\hat{\delta}_{i-1}$ is unstable due to small samples.
+The logic: if $\hat{\delta}>0$ (more events in intervention), then events suggest intervention and non-events suggest control.
+If $\hat{\delta}<0$ (fewer events in intervention), then events suggest control and non-events suggest intervention.
+The factor of $0.5$ before $c_{i}\cdot\hat{\delta}_{i-1}$ ensures $\lambda_{i}\in[0,1]$.
 2.4 Worked Example
@@
 Allocation is 1:1 ($p=0.5$).
-Assume burn-in is complete ($c=1$).
+Assume burn-in is complete ($c_{i}=1$).
 We look back at patients 1–199.
@@
 We bet $\lambda=0.5+0.5\times(-0.054)=0.473$ on intervention and $1-\lambda=0.527$ on control.
-Assignment revealed and patient indeed is in control group.
-We guessed right!
-Our multiplier will be $0.527/0.5=1.054$, hence wealth grows 5.4%.
-Now patient 201 has been enrolled.
-First, we update the counts with patient 200’s data.
-Intervention arm now has 100 patients, 35 events (rate = 35.0%), and control arm has 100 patients, 41 events (rate = 41.0%).
-The $\hat{\delta}_{200}=0.350-0.410=-0.060$.
-Patient 201 does not have an event.
+Assignment revealed: control.
+We guessed right.
+Multiplier: $0.527/0.5=1.054$.
+Wealth grows 5.4%.
+Patient 201: Update counts—intervention has 100 patients, 35 events (35.0%); control has 100 patients, 41 events (41.0%).
+$\hat{\delta}_{200}=-0.060$.
+Patient 201 has no event.
 Non-events are more common in intervention (65.0% vs 59.0%), so probably intervention.
-We will bet once again.
-This time we bet $\lambda=0.5-0.5\times(-0.060)=0.530$ on intervention and $1-\lambda=0.470$ on control.
-Finally, assignment is revealed and we guessed right again: patient is in intervention.
-Our multiplier =0.530/0.5=1.060=0.530/0.5=1.060.
-Wealth grows another 6.0%.
-Finally, patient 202 has been enrolled.
-We update counts once more with intervention having 101 patients and 35 events (rate = 34.7%) and control having 100 patients with 41 events (rate = 41.0%).
-Our $\hat{\delta}_{201}=0.347-0.410=-0.063$.
-Patient 202 has an event (dies).
-Events cluster in control, so we bet toward control: $\lambda=0.5+0.5\times(-0.063)=0.469$ on intervention and $1-\lambda=0.531$ on control.
-Assignment is revealed and we guessed wrong—patient is in intervention group!
-Our multiplier will be $0.469/0.5=0.938$.
+Bet: $\lambda=0.5-0.5\times(-0.060)=0.530$.
+Assignment revealed: intervention.
+Multiplier: $0.530/0.5=1.060$.
+Wealth grows 6.0%.
+Patient 202: Update counts—intervention has 101 patients, 35 events (34.7%); control has 100 patients, 41 events (41.0%).
+$\hat{\delta}_{201}=-0.063$.
+Patient 202 has an event.
+Bet toward control: $\lambda=0.5+0.5\times(-0.063)=0.469$.
+Assignment revealed: intervention.
+Wrong guess.
+Multiplier: $0.469/0.5=0.938$.
 Wealth shrinks 6.2%.
-Cumulative wealth is the product of all multipliers:
+Cumulative wealth:
 $W_{202}=W_{199}\times 1.054\times 1.060\times 0.938=W_{199}\times 1.048$ (6)
 Despite one wrong guess, wealth grew 4.8% over these three patients.
-Even when intervention works, some intervention patients die.
-We guess wrong sometimes.
-But if intervention truly reduces mortality, events cluster in control and non-events cluster in intervention — we are right more often than wrong, and wealth grows on average.
-Under the null hypothesis (no effect), right and wrong guesses balance out. Wealth fluctuates randomly around 1.
-In summary: the outcome determines $\lambda$ (which direction to bet); the treatment arm determines which multiplier formula to apply (whether the bet paid off).
-The machinery is agnostic to whether the effect size is following the expected direction; it simply keeps betting according to what is more likely.
+Under the alternative, correct guesses outnumber incorrect ones on average and wealth grows.
+Under the null, right and wrong guesses balance out and wealth fluctuates around 1.
 2.5 Validity
 Theorem 1.
-Under the null hypothesis, the wealth process $(W_{n})$ is a test martingale — its expected value cannot grow.
+Under the null hypothesis, the wealth process $(W_{n})$ is a nonnegative martingale.
 Proof.
-Under the null, outcome and treatment are independent. So after seeing the outcome, treatment assignment is still just a coin flip with probability $p$. The expected wealth multiplier is:
-$\mathbb{E}[\text{multiplier}]$ $=p\times\frac{\lambda}{p}+(1-p)\times\frac{1-\lambda}{1-p}$ (6)
-$=\lambda+(1-\lambda)$ (7)
-$=1$ (8)
-No matter how we choose our bet $\lambda$, we can’t systematically grow wealth when the null is true. ∎
+Under the null, outcome and treatment are independent.
+After observing outcome $Y_{i}$, treatment assignment remains a coin flip with $P(T_{i}=1)=p$.
+The expected multiplier given any bet $\lambda_{i}$ is:
+$\mathbb{E}[\text{multiplier}\mid\lambda_{i}]$ $=p\times\frac{\lambda_{i}}{p}+(1-p)\times\frac{1-\lambda_{i}}{1-p}$ (7)
+$=\lambda_{i}+(1-\lambda_{i})=1$ (8)
+Thus $\mathbb{E}[W_{i}\mid W_{i-1},\lambda_{i}]=W_{i-1}$. ∎
 Corollary 1.
-For any stopping time $\tau$, $\mathbb{E}[W_{\tau}]\leq 1$ under the null.
-By Markov’s inequality, $P(W_{\tau}\geq 1/\alpha)\leq\alpha$.
-Thus rejecting when $W_{\tau}\geq 1/\alpha$ controls Type I error at level $\alpha$, regardless of the stopping rule.
+By Ville’s inequality (ville1939), for any nonnegative martingale starting at 1:
+$\Pr_{H_{0}}\left(\sup_{n\geq 1}W_{n}\geq\frac{1}{\alpha}\right)\leq\alpha$ (9)
+Thus rejecting when wealth ever exceeds $1/\alpha$ controls Type I error at level $\alpha$, regardless of when or why monitoring stops.
 3 Simulation Studies
-We evaluated e-RT operating characteristics by simulation.
+We evaluated operating characteristics by simulation.
 For each scenario, we calculated the sample size required for a chi-square test to achieve the target power at $\alpha=0.05$, then ran 5,000 simulated trials at that sample size.
@@
 Table 1 presents Type I error and power for trials designed to detect 5% or 10% absolute risk reductions (ARR) with 80% or 90% power.
-Table 1: e-RT Operating Characteristics
-ARR Target Power $n$ Type I Error e-RT Power Median Crossing
+Table 1: Operating Characteristics for Binary e-process
+ARR Target Power $n$ Type I Error Binary e-process Power Median Crossing
 5% 80% 2942 0.032 48.6% 1392 (47%)
@@
 Power was approximately 50% for trials designed with 80% power, and 63–66% for trials designed with 90% power.
-When the e-RT rejected the null, it did so at approximately half the planned sample size (median crossing 47–56% of total enrollment).
+When the process rejected the null, it did so at approximately half the planned sample size (median crossing 47–56% of total enrollment).
 3.2 Interpretation
-The e-RT is not a replacement for the primary analysis—it is a continuous monitoring tool.
-A trial designed for 80% power with a traditional test gains a 50% chance of stopping early via e-RT, at roughly the halfway point.
-If the threshold is not crossed, the trial proceeds to completion and the planned analysis is conducted.
+This method seems to be appealing as a continuous monitoring tool.
+A trial designed for 80% power with a traditional test gains a 50% chance of stopping early, at roughly the halfway point.
+If the threshold is not crossed, the trial proceeds to completion and the planned analysis is conducted with no penalty.
 This represents a free option: anytime-valid monitoring with no alpha spending and no pre-specified interim looks.
-The “cost” is that the e-RT alone has lower power than a fixed-sample test.
+The “cost” is that this method alone has lower power than a fixed-sample test.
 But when layered on top of a properly powered trial, it provides early stopping when effects are larger than anticipated.
 3.3 Trajectory Examples
-Figure 1 shows example e-RT trajectories from 30 simulated trials under the null hypothesis (both arms 40% event rate).
+Figure 1 shows example wealth trajectories from 30 simulated trials under the null hypothesis (both arms 40% event rate).
 Sample sizes correspond to trials designed for 80% power (n = 712, left) and 90% power (n = 954, right) to detect a 10% ARR.
@@
 The downward drift over time reflects the accumulating “cost” of betting on noise.
-Figure 1: e-RT trajectories under the null hypothesis.
+Figure 1: Wealth trajectories under the null hypothesis.
 Left: n = 712 (80% power design).
@@
 Most trajectories cross the rejection threshold before enrollment completes, and many reach values far exceeding 20 — providing strong evidence against the null.
-Figure 2: e-RT trajectories under the alternative hypothesis (true ARR = 10%).
-Panel A: n = 712 (80% power design).
-Panel B: n = 954 (90% power design).
+Figure 2: E-processes trajectories under the alternative hypothesis (true ARR = 10%).
+Left: n = 712 (80% power design).
+Right: n = 954 (90% power design).
 Dashed red line: rejection threshold ($1/\alpha=20$).
 Under the alternative, approximately half of trajectories cross the threshold, typically around the midpoint of enrollment.
+4 Continuous Outcomes
+The previous approach treats each patient as a single Bernoulli trial: the outcome (event vs.
+no event) is observed, and we bet on which arm that patient came from.
+For continuous endpoints, the logic is the same but the signal is richer.
+Each patient now contributes a continuous measurement (for example, ventilator-free days, change in biomarker, or a physiologic score), and the betting strategy uses how extreme that value is relative to past data.
+Therefore, defining wager is slightly more granular.
+One may extend this to continuous endpoint.
+The validity still comes only from randomization: under the null hypothesis, the distribution of the continuous outcome is the same in both arms, so the outcome does not help to predict treatment assignment.
+4.1 Setup
+At each enrollment $i=1,2,\ldots$, the trial generates two pieces of information.
+First, the randomization mechanism assigns the patient to one of the two arms, denoted by $T_{i}\in\{0,1\}$ (with 0 for control and $1$ for intervention), using a known allocation probability $p$ (typically $p=0.5$ for 1:1 randomization, although equal allocation is not required).
+Second, once follow-up is complete, we observe a continuous outcome $Y_{i}\in\mathbb{R}$ such as ventilator–free days, a biomarker concentration, or a physiologic measurement.
+So far, same as before but using a continuous endpoint.
+Under the null hypothesis of no treatment effect, the distribution of $Y_{i}$ is identical in both arms, and hence the outcome carries no information about the treatment assignment; formally, $Y_{i}\perp T_{i}$ under $H_{0}$.
+This single independence relationship is the foundation of the continuous randomization e-process.
+The idea mirrors the binary approach: each patient creates a small betting game in which we observe $Y_{i}$, form a data-driven guess about which arm the patient is more likely to belong to, and then update our wealth once the actual assignment $T_{i}$ is revealed.
+If the null is true, these guesses cannot systematically win because outcomes are uninformative about treatment.
+If the alternative is true, outcomes begin to separate between arms, the bets gain predictive power, and wealth grows accordingly.
+Viewed this way, extending the randomization e-process from binary to continuous outcomes may not require additional assumptions; it merely replaces the event/no-event signal with a continuous measure of extremeness relative to past observations, while preserving the anytime-valid martingale structure.
+The difference in how the wager is defined.
+4.2 Betting strategy for continuous outcomes
+We can use a data-driven rule with three ingredients:
+  1. 1.
+    A center and scale for the outcomes observed so far. Outcomes con be all all over the place. We need to standardize.
+  2. 2.
+    A standardized residual for the new outcome. As the reader will see we bet proportionally to how far the result is away from the center reference we chose. This also needs to be standardized.
+  3. 3.
+    A smooth map from that residual into a betting fraction $\lambda_{i}\in(0,1)$.
+If the measurement is an outlier, a huge wager would be placed because difference between measurement and the center of the scale would be huge, we need to muffle it to something between $(0,1)$.
+In simpler terms: Each new outcome is first compared to the past outcomes to understand how “unusual” it is.
+To do this, we anchor the outcome to a robust center and a robust scale: the median gives the center and the MAD (median absolute deviation) gives the spread.
+We then compute a standardized value: how far above or below the median this new observation is, measured in MAD units.
+That standardized number is what guides how aggressively we bet.
+A large positive value means “this looks unusually high compared with past outcomes,” a large negative value means the opposite, and values near zero mean “this one looks typical.” The point of the MAD is simple: it behaves well even when early data are messy or skewed.
+It stops a single extreme value from blowing up the bet and keeps the e-process stable while the trial is still young.
+This likely comes at the cost of reduced power.
+At step $i$, using all previous outcomes $Y_{1},\dots,Y_{i-1}$, we compute:
+$m_{i-1}$ $=\text{median}(Y_{1},\dots,Y_{i-1}),$ (10)
+$s_{i-1}$ $=\text{MAD}(Y_{1},\dots,Y_{i-1}),$ (11)
+where MAD is the median absolute deviation.
+MAD is defined as $\mathrm{MAD}=\operatorname{median}(|Y_{i}-\operatorname{median}(Y)|)$.
+These are robust to outliers and skewness.
+If $s_{i-1}$ is zero or not finite, we set $s_{i-1}=1$ to avoid degeneracy.
+For the new patient, we form a standardized residual:
+$r_{i}=\frac{Y_{i}-m_{i-1}}{s_{i-1}}.$ (12)
+This means that the patient residual $r_{i}$ is the observed value ($Y_{i}$) minus the median observed so far ($s_{i-1}$ divided by the MAD ($s_{i-1}$).
+Note how this uses information for patients before $i$, keeping the martingale.
+We then squash this standardized value into the interval (1,1)(-1,1) using
+$g_{i}=\frac{r_{i}}{1+|r_{i}|}.$
+This is a simple monotone transformation: for moderate values $g_{i}\approx r_{i}$, while very large positive or negative residuals are shrunk toward +1+1 or 1-1.
+The only purpose of this step is to prevent a single extreme observation from forcing an almost all-in bet.
+Next, we ramp up the betting strength over time. Let
+$c_{i}=\min\left\{1,\max\left(0,\frac{i-\text{burn-in}}{\text{ramp}}\right)\right\},$ (13)
+where burn-in is the number of initial patients during which we essentially do not bet, and ramp controls how quickly we move from very cautious betting to our maximum aggressiveness.
+Those concepts are exactly like the binary approach.
+Finally, we cap the maximum betting strength at $c_{\max}\in(0,0.5]$ to avoid pathological bets.
+The betting fraction $\lambda_{i}$ is then
+$\lambda_{i}=0.5+c_{i}c_{\max}g_{i}.$ (14)
+By construction, $\lambda_{i}\in(0,1)$ and is predictable: it depends only on past outcomes and the new $Y_{i}$, not on $T_{i}$.
+Intuitively:
+  • If $Y_{i}$ is close to the historical median, $g_{i}\approx 0$ and $\lambda_{i}\approx 0.5$: we essentially do not bet.
+  • If $Y_{i}$ is unusually low (for example, very high ventilator-free days in a trial where we expect intervention to improve outcomes), then $g_{i}<0$ or $g_{i}>0$ depending on how we code the outcome and the expected direction of benefit.
+The sign of $g_{i}$ determines which arm we lean towards.
+  • Early in the trial, $c_{i}$ is small, so even unusual observations lead to mild bets. As data accumulates, $c_{i}$ approaches 1 and the bets become more confident.
+4.3 Wealth update
+Treatment is still randomized with probability $p$ of intervention. After we choose $\lambda_{i}$, the wealth updates exactly as in the binary approach:
+$W_{i}=W_{i-1}\times\begin{cases}\lambda_{i}/p&\text{if }T_{i}=1\ (\text{intervention})\\ (1-\lambda_{i})/(1-p)&\text{if }T_{i}=0\ (\text{control})\end{cases}$ (15)
+with $W_{0}=1$.
+The only difference from the binary case is how we choose $\lambda_{i}$, as we saw.
+For binary outcomes, $\lambda_{i}$ depends on the event indicator and past event rates by arm.
+For continuous outcomes, $\lambda_{i}$ depends on how extreme $Y_{i}$ is relative to past outcomes.
+4.4 Worked intuition
+Imagine a trial where the outcome is ventilator-free days, and higher is better.
+Suppose that after 100 patients, the median and MAD of $Y$ are roughly stable.
+Patient 101 has an unusually high number of ventilator-free days compared with this distribution.
+The standardized residual $r_{101}$ is positive and large, so $g_{101}\approx 0.8$ and, after burn-in, $c_{101}\approx 1$.
+With $c_{\max}=0.6$, we might get $\lambda_{101}\approx 0.5+0.6\times 0.8\approx 0.98$: we strongly bet that this patient was in the intervention arm.
+If they indeed were, wealth increases by roughly a factor of $0.98/0.5\approx 2$ for this one patient.
+If not, wealth shrinks by $(1-0.98)/0.5\approx 0.04$.
+Under the null, high values like this are just as likely in control as in intervention: we win and lose in balance, and wealth does not grow on average.
+Under a true benefit, such favorable outliers cluster in the intervention arm, and the bets pay off more often than not.
+4.5 Validity
+The key point is that validity does not depend on the choice of median, MAD, or the specific transformation $g_{i}$. It depends only on the fact that:
+  1. 1.
+    $\lambda_{i}$ is chosen before observing $T_{i}$ and depends only on past data and $Y_{i}$;
+  2. 2.
+    under the null, $T_{i}$ is independent of $Y_{i}$ with $\mathbb{P}(T_{i}=1)=p$.
+Theorem 2.
+Under the null hypothesis of no treatment effect, the e-process wealth process $(W_{i})$ is a test martingale: for all $i$,
+$\mathbb{E}[W_{i}\mid\mathcal{F}_{i-1}]\leq W_{i-1},$
+where $\mathcal{F}_{i-1}$ is the sigma-field generated by all observations up to step i1i-1.
+Proof.
+Condition on $\mathcal{F}_{i-1}$ and $Y_{i}$. The bet $\lambda_{i}$ is now fixed. Under the null, $T_{i}$ is independent of $Y_{i}$ and
+$\mathbb{P}(T_{i}=1\mid\mathcal{F}_{i-1},Y_{i})=p,\quad\mathbb{P}(T_{i}=0\mid\mathcal{F}_{i-1},Y_{i})=1-p.$
+The conditional expectation of the wealth multiplier is:
+$\mathbb{E}\!\left[\frac{W_{i}}{W_{i-1}}\,\middle|\,\mathcal{F}_{i-1},Y_{i}\right]$ $=p\cdot\frac{\lambda_{i}}{p}+(1-p)\cdot\frac{1-\lambda_{i}}{1-p}$ (16)
+$=\lambda_{i}+(1-\lambda_{i})$ (17)
+$=1.$ (18)
+Thus $\mathbb{E}[W_{i}\mid\mathcal{F}_{i-1},Y_{i}]=W_{i-1}$, and taking expectations over $Y_{i}$ yields $\mathbb{E}[W_{i}\mid\mathcal{F}_{i-1}]=W_{i-1}$.
+This shows that $(W_{i})$ is a martingale with unit expectation under the null.
+∎
+As in the binary case, Ville’s inequality implies that for any stopping time $\tau$,
+$\mathbb{P}(W_{\tau}\geq 1/\alpha)\leq\alpha,$
+so rejecting the null when $W_{\tau}\geq 1/\alpha$ controls Type I error at level $\alpha$, regardless of the stopping rule.
+4.6 Simulation overview
+We evaluated this using the same design philosophy as for binary approach.
+For a given standardized effect size (Cohen’s $d$) and target power, we first computed the fixed-sample size required for a two-sample $t$-test at $\alpha=0.05$.
+We then simulated trials with that sample size, assigning patients 1:1 to intervention or control, with outcomes drawn from normal distributions of equal variance and means differing by $d$ under the alternative.
+The test was run sequentially with a burn-in period and ramp (for example, burn-in = 20 patients, ramp = 50 patients, $c_{\max}=0.6$).
+Under the null (no mean difference between arms), Type I error was close to or below the nominal level.
+Under the alternative, the implementation for continuous endpoints rejected the null with moderate power and typically crossed the threshold at an intermediate sample size (Table 2).
+Again, we believe that this may not be a replacement for the fixed-sample $t$-test but a conservative, anytime-valid monitoring tool that can trigger early stopping when effects are larger or clearer than anticipated.
+Table 2: Operating Characteristics for Continuous Outcomes
+Cohen’s $d$ Target Power $n$ Type I Error e-process Power Median Crossing
+0.20 80% 788 0.042 14.1% 66
+0.40 80% 200 0.038 33.6% 66
+0.60 80% 90 0.037 55.1% 63
+0.20 90% 1054 0.043 14.1% 67
+0.40 90% 266 0.043 34.8% 66
+0.60 90% 120 0.045 58.5% 65
+A visual representation can be found in figure below.
+Figure 3: Trajectories of the continuous randomization e-process for a trial designed to detect a standardized mean difference of $d=0.40$ with 80% power.
+Left: trajectories under the null hypothesis ($d=0$), where wealth wanders near or below 1 and rarely approaches the rejection threshold ($1/\alpha=20$).
+Right: trajectories under the alternative hypothesis ($d=0.40$), where wealth grows systematically and many paths cross the rejection threshold before the planned sample size is reached.
+Figure 3 illustrates 30 simulated trajectories of the process for continuous endpoints under a design targeting a standardized effect size of $d=0.40$ with 80% power.
+Under the null hypothesis (left panel), wealth fluctuates around 1 and gradually drifts downward as repeated small bets accumulate against noise.
+No trajectory crosses the rejection threshold of $1/\alpha=20$.
+Under the corresponding alternative (right panel), outcomes begin to separate between arms, the bets become systematically correct, and the wealth grows.
+Some trajectories cross the rejection threshold before the planned sample size, demonstrating the potential for early stopping.
+5 Time-to-Event Outcomes
+Clinical trials often use time-to-event endpoints (e.g., overall survival), usually analyzed via the Log-Rank test or Cox proportional hazards models.
+These traditional methods require assumptions about proportional hazards or require waiting for a specific number of events.
+One can extend the randomization e-process to survival data, constructing a sequential Log-Rank test that updates wealth at every observed event.
+pmlr-v146-grunwald21a developed a safe logrank test using e-values under a proportional hazards model with a prior on the hazard ratio.
+We attempted to construct a nonparametric approach where validity derives solely from randomization, not from a correctly specified hazard model.
+5.1 Setup and Martingale Construction
+Let $N$ patients be randomized to treatment ($T=1$) or control ($T=0$).
+We observe outcomes over time.
+The "time" scale here is the distinct ordering of events.
+Let $t_{1}<t_{2}<\dots<t_{k}$ denote the times at which events occur.
+At any event time $t_{j}$, we define the risk set $\mathcal{R}_{j}$ as the set of patients who have not yet had an event and have not been censored.
+Let $Y_{1}(t_{j})$ and $Y_{0}(t_{j})$ be the number of patients at risk in the treatment and control arms, respectively.
+Under the null hypothesis of no treatment effect, the probability that the event at time $t_{j}$ comes from the treatment arm, conditional on a failure occurring within $\mathcal{R}_{j}$, is simply the proportion of treated patients at risk:
+$p_{j}=\frac{Y_{1}(t_{j})}{Y_{1}(t_{j})+Y_{0}(t_{j})}.$ (19)
+Let $X_{j}$ be the indicator that the event at $t_{j}$ is a treated patient ($X_{j}=1$ if treated, 0 if control).
+Under the null, $X_{j}$ is a Bernoulli trial with probability $p_{j}$.
+We construct the martingale increment (score) as:
+$U_{j}=X_{j}-p_{j}.$ (20)
+Note that $\mathbb{E}[U_{j}|\mathcal{R}_{j}]=0$.
+5.2 Betting Strategy
+We wager on the sign of $U_{j}$.
+If the treatment is beneficial, events will occur more slowly in the treatment arm than expected under the null.
+Thus, the observed number of treatment events will be lower than the expected number, leading to a negative trend in the cumulative sum of $U_{j}$.
+We define the cumulative Log-Rank score at step j1j-1 as $Z_{j-1}=\sum_{k=1}^{j-1}U_{k}$. Our betting strategy targets this trend:
+$\lambda_{j}=\text{sign}(Z_{j-1})\cdot c_{j}\cdot\lambda_{\max},$ (21)
+where $c_{j}\in[0,1]$ is a ramping function similar to previous sections, and $\lambda_{\max}<1$ is a cap on betting aggressiveness.
+The wealth update at event $j$ is:
+$W_{j}=W_{j-1}\times(1+\lambda_{j}U_{j}).$ (22)
+Because $\mathbb{E}[U_{j}]=0$, the expected multiplicative factor is 1 under the null. Thus, $(W_{j})$ is a test martingale.
+Note that unlike the binary and continuous approaches, this betting strategy uses only the sign of the cumulative score $Z_{j-1}$, not its magnitude.
+Once evidence favors one direction, the bet size is fixed at $\lambda_{\max}$ regardless of how strong the accumulated evidence is.
+This mirrors the Kelly criterion in betting theory: the optimal wager size depends on the expected edge, and $\lambda_{\max}=0.25$ is calibrated for moderate effects (HR $\approx 0.80$(kelly1956).
+The connection between the wager, $\lambda$, and Kelly’s ideas on fraction of betting needs to be further explored.
+In brief, it makes sense that wager should be higher when prospects of winning are more favorable.
+For time-to-event, however, adapting response to events may take a long time, and a fixed wager may be preferable.
+The parameters burn-in, ramp, and $\lambda_{\max}$ are set arbitrarily (burn-in =30=30, ramp =50=50, and $\lambda_{\max}=0.25$ in simulations).
+Different choices will yield different operating characteristics.
+The validity of the test does not depend on these choices—only efficiency does.
+5.3 Handling Staggered Entry
+In clinical practice, patients are recruited over time (staggered entry), whereas this simplified simulation generates survival times simultaneously.
+However, the Log-Rank test and this betting strategy rely solely on the rank ordering of events based on “time on study.”
+We verified the validity of this simplification by simulating two scenarios with $N=631$ and a true Hazard Ratio of 0.80: (1) simultaneous entry, and (2) staggered entry where patients were recruited uniformly over 12 months, and analysis was performed on calculated study time ($T_{\text{event}}-T_{\text{entry}}$).
+The resulting distributions of final e-values were similar (Simultaneous Median $E\approx 25.1$, Staggered Median $E\approx 23.6$; Power $\approx 54\%$ and $53\%$ respectively).
+This confirms that the sequential e-process remains valid for real-world staggered designs provided the analysis utilizes time-since-randomization.
+5.4 Simulation Results
+We simulated survival trials comparing exponential survival times with a Hazard Ratio (HR) of 0.80.
+Targeted power was 80% using a standard Log-Rank design, which requires approximately 631 events.
+A total of $N=631$ patients (assuming no censoring) were used for simulations.
+Results are shown in Table 3.
+Table 3: Operating Characteristics (N=631)
+True HR Target HR Type I Error Power Median Events to Stop
+1.00 (Null) 0.80 0.039
+0.80 (Alt) 0.80 62.8% 329 (52% of N)
+Under the alternative (HR=0.80), the e-survival process achieved 62.8% power to reject the null, with a median stopping time of 329 events. Examples are shown in Figure 4.
+Figure 4: Trajectories of the e-Survival process for a trial designed to detect a Hazard Ratio of 0.80 with 80% power ($N=631$).
+Left: trajectories under the null hypothesis (HR = 1.00), where wealth fluctuates randomly.
+Right: trajectories under the alternative hypothesis (HR = 0.80), where wealth grows systematically.
+The red dashed line represents the rejection threshold ($1/\alpha=20$).
 6 Discussion
-The e-RT is a nonparametric sequential test for randomized trials based on the betting framework for e-values.
+The binary implementation is a nonparametric sequential test for randomized trials based on the betting framework for e-values (i-bet pmlr-v177-duan22a).
 The method requires only that treatment assignment is randomized — no distributional assumptions about outcomes are needed.
 This makes it a robust complement to model-based analyses.
+It can be extended to continuous endpoints and time-to-event analyses.
 6.1 Operating characteristics
-Simulations demonstrate that e-RT controls Type I error conservatively (2–3% vs nominal 5%) while providing approximately 50% power for early stopping in trials designed with 80% power, and 63–66% power in trials designed with 90% power.
+Simulations demonstrate that the methos properly controlled for Type I error (2–3% vs nominal 5%) while providing approximately 50% power for early stopping in trials designed with 80% power, and 63–66% power in trials designed with 90% power.
 When early stopping occurs, it happens at roughly half the planned sample size.
+Both the continuous and the survival implementations had similar characteristics: Power was traded for continuous Type 1 error control.
+All simulations showed that Type 1 error was exceptionally well controlled but power was nominally low.
 These results should be interpreted carefully.
-The e-RT is not designed to replace the primary analysis of a trial; at least not yet.
-Rather, it provides a continuous monitoring option that requires no alpha spending and no prespecified interim analysis schedule.
-If the e-RT crosses its threshold, one could consider stopping the trial.
+From a clinical trial perspective, it is uncertain whether those methods can replace traditional frequentist or Bayesian paradigms, but they may provide a continuous monitoring option that requires no alpha spending and no prespecified interim analysis schedule.
+If the e-process crosses its threshold, one could consider stopping the trial, for example.
 If it does not cross, the trial may proceed to its planned conclusion and primary analysis.
+6.2 What is the null hypothesis being tested?
+This approach tests whether treatment assignment can be predicted from outcomes—equivalently, whether outcomes are exchangeable between arms.
+Under the null, $Y_{i}\perp T_{i}$ at each observation: knowing the outcome provides no information about which arm the patient belongs to.
+This is neither Fisher’s sharp null (every individual has exactly zero treatment effect) nor the weak null of equal population means.
+A useful analogy is a casino.
+Under the null, the house is fair: no betting strategy can systematically grow wealth.
+The e-value quantifies accumulated evidence that the game is beatable.
+Rejecting the null means we have found a profitable strategy—outcomes predict assignments better than chance.
+The user is also referenced to the pioneer lessons by ramdas2021game.
+This framing clarifies both the method’s strength and its limitation.
+The strength is generality: any departure from exchangeability—constant effects, heterogeneous effects, or even randomization failures—makes outcomes informative and wealth grows.
+The limitation is that the test detects only departures that a cumulative backward-looking betting strategy can exploit.
+In practical terms, the null hypothesis is best understood as: “there exists no stable, learnable relationship between outcomes and treatment assignment that persists long enough for an adaptive strategy to exploit.” This is appropriate for most trials where effects are expected to be consistent over enrollment, but investigators should be aware that non-stationary effects represent a blind spot.
 6.3 Relationship to existing work
@@
 E-values and e-processes have been extensively studied (vovk2021; ramdas2022; ramdas2025).
+pmlr-v177-duan22a introduced interactive rank testing by betting (i-bet), which applies the betting framework directly to randomized experiments: an analyst sequentially bets on treatment assignments based on observed outcomes, with wealth forming a test martingale under the null.
+The binary approach implements this framework with a specific adaptive betting strategy tied to outcome values.
+Betting approaches have been established for estimating means of bounded random variables (waudbysmith2023estimating).
+The continuous extension adapts these principles to the two-sample randomization setting using a standardization strategy.
 koning2025 develops e-values for group invariance, including permutation tests, using batch-based likelihood ratio statistics normalized by permutation expectations.
-Our construction differs in using patient-by-patient betting with adaptive learning of the effect size.
-These approaches share theoretical foundations but yield different procedures.
+pmlr-v146-grunwald21a developed the ‘Safe Log-rank Test’ based on evaluating likelihood ratios with specific priors on the hazard ratio to ensure growth rate optimality.
+In contrast, the survival approach constructs a linear test martingale directly from the score function using an adaptive betting strategy.
+Rather than relying on likelihood integration or specific priors, the survival approach process uses a heuristic ‘plug-in’ estimate of the effect direction, modulated by a ramping function.
+This offers a computationally simple, algorithmic alternative that derives validity strictly from the randomization probabilities within the risk set, without requiring the full apparatus of partial likelihood theory.
+It is possible that some of the concepts here were discussed by other authors in different contexts that were not immediately available for this author.
+Reader is encouraged to reach out if that is the case, and the author will happily adjust accordingly.
 6.4 Limitations
@@
 This is an experimental method under development.
-So far, this has only been tested for binary outcomes.
-It is reasonable to assume that extending to continuous endpoints would be feasible; this is planned.
-It is uncertain how this method could be applied for time-to-event endpoints.
-Hypothetically, we could hold the bet value for all patients randomized in a period until outcomes arrive, and approaches such as restricted mean survival time could be used.
-This is also under development.
-There are other limitations.
-The e-RT tests only whether event rates differ between arms.
-It does not directly estimate treatment effects or provide confidence intervals.
+Simulations are not exhaustive and were done for binary, continuous endpoints, and time to event.
+It is uncertain how this method would behave in more complex models, including competing risk models.
+The methods proposed test only whether there are differences between arms.
+They do not directly estimate treatment effects or provide confidence intervals.
 The adaptive learning of $\hat{\delta}$ requires a burn-in period during which little evidence accumulates.
 Third, for trials where parametric assumptions are plausible, model-based sequential methods will generally have better power.
-Fourth, our simulations used a specific betting strategy (burn-in = 50, ramp = 100); other choices may yield different operating characteristics.
+Fourth, our simulations used specific betting strategies; other choices may yield different operating characteristics.
 Hypothetically, it could be possible to increase betting aggressiveness over time as evidence accumulates; this may result in better power.
@@
 This is also under development.
+6.5 Future Directions
+Several extensions merit exploration.
+First, the current betting strategy uses a cumulative estimate $\hat{\delta}$ that weights all historical observations equally.
+This makes the method vulnerable to time-varying effects: if treatment benefit reverses to harm mid-trial, the strategy continues betting on stale information and wealth erodes despite continuous violation of exchangeability.
+Adaptive weighting schemes—such as exponential decay, rolling windows, or hybrid estimators blending long-term and recent signals—could improve robustness to non-stationary effects.
+Second, the betting intensity could itself adapt to recent performance: increasing $\lambda$ during sustained wealth growth (exploiting a confirmed edge) and dampening it following drawdowns (protecting against regime change).
+This mirrors Kelly criterion extensions that incorporate drawdown constraints.
+These refinements trade power under stable effects against robustness to drift, and the optimal balance likely depends on the clinical context.
+This is under development.
 6.6 Conclusion
-The e-RT provides anytime-valid sequential inference for randomized trials using only the guarantee of randomization.
-Its validity is unconditional on the data-generating process, making it a robust tool for trial monitoring.
+E-processes provide conceptually valid anytime-valid sequential inference for randomized trials using only the guarantee of randomization.
+Its validity is unconditional on the data-generating process, making it a potentially useful tool for trial monitoring.
 While it trades power for this robustness, it offers a valuable complement to traditional analysis methods.
+7 Disclaimers and Version Control
 7.1 Disclaimer
 This is an experimental method under development.
-Application to real patients should only be considered under surveillance from an experienced statistician.
+Application to real patients should only be considered under surveillance from an experienced statistician and remain strongly discouraged at this point by the author.
 The author is not responsible for consequences of use of this method.
+7.2 LLM use statement
+Large language models were extensively used in this work.
+The author had the idea that perhaps the e-value and e-process machinery could be used to bet against randomization which would result in a continuous trial monitoring tool.
+They uploaded the references in this manuscript to Gemini 3.0 Pro for brainstorming, which quickly resulted in a preliminary version.
+This was refined, tested, and debugged using Claude 4.5 Opus and ChatGPT 5.1 Pro.
+Gemini 3.0 Pro aided with coding for survival approach.
+Claude 4.5 Opus provided final spellchecking and organization of this draft, which remains suboptimal despite the efforts of several LLMs.
+7.3 Acknowledgments
+The author is thankful to Aaditya Ramdas for their thoughtful comments on the first version and for pointing out previous literature to the author.
 7.4 Version Control
@@
   2. 2.
-    Second Version (this version): Minor text adjustments; removed claim on sharp null.
+    Second Version (Dec 07, 2025): Minor text adjustments; removed claim on sharp null.
+  3. 3.
+    Third Version (this version): Added continuous and survival endpoints; text adjustments.
 Appendix A R Code
@@
 ggsave("traj_null_10pct_90pow.pdf", p4, width = 8, height = 5)
+# e-RTC (Randomization E-Process for Continuous Outcomes)
+# Type I Error and Power Simulations (Unified)
+setwd("")
+suppressPackageStartupMessages({
+library(tidyverse)
+# Core e-RTC function
+# treatment: 0/1 vector (0 = control, 1 = intervention)
+# outcome : continuous outcome (numeric)
+# p : allocation prob for treatment (default 0.5)
+# burn_in : number of patients before bets start ramping up
+# ramp : additional patients over which betting strength ramps to max
+# c_max : maximum betting strength (0 < c_max <= 0.5 recommended)
+compute_eRTC <- function(treatment, outcome,
+p = 0.5,
+burn_in = 20,
+ramp = 50,
+c_max = 0.6) {
+n <- length(treatment)
+if (is.factor(treatment)) treatment <- as.numeric(treatment) - 1
+wealth <- numeric(n)
+wealth[1] <- 1
+for (i in 2:n) {
+# Past outcomes
+out_prev <- outcome[1:(i-1)]
+# Require enough past data before betting
+if ((i - 1) < burn_in) {
+wealth[i] <- wealth[i-1]
+next
+# Robust center and scale: median + MAD
+med_prev <- median(out_prev, na.rm = TRUE)
+mad_prev <- mad(out_prev, center = med_prev, constant = 1, na.rm = TRUE)
+if (!is.finite(mad_prev) $|$ mad_prev <= 0) mad_prev <- 1
+# Current outcome
+y_i <- outcome[i]
+# Standardized residual, squashed to (-1, 1)
+r_i <- y_i - med_prev
+s_i <- r_i / mad_prev
+g_i <- s_i / (1 + abs(s_i)) # in (-1,1), robust to outliers
+# Ramping betting strength
+ramp_frac <- max(0, min(1, (i - burn_in) / ramp))
+c_i <- ramp_frac * c_max
+lambda <- 0.5 + c_i * g_i
+lambda <- max(0.001, min(0.999, lambda))
+# Wealth update based on actual assignment
+if (treatment[i] == 1) {
+multiplier <- lambda / p
+} else {
+multiplier <- (1 - lambda) / (1 - p)
+wealth[i] <- wealth[i-1] * multiplier
+return(wealth)
+# Simulate a single trial with continuous outcomes
+# n : total number of patients
+# p_trt : probability of being assigned to treatment (default 0.5)
+# mu_ctrl : mean in control arm
+# mu_trt : mean in treatment arm
+# sd : common standard deviation (default 1)
+simulate_trial_continuous <- function(n,
+p_trt = 0.5,
+mu_ctrl = 0,
+mu_trt = 0,
+sd = 1) {
+treatment <- rbinom(n, 1, p_trt)
+outcome <- numeric(n)
+outcome[treatment == 0] <- rnorm(sum(treatment == 0), mean = mu_ctrl, sd = sd)
+outcome[treatment == 1] <- rnorm(sum(treatment == 1), mean = mu_trt, sd = sd)
+return(data.frame(treatment = treatment, outcome = outcome))
+# Unified simulation function for e-RTC
+# n_sims : number of simulated trials
+# mu_ctrl : control mean (default 0)
+# mu_trt : treatment mean (NULL => null, mu_trt = mu_ctrl)
+# hypothesized_d: Cohen’s d used for sample size calc (delta/sd)
+# target_power : design power (0.80, 0.90)
+# alpha : significance level
+# burn_in, ramp, c_max: e-RTC tuning parameters
+simulate_eRTC <- function(n_sims = 5000,
+mu_ctrl = 0,
+mu_trt = NULL,
+hypothesized_d,
+target_power = 0.80,
+alpha = 0.05,
+burn_in = 20,
+ramp = 50,
+c_max = 0.6) {
+if (missing(hypothesized_d)) {
+stop("Mustspecifyhypothesized_d(Cohen’sdfordesign).")
+# Treat SD = 1 for design (Cohen’s d = delta / sd)
+sd_true <- 1
+delta_design <- hypothesized_d * sd_true
+# Sample size via t-test (two-sample, equal sizes)
+$s$ <- power.t.test(delta = delta_design,
+sd = sd_true,
+power = target_power,
+sig.level = alpha,
+type = "two.sample",
+alternative = "two.sided")
+n_per_arm <- ceiling(ss\$n)
+n_patients <- 2 * n_per_arm
+# Determine true mu_trt (null if not specified)
+if (is.null(mu_trt)) {
+mu_trt <- mu_ctrl # Null is true
+true_d <- 0
+sim_type <- "TypeIError"
+} else {
+true_d <- (mu_trt - mu_ctrl) / sd_true
+sim_type <- "Power"
+[…]
+[…]
+mu_ctrl, hypothesized_d, target_power * 100))
+[…]
+[…]
+mu_ctrl, mu_trt, true_d))
+[…]
+rejections <- 0
+first_crossing <- numeric(n_sims)
+final_evalues <- numeric(n_sims)
+pb <- txtProgressBar(min = 0, max = n_sims, style = 3)
+for (sim in 1:n_sims) {
+trial <- simulate_trial_continuous(n_patients,
+mu_ctrl = mu_ctrl,
+mu_trt = mu_trt,
+sd = sd_true)
+wealth <- compute_eRTC(trial\$treatment,
+trial\$outcome,
+p = 0.5,
+burn_in = burn_in,
+ramp = ramp,
+c_max = c_max)
+final_evalues[sim] <- wealth[n_patients]
+crossing <- which(wealth >= 1 / alpha)
+if (length(crossing) > 0) {
+rejections <- rejections + 1
+first_crossing[sim] <- crossing[1]
+} else {
+first_crossing[sim] <- NA
+setTxtProgressBar(pb, sim)
+close(pb)
+rejection_rate <- rejections / n_sims
+se <- sqrt(rejection_rate * (1 - rejection_rate) / n_sims)
+median_stop <- median(first_crossing, na.rm = TRUE)
+[…]
+[…]
+[…]
+rejection_rate - 1.96 * se,
+rejection_rate + 1.96 * se))
+if (sim_type == "Power") {
+[…]
+[…]
+} else {
+[…]
+return(list(
+sim_type = sim_type,
+rejection_rate = rejection_rate,
+se = se,
+n_per_arm = n_per_arm,
+n_patients = n_patients,
+mu_ctrl = mu_ctrl,
+mu_trt = mu_trt,
+hypothesized_d = hypothesized_d,
+true_d = true_d,
+target_power = target_power,
+alpha = alpha,
+median_crossing = median_stop,
+first_crossing = first_crossing,
+final_evalues = final_evalues
+# Plot trajectories under null / alternative
+# n_trials : number of trajectories to overlay
+# mu_ctrl, mu_trt, hypothesized_d, target_power, alpha, burn_in, ramp, c_max
+plot_trajectories_eRTC <- function(n_trials = 30,
+mu_ctrl = 0,
+mu_trt = NULL,
+hypothesized_d,
+target_power = 0.80,
+alpha = 0.05,
+burn_in = 20,
+ramp = 50,
+c_max = 0.6,
+title = NULL) {
+if (missing(hypothesized_d)) {
+stop("Mustspecifyhypothesized_d(Cohen’sd).")
+sd_true <- 1
+delta_design <- hypothesized_d * sd_true
+$s$ <- power.t.test(delta = delta_design,
+sd = sd_true,
+power = target_power,
+sig.level = alpha,
+type = "two.sample",
+alternative = "two.sided")
+n_patients <- 2 * ceiling(ss\$n)
+if (is.null(mu_trt)) {
+mu_trt <- mu_ctrl # Null
+true_d <- 0
+} else {
+true_d <- (mu_trt - mu_ctrl) / sd_true
+if (is.null(title)) {
+if (true_d == 0) {
+title <- sprintf(
+"e-RTCUnderNull(n=%d,designedford=%.2f)",
+n_patients, hypothesized_d
+} else {
+title <- sprintf(
+"e-RTCUnderAlternative(n=%d,trued=%.2f,%.0f%%powerdesign)",
+n_patients, true_d, target_power * 100
+trajectories <- vector("list", n_trials)
+for (i in 1:n_trials) {
+trial <- simulate_trial_continuous(n_patients,
+mu_ctrl = mu_ctrl,
+mu_trt = mu_trt,
+sd = sd_true)
+wealth <- compute_eRTC(trial\$treatment,
+trial\$outcome,
+p = 0.5,
+burn_in = burn_in,
+ramp = ramp,
+c_max = c_max)
+trajectories[[i]] <- data.frame(
+patient = 1:n_patients,
+wealth = wealth,
+trial = i
+df <- bind_rows(trajectories)
+p <- ggplot(df, aes(x = patient, y = wealth, group = trial)) +
+geom_line(alpha = 0.4, color = "steelblue") +
+geom_hline(yintercept = 1 / alpha, linetype = "dashed", color = "red") +
+geom_hline(yintercept = 1, linetype = "dotted", color = "gray50") +
+scale_y_log10() +
+labs(
+title = title,
+subtitle = sprintf("mu_ctrl=%.2f,mu_trt=%.2f,trued=%.2f",
+mu_ctrl, mu_trt, true_d),
+x = "Patient",
+y = "e-value(logscale)"
+annotate("text",
+x = n_patients * 0.95,
+y = (1 / alpha) * 1.5,
+label = sprintf("1/alpha=%.0f", 1 / alpha),
+color = "red", hjust = 1) +
+theme_minimal() +
+theme(
+plot.title = element_text(face = "bold"),
+panel.grid.minor = element_blank()
+return(p)
+# Run all scenarios (grid over d and power)
+# ds : vector of Cohen’s d (e.g. c(0.2, 0.4, 0.6))
+# target_powers: vector of target powers (e.g. c(0.80, 0.90))
+run_all_eRTC <- function(n_sims = 5000,
+mu_ctrl = 0,
+ds = c(0.2, 0.4, 0.6),
+target_powers = c(0.80, 0.90),
+alpha = 0.05,
+burn_in = 20,
+ramp = 50,
+c_max = 0.6,
+seed = 42) {
+set.seed(seed)
+results <- list()
+scenarios <- expand.grid(
+hypothesized_d = ds,
+target_power = target_powers,
+stringsAsFactors = FALSE
+all_results <- data.frame()
+for (i in 1:nrow(scenarios)) {
+hyp_d <- scenarios\$hypothesized_d[i]
+tgt_pow <- scenarios\$target_power[i]
+[…]
+[…]
+hyp_d, tgt_pow * 100))
+[…]
+# Type I error
+[…]
+t1 <- simulate_eRTC(
+n_sims = n_sims,
+mu_ctrl = mu_ctrl,
+mu_trt = NULL, # Null true
+hypothesized_d = hyp_d,
+target_power = tgt_pow,
+alpha = alpha,
+burn_in = burn_in,
+ramp = ramp,
+c_max = c_max
+# Power
+[…]
+pow <- simulate_eRTC(
+n_sims = n_sims,
+mu_ctrl = mu_ctrl,
+mu_trt = mu_ctrl + hyp_d, # sd = 1, so delta = d
+hypothesized_d = hyp_d,
+target_power = tgt_pow,
+alpha = alpha,
+burn_in = burn_in,
+ramp = ramp,
+c_max = c_max
+all_results <- rbind(all_results, data.frame(
+hypothesized_d = hyp_d,
+target_power = tgt_pow,
+n_patients = t1\$n_patients,
+type1_error = t1\$rejection_rate,
+type1_se = t1\$se,
+eRTC_power = pow\$rejection_rate,
+power_se = pow\$se,
+median_crossing = pow\$median_crossing
+results\$summary <- all_results
+[…]
+[…]
+[…]
+print(all_results %>%
+mutate(
+hypothesized_d = sprintf("%.2f", hypothesized_d),
+target_power = sprintf("%.0f%%", target_power * 100),
+type1_error = sprintf("%.3f", type1_error),
+eRTC_power = sprintf("%.1f%%", eRTC_power * 100)
+select(hypothesized_d, target_power, n_patients,
+type1_error, eRTC_power, median_crossing))
+return(results)
+# Example run (interactive)
+if (interactive()) {
+# Grid of d and power
+results_eRTC <- run_all_eRTC(
+n_sims = 2000, # lower for quick tests; bump to 5000+ for final
+mu_ctrl = 0,
+ds = c(0.2, 0.4, 0.6),
+target_powers = c(0.80, 0.90),
+alpha = 0.05,
+burn_in = 20,
+ramp = 50,
+c_max = 0.6
+# Trajectory plots - Alternative examples
+p_alt_04_80 <- plot_trajectories_eRTC(
+mu_ctrl = 0,
+mu_trt = 0.4, # true d = 0.4
+hypothesized_d = 0.4,
+target_power = 0.80,
+alpha = 0.05
+ggsave("traj_eRTC_alt_d0.4_80pow.pdf", p_alt_04_80, width = 8, height = 5)
+# Trajectory plots - Null
+p_null_04_80 <- plot_trajectories_eRTC(
+mu_ctrl = 0,
+mu_trt = NULL, # null
+hypothesized_d = 0.4,
+target_power = 0.80,
+alpha = 0.05
+ggsave("traj_eRTC_null_d0.4_80pow.pdf", p_null_04_80, width = 8, height = 5)
+# e-Survival: Sequential Log-Rank Test (Unified)
+library(tidyverse)
+library(survival) # Used for standard verification if needed
+# Core e-Survival function
+# time: vector of event times
+# status: vector of status (1=event, 0=censored)
+# treatment: vector of assignments (1=trt, 0=ctrl)
+# burn_in: events to observe before betting
+# ramp: number of events to ramp up betting intensity
+# lambda: maximum betting scalar (0 < lambda < 1)
+compute_eSurvival <- function(time, status, treatment,
+burn_in = 20,
+ramp = 50,
+lambda_max = 0.2) {
+n <- length(time)
+# 1. Create a dataframe and sort by time (events happen sequentially)
+df <- data.frame(time, status, treatment) %>%
+arrange(time)
+# Vectors for processing
+T_sorted <- df\$time
+S_sorted <- df\$status
+A_sorted <- df\$treatment
+wealth <- numeric(n)
+wealth[1] <- 1
+# Running Log-Rank Score (Observed - Expected)
+cumulative_Z <- 0
+# Track active Risk Set manually for speed/simplicity
+# Initially, everyone is at risk (assuming time 0 start)
+risk_trt <- sum(treatment == 1)
+risk_ctrl <- sum(treatment == 0)
+# We need to process events one by one.
+# Note: This simple loop assumes no tied times for clarity.
+# (Ties are rare in continuous simulation).
+for (i in 1:n) {
+# --- A. Betting Step (Before observing outcome i) ---
+# Determine bet direction based on history
+if (i > burn_in) {
+# Ramping factor
+c_i <- max(0, min(1, (i - burn_in) / ramp))
+# Direction:
+# If cumulative_Z is negative, Treatment has FEWER events than expected (Benefit).
+# We bet that this trend continues.
+# If Observed (1) - Expected (p) is likely negative, we want a NEGATIVE bet
+# to make the wealth multiplier (1 + bet * U) > 1.
+# Simple strategy: proportional to the sign of the current trend
+bet_direction <- sign(cumulative_Z)
+if (bet_direction == 0) bet_direction <- 0
+# The bet:
+b_i <- c_i * lambda_max * bet_direction
+} else {
+b_i <- 0
+# --- B. Observation Step ---
+is_event <- S_sorted[i] == 1
+tr_is_event <- (A_sorted[i] == 1)
+# Probability that *if* an event happens, it is in Treatment arm
+# p_null = (Risk Treatment) / (Total Risk)
+total_risk <- risk_trt + risk_ctrl
+if (total_risk > 0) {
+p_null <- risk_trt / total_risk
+} else {
+p_null <- 0.5
+# --- C. Update Wealth (Only on events, not censorings) ---
+if (is_event) {
+# U_i: The "score" contribution (Observed - Expected)
+# If Treatment event: 1 - p
+# If Control event: 0 - p
+obs <- ifelse(tr_is_event, 1, 0)
+U_i <- obs - p_null
+# Martingale update
+# Constraint: ensure (1 + b_i * U_i) is non-negative.
+# Since |U_i| < 1 and lambda_max < 1, this is generally safe.
+multiplier <- 1 + b_i * U_i
+# Update history
+cumulative_Z <- cumulative_Z + U_i
+if (i > 1) {
+wealth[i] <- wealth[i-1] * multiplier
+} else {
+wealth[i] <- multiplier
+} else {
+# Censored event: No bet made, wealth carries forward
+if (i > 1) wealth[i] <- wealth[i-1]
+# --- D. Update Risk Set for next step ---
+# The person at index i is removed from risk set (event or censored)
+if (tr_is_event) {
+risk_trt <- max(0, risk_trt - 1)
+} else {
+risk_ctrl <- max(0, risk_ctrl - 1)
+# Map wealth back to original unsorted indices is tricky in sequential plots,
+# so we return the Sequential Wealth (wealth over event time).
+return(data.frame(
+event_num = 1:n,
+time = T_sorted,
+wealth = wealth
+# Simulate a single survival trial (Weibull)
+# HR < 1 implies Treatment is better (longer survival)
+simulate_trial_survival <- function(n,
+HR = 1,
+shape = 1.2,
+scale = 10,
+cens_prop = 0.0) {
+# 1. Treatments
+treatment <- rbinom(n, 1, 0.5)
+# 2. True Times (Weibull AFT)
+# HR in Cox model corresponds roughly to scale change in Weibull
+# Scale_trt = Scale_ctrl / (HR^(1/shape))
+scale_trt <- scale / (HR^(1/shape))
+U <- runif(n)
+true_time <- numeric(n)
+true_time[treatment == 0] <- scale * (-log(U[treatment==0]))^(1/shape)
+true_time[treatment == 1] <- scale_trt * (-log(U[treatment==1]))^(1/shape)
+# 3. Censoring
+# Simple independent uniform censoring to achieve approx proportion
+if (cens_prop > 0) {
+C <- runif(n, 0, 2 * scale) # broad censoring distribution
+# Adjust range to hit target prop roughly if needed, keeping simple here
+time <- pmin(true_time, C)
+status <- as.numeric(true_time <= C)
+} else {
+time <- true_time
+status <- rep(1, n)
+return(list(time = time, status = status, treatment = treatment))
+# Unified simulation function
+simulate_eSurvival <- function(n_sims = 1000,
+n_patients = NULL, # Change to NULL to trigger auto-calc
+HR_true = 1,
+target_HR = 0.7,
+target_power = 0.80, # Added for flexibility
+alpha = 0.05,
+burn_in = 30,
+ramp = 50,
+lambda_max = 0.25) {
+# Determine simulation type
+if (HR_true == 1) {
+sim_type <- "TypeIError(Null)"
+} else {
+sim_type <- "Power(Alternative)"
+# Calculate Sample Size if not provided
+if (is.null(n_patients)) {
+# Schoenfeld formula for events
+z_alpha <- qnorm(1 - alpha/2)
+z_beta <- qnorm(target_power)
+# Total events needed
+req_events <- ceiling( 4 * ((z_alpha + z_beta) / log(target_HR))^2 )
+# Assuming no censoring, Patients = Events
+n_patients <- req_events
+[…]
+target_power*100, target_HR, n_patients))
+[…]
+[…]
+[…]
+rejections <- 0
+first_crossing <- numeric(n_sims)
+pb <- txtProgressBar(min = 0, max = n_sims, style = 3)
+for (sim in 1:n_sims) {
+trial <- simulate_trial_survival(n_patients, HR = HR_true)
+res <- compute_eSurvival(trial\$time, trial\$status, trial\$treatment,
+burn_in = burn_in, ramp = ramp, lambda_max = lambda_max)
+# Check stopping
+crossing <- which(res\$wealth >= 1/alpha)
+if (length(crossing) > 0) {
+rejections <- rejections + 1
+first_crossing[sim] <- crossing[1]
+} else {
+first_crossing[sim] <- NA
+setTxtProgressBar(pb, sim)
+close(pb)
+rate <- rejections / n_sims
+se <- sqrt(rate * (1 - rate) / n_sims)
+median_stop <- median(first_crossing, na.rm = TRUE)
+[…]
+[…]
+if (sim_type == "Power(Alternative)") {
+[…]
+return(list(rate = rate, median_stop = median_stop, n_patients = n_patients))
+# Plot Trajectories
+plot_trajectories_surv <- function(n_trials = 20,
+n_patients = 400,
+HR_true = 0.7,
+alpha = 0.05,
+title = "e-SurvivalTrajectories") {
+trajectories <- list()
+for (i in 1:n_trials) {
+trial <- simulate_trial_survival(n_patients, HR = HR_true)
+res <- compute_eSurvival(trial\$time, trial\$status, trial\$treatment)
+res\$trial <- i
+trajectories[[i]] <- res
+df <- bind_rows(trajectories)
+ggplot(df, aes(x = event_num, y = wealth, group = trial)) +
+geom_line(alpha = 0.5, color = "darkgreen") +
+geom_hline(yintercept = 1/alpha, linetype = "dashed", color = "red") +
+scale_y_log10() +
+labs(
+title = title,
+subtitle = sprintf("TrueHR=%.2f", HR_true),
+x = "NumberofEvents",
+y = "e-value(logscale)"
+annotate("text", x = n_patients, y = 1/alpha,
+label = sprintf("Threshold%.0f", 1/alpha), vjust = -0.5, color="red") +
+theme_minimal()
+# Run Scenarios
+if (interactive()) {
+# Define design parameters
+target_hr <- 0.80
+# 1. Check Type I Error (Null: HR = 1.0)
+# Passing n_patients = NULL triggers the auto-calculation based on target_HR
+simulate_eSurvival(
+n_sims = 2000,
+n_patients = NULL, # Auto-calculate N for 80% power
+HR_true = 1.0, # Reality: Null
+target_HR = target_hr,
+alpha = 0.05
+# 2. Check Power (Alternative: HR = 0.7)
+simulate_eSurvival(
+n_sims = 2000,
+n_patients = NULL, # Auto-calculate N for 80% power
+HR_true = 0.8, # Reality: Alternative works
+target_HR = target_hr,
+alpha = 0.05
+# 3. Plot Trajectories
+# We calculate N explicitly here just to pass it to the plot function
+req_events <- ceiling(4 * ((qnorm(0.975) + qnorm(0.80)) / log(target_hr))^2)
+p <- plot_trajectories_surv(
+n_trials = 30,
+n_patients = req_events,
+HR_true = 0.8,
+alpha = 0.05,
+title = sprintf("e-Survival(TargetHR=%.2f,N=%d)", target_hr, req_events)
+print(p)
+set.seed(2)
+p1<-plot_trajectories_surv(
+n_patients = ceiling(4 * ((qnorm(0.975) + qnorm(0.80)) / log(0.8))^2),
+HR_true = 0.8)
+ggsave("traj_survivaleRT_08.pdf",p1,
+width = 8, height = 5)
+p2<-plot_trajectories_surv(
+n_patients = ceiling(4 * ((qnorm(0.975) + qnorm(0.80)) / log(0.8))^2),
+HR_true = 1)
+p2
+ggsave("traj_survivaleRT_null.pdf",p2,
+width = 8, height = 5)
+# APPENDIX: Verification of Staggered vs. Simultaneous Entry
+# This script demonstrates that the "Simultaneous Entry" assumption used in the
+# main simulation yields statistically identical results to a realistic
+# "Staggered Entry" design, provided the analysis is based on "Time on Study".
+verify_staggered_equivalence <- function(n_sims = 1000,
+n_patients = 631, # Updated to match HR=0.8 scenario
+HR_true = 0.8, # Updated to match HR=0.8 scenario
+recruit_period = 12, # Recruitment over 12 months
+alpha = 0.05) {
+[…]
+# Storage for final e-values
+final_e_simultaneous <- numeric(n_sims)
+final_e_staggered <- numeric(n_sims)
+pb <- txtProgressBar(min = 0, max = n_sims, style = 3)
+for (i in 1:n_sims) {
+# 1. Generate the underlying biology (Survival Times)
+# -- Method A: Simultaneous (The Simple Way) --
+trial_sim <- simulate_trial_survival(n_patients, HR = HR_true)
+res_sim <- compute_eSurvival(trial_sim\$time, trial_sim\$status, trial_sim\$treatment)
+final_e_simultaneous[i] <- tail(res_sim\$wealth, 1)
+# -- Method B: Staggered (The "Real World" Way) --
+# We generate fresh biology for the staggered arm to compare distributions
+trial_stag <- simulate_trial_survival(n_patients, HR = HR_true)
+# ... but we add "Recruitment Time" (Uniform over 12 months)
+entry_time <- runif(n_patients, 0, recruit_period)
+# The event happens at this Calendar Time:
+calendar_event_time <- entry_time + trial_stag\$time
+# The analyst calculates "Time on Study" (Analysis Time):
+# Analysis Time = Calendar Event Time - Entry Time
+calculated_study_time <- calendar_event_time - entry_time
+# Run e-process on the calculated study time
+# Note: compute_eSurvival handles the sorting of events internally
+res_stag <- compute_eSurvival(calculated_study_time,
+trial_stag\$status,
+trial_stag\$treatment)
+final_e_staggered[i] <- tail(res_stag\$wealth, 1)
+setTxtProgressBar(pb, i)
+close(pb)
+# 2. Compare Results
+df_res <- data.frame(
+e_value = c(final_e_simultaneous, final_e_staggered),
+Method = rep(c("Simultaneous(Simplified)", "Staggered(Realistic)"), each = n_sims)
+# Summary Stats
+print(df_res %>%
+group_by(Method) %>%
+summarise(
+Mean_E = mean(e_value),
+Median_E = median(e_value),
+Power_Estimated = mean(e_value >= 1/alpha)
+# 3. Visualization
+p <- ggplot(df_res, aes(x = e_value, fill = Method)) +
+geom_density(alpha = 0.5, color = NA) +
+scale_x_log10() +
+labs(
+title = "EquivalenceofSimultaneousvs.StaggeredEntry",
+subtitle = sprintf("Distributionoffinale-values(N=%d,HR=%.2f)", n_patients, HR_true),
+x = "Finale-value",
+y = "Density"
+theme_minimal() +
+theme(legend.position = "bottom")
+return(p)
+# Run the verification
+if (interactive()) {
+# Matching the N=631 / HR=0.8 scenario from the paper
+p_verify <- verify_staggered_equivalence(
+n_sims = 1000,
+n_patients = 631,
+HR_true = 0.8
+print(p_verify)
+# Optional: Save for supplement
+ggsave("supp_staggered_equivalence.pdf", p_verify, width = 8, height = 5)

### v4

@@
 Under the null hypothesis, randomization ensures that assignments are independent of outcomes, so no betting strategy can systematically accumulate wealth.
-We discuss potential methods for trial monitoring based on previous efforts.
-For binary outcomes, our construction implements the i-bet framework with a specific adaptive betting strategy.
-We present methods with unified notation, validity proofs, and simulation studies demonstrating Type I error control and power characteristics.
+We propose e-RT (e-value Randomized Trial), a method for prospective sequential monitoring of randomized trials.
+Like i-bet (pmlr-v177-duan22a), e-RT uses betting martingales for inference, but differs in key respects: e-RT monitors sequentially as patients enroll rather than analyzing completed trial data; it requires no covariates or working models; and it uses an adaptive betting fraction derived from running outcome estimates rather than fixed wagers.
+This yields a method with minimal assumptions suitable for real-time trial monitoring.
 2 Overall construction
@@
 3 Simulation Studies
-We evaluated operating characteristics by simulation.
+We evaluated operating characteristics or e-RT by simulation.
 For each scenario, we calculated the sample size required for a chi-square test to achieve the target power at $\alpha=0.05$, then ran 5,000 simulated trials at that sample size.
@@
 4 Continuous Outcomes
-The previous approach treats each patient as a single Bernoulli trial: the outcome (event vs.
+The e-RT treats each patient as a single Bernoulli trial: the outcome (event vs.
 no event) is observed, and we bet on which arm that patient came from.
@@
 Therefore, defining wager is slightly more granular.
-One may extend this to continuous endpoint.
+One may extend this to continuous endpoint which we shall call e-RTC.
 The validity still comes only from randomization: under the null hypothesis, the distribution of the continuous outcome is the same in both arms, so the outcome does not help to predict treatment assignment.
@@
 4.6 Simulation overview
-We evaluated this using the same design philosophy as for binary approach.
+We evaluated e-RTC using the same design philosophy as for binary approach.
 For a given standardized effect size (Cohen’s $d$) and target power, we first computed the fixed-sample size required for a two-sample $t$-test at $\alpha=0.05$.
@@
 We attempted to construct a nonparametric approach where validity derives solely from randomization, not from a correctly specified hazard model.
+We call this e-survival.
 5.1 Setup and Martingale Construction
@@
 6 Discussion
-The binary implementation is a nonparametric sequential test for randomized trials based on the betting framework for e-values (i-bet pmlr-v177-duan22a).
+The e-Rt is an nonparametric sequential test for randomized trials based on the betting framework for e-values (i-bet pmlr-v177-duan22a).
 The method requires only that treatment assignment is randomized — no distributional assumptions about outcomes are needed.
@@
 6.1 Operating characteristics
-Simulations demonstrate that the methos properly controlled for Type I error (2–3% vs nominal 5%) while providing approximately 50% power for early stopping in trials designed with 80% power, and 63–66% power in trials designed with 90% power.
+Simulations demonstrate that the method properly controlled for Type I error (2–3% vs nominal 5%) while providing approximately 50% power for early stopping in trials designed with 80% power, and 63–66% power in trials designed with 90% power.
 When early stopping occurs, it happens at roughly half the planned sample size.
-Both the continuous and the survival implementations had similar characteristics: Power was traded for continuous Type 1 error control.
+Both the continuous (e-RTC) and the survival (e-survival) implementations had similar characteristics: Power was traded for continuous Type 1 error control.
 All simulations showed that Type 1 error was exceptionally well controlled but power was nominally low.
@@
 This offers a computationally simple, algorithmic alternative that derives validity strictly from the randomization probabilities within the risk set, without requiring the full apparatus of partial likelihood theory.
+The binary e-RT shares the same martingale foundation as i-bet but differs in key respects: it operates prospectively as patients enroll rather than retrospectively on completed data; it requires no covariates or working models; and it uses betting fractions that adapt continuously to running outcome estimates rather than fixed-magnitude wagers guided by covariate-based predictions.
+This yields a simpler method that may be suitable for real-time trial monitoring.
 It is possible that some of the concepts here were discussed by other authors in different contexts that were not immediately available for this author.
@@
   3. 3.
-    Third Version (this version): Added continuous and survival endpoints; text adjustments.
+    Third Version (Dec 11, 2025): Added continuous and survival endpoints; text adjustments.
+  4. 4.
+    Forth Version (This version): Correction on signal direction on -RTC. Text adjustments
 Appendix A R Code
@@
 g_i <- s_i / (1 + abs(s_i)) # in (-1,1), robust to outliers
+# Learn direction from past data (NEW)
+trt_prev <- treatment[1:(i-1)]
+mean_trt <- mean(out_prev[trt_prev == 1])
+mean_ctrl <- mean(out_prev[trt_prev == 0])
+# Handle edge cases early in trial
+if (is.nan(mean_trt) $|$ is.nan(mean_ctrl)) {
+direction <- 0
+} else {
+direction <- sign(mean_trt - mean_ctrl)
 # Ramping betting strength
@@
 c_i <- ramp_frac * c_max
-lambda <- 0.5 + c_i * g_i
+lambda <- 0.5 + c_i * g_i * direction
 lambda <- max(0.001, min(0.999, lambda))

### v5

@@
 When these assumptions fail, or when trials adapt in ways not fully prespecified, validity guarantees may erode.
-Monitoring clinical trials, specially in the context of acutely ill patients, is of paramount importance.
+Monitoring clinical trials, especially in the context of acutely ill patients, is of paramount importance.
 Traditional methods include $\alpha-spending$ functions.
@@
   1. 1.
-    A center and scale for the outcomes observed so far. Outcomes con be all all over the place. We need to standardize.
+    A center and scale for the outcomes observed so far. Outcomes can be all over the place. We need to standardize.
   2. 2.
@@
 0.60 90% 120 0.045 58.5% 65
-A visual representation can be found in figure below.
+A visual representation is shown in Figure 3.
 Figure 3: Trajectories of the continuous randomization e-process for a trial designed to detect a standardized mean difference of $d=0.40$ with 80% power.
@@
 The red dashed line represents the rejection threshold ($1/\alpha=20$).
+6 Multi-State Models
+Clinical trials in critical care increasingly use multi-state models to capture patient trajectories.
+A patient in the ICU may improve to a general ward, be discharged home, or die—and some of these transitions can reverse.
+Traditional analyses model the full transition matrix or use competing risks methods, requiring assumptions about the dependence structure between outcomes.
+The e-RT framework extends naturally to this setting.
+Rather than modeling the Markov structure, we classify each transition as “good” (recovery-oriented) or “bad” (deterioration-oriented) and bet on this binary outcome.
+The method is agnostic to the underlying stochastic process: we simply ask whether good transitions predict treatment assignment.
+6.1 Setup
+Consider a four-state model common in ICU trials:
+  • State 1: General Ward
+  • State 2: ICU
+  • State 3: Home (absorbing)
+  • State 4: Dead (absorbing)
+Patients begin in the ICU and are followed for a fixed period (e.g., 28 days). Each day, a patient may transition between states according to arm-specific transition probabilities.
+We define good transitions as those representing clinical improvement:
+  • ICU $\to$ Ward (step-down from intensive care)
+  • Ward $\to$ Home (discharge)
+All other transitions—including ICU $\to$ Dead, Ward $\to$ Dead, and Ward $\to$ ICU (readmission)—are classified as bad.
+The null hypothesis is that the rate of good transitions is equal between arms:
+$H_{0}:P(\text{good transition}\mid\text{treatment})=P(\text{good transition}\mid\text{control})$ (23)
+6.2 Betting Strategy
+Each state transition generates one betting opportunity.
+Unlike the binary e-RT where each patient contributes one observation, here a single patient may contribute multiple transitions as they move through states.
+A patient who goes ICU $\to$ Ward $\to$ ICU $\to$ Ward $\to$ Home contributes four transitions and four bets.
+Let $\hat{\delta}_{i-1}$ be the difference in good-transition rates between arms, estimated from all transitions observed before transition $i$:
+[…]
+The betting fraction follows the same logic as binary e-RT:
+[…]
+where $c_{i}$ ramps from 0 to 1 over a burn-in period, exactly as before.
+The wealth update remains:
+$W_{i}=W_{i-1}\times\begin{cases}\lambda_{i}/0.5&\text{if treatment arm}\\ (1-\lambda_{i})/0.5&\text{if control arm}\end{cases}$ (26)
+6.3 Validity
+The martingale property holds by the same argument as before.
+Under the null, treatment assignment is independent of transition quality.
+Given any transition (good or bad), the probability it came from the treatment arm is 0.5.
+The expected wealth multiplier is therefore 1, and $(W_{i})$ is a test martingale.
+The key insight is that validity does not depend on the Markov assumption or any specific transition structure.
+We have simply replaced “patient with event” with “transition classified as good” and applied the same betting logic.
+The data-generating process happens to be Markovian in our simulations, but the test would remain valid for any process that generates a sequence of binary-classified transitions from randomized patients.
+6.4 Simulation Study
+We simulated trials with the following daily transition probabilities:
+Table 4: Daily Transition Probabilities
+To State
+From State Ward ICU Home Dead
+Control
+Ward 0.880 0.070 0.030 0.020
+ICU 0.070 0.915 0.000 0.015
+Treatment
+Ward 0.870 0.050 0.050 0.030
+ICU 0.090 0.900 0.000 0.010
+Treatment improves recovery transitions: ICU $\to$ Ward increases from 7% to 9% daily, and Ward $\to$ Home increases from 3% to 5% daily. The effect on mortality is modest.
+All patients started in the ICU and were followed for 28 days. At day 28, the expected state distributions were:
+Table 5: Day 28 State Distribution
+Arm Ward ICU Home Dead
+Control 20.8% 26.3% 18.8% 34.1%
+Treatment 16.9% 16.6% 33.5% 33.0%
+Difference $3.9\%$ $9.7\%$ $4.7\% 1.1$$1\%$
+Treatment substantially increases discharge home ($14.7\%$) while barely changing mortality ($1.1\%$).
+This is precisely the scenario where trajectory-based analysis adds value: a mortality-only endpoint would detect little effect, while multi-state analysis captures the clinically meaningful improvement in recovery.
+We used burn-in = 30 transitions, ramp = 50 transitions, and threshold $1/\alpha=20$. Results from 1,000 simulated trials at $N=1000$ patients:
+Table 6: Operating Characteristics for e-RTms ($N=1000$)
+Metric Value
+Type I Error 0.30% (SE: 0.17%)
+Power 89.3% (SE: 1.0%)
+Median transitions per trial 2017
+Median crossing 898 transitions (45%)
+Type I error was well below the nominal 5%, consistent with the conservative behavior observed in other e-RT variants.
+Power was 89% at $N=1000$; sample size search indicated approximately $N=800$ for 80% power.
+The median trial generated approximately 2000 transitions from 1000 patients (roughly 2 transitions per patient).
+When the threshold was crossed, it occurred at 898 transitions—45% of the way through the trial.
+6.5 Trajectory Examples
+Figure 5 shows 30 example wealth trajectories under the null and alternative hypotheses.
+Under the null (left panel), wealth fluctuates around 1 with the characteristic downward drift from betting on noise.
+Under the alternative (right panel), wealth grows systematically as good transitions accumulate preferentially in the treatment arm.
+Figure 5: Trajectories of e-RTms for a trial with $N=1000$ patients.
+Left: under the null hypothesis (equal transition matrices), wealth fluctuates randomly.
+Right: under the alternative hypothesis, wealth grows as treatment improves recovery transitions.
+Dashed red line: rejection threshold ($1/\alpha=20$).
+6.6 Interpretation
+The e-RTms approach offers a philosophically clean handling of multi-state outcomes.
+By collapsing the transition matrix into a single binary classification, we sidestep the complexities of competing risks and counterfactual reasoning.
+We do not ask “what would have happened to this patient had they been in the other arm?” We simply ask “does knowing whether this transition was good help predict which arm the patient was in?”
+This comes at a cost: we lose information about which specific transitions drive the effect.
+A treatment that dramatically improves ICU $\to$ Ward transitions but worsens Ward $\to$ Home transitions might show no overall effect if the good-transition rates balance.
+For monitoring purposes, this limitation is acceptable—the goal is to detect any departure from exchangeability, not to dissect the mechanism.
+Detailed transition-specific analyses can follow at trial completion.
+e-RT is a monitoring, not a reporting, method.
+The method is most valuable when treatment affects trajectory without substantially changing mortality— the pattern in our simulation.
+For trials where mortality is the dominant signal, e-survival may be more powerful.
+For trials where the clinical question is “does treatment help patients recover faster?”, e-RTms provides a natural framework.
 7 Discussion
-The e-Rt is an nonparametric sequential test for randomized trials based on the betting framework for e-values (i-bet pmlr-v177-duan22a).
+The e-RT is a nonparametric sequential test for randomized trials based on the betting framework for e-values (i-bet pmlr-v177-duan22a).
 The method requires only that treatment assignment is randomized — no distributional assumptions about outcomes are needed.
@@
   4. 4.
-    Forth Version (This version): Correction on signal direction on -RTC. Text adjustments
+    Fourth Version (Dec 17, 2025): Correction on signal direction on e-RTC. Text adjustments
+  5. 5.
+    Fifth Version (Dec 31, 2025): Added multi-state extension (e-RTms).
 Appendix A R Code
-# e-RT Simulations: Type I Error and Power (Unified)
+# e-RT: Sequential Randomization Tests Using e-values
+# Supplementary R Code
 library(tidyverse)
-# Core e-RT function
+# --- e-RT: Binary Outcomes ---
 compute_eRT <- function(treatment, outcome, p = 0.5, burn_in = 50, ramp = 100) {
@@
 c_i <- max(0, min(1, (i - burn_in) / ramp))
-if (outcome[i] == 1) {
-lambda <- 0.5 + 0.5 * c_i * delta_hat
+lambda <- if (outcome[i] == 1) {
+0.5 + 0.5 * c_i * delta_hat
 } else {
-lambda <- 0.5 - 0.5 * c_i * delta_hat
+0.5 - 0.5 * c_i * delta_hat
 lambda <- max(0.001, min(0.999, lambda))
-if (treatment[i] == 1) {
-multiplier <- lambda / p
-} else {
-multiplier <- (1 - lambda) / (1 - p)
+multiplier <- if (treatment[i] == 1) lambda / p else (1 - lambda) / (1 - p)
 wealth[i] <- wealth[i-1] * multiplier
 return(wealth)
-# Simulate a single trial
 simulate_trial <- function(n, p_trt = 0.5, rate_trt, rate_ctrl) {
@@
 outcome[treatment == 0] <- rbinom(sum(treatment == 0), 1, rate_ctrl)
-return(data.frame(treatment = treatment, outcome = outcome))
-# Unified simulation function
-simulate_eRT <- function(n_sims = 5000,
-p_ctrl,
-p_trt = NULL,
-hypothesized_ARR = NULL,
-target_power = 0.80,
-alpha = 0.05,
-burn_in = 50,
-ramp = 100) {
-# Determine hypothesized ARR for sample size calculation
+data.frame(treatment = treatment, outcome = outcome)
+simulate_eRT <- function(n_sims = 5000, p_ctrl, p_trt = NULL,
+hypothesized_ARR = NULL, target_power = 0.80,
+alpha = 0.05, burn_in = 50, ramp = 100) {
 if (is.null(hypothesized_ARR) && is.null(p_trt)) {
 stop("Mustspecifyeitherp_trtorhypothesized_ARR")
-if (is.null(hypothesized_ARR)) {
-hypothesized_ARR <- p_ctrl - p_trt
-# Calculate sample size based on hypothesized effect
+if (is.null(hypothesized_ARR)) hypothesized_ARR <- p_ctrl - p_trt
 $s$ <- power.prop.test(p1 = p_ctrl, p2 = p_ctrl - hypothesized_ARR,
@@
 n_patients <- 2 * n_per_arm
-# Determine true p_trt (null if not specified)
 if (is.null(p_trt)) {
-p_trt <- p_ctrl # Null is true
+p_trt <- p_ctrl
 true_ARR <- 0
-sim_type <- "TypeIError"
 } else {
 true_ARR <- p_ctrl - p_trt
-sim_type <- "Power"
-[…]
-[…]
-p_ctrl, hypothesized_ARR, target_power * 100))
-[…]
-[…]
-p_ctrl, p_trt, true_ARR))
-[…]
 rejections <- 0
@@
 trial <- simulate_trial(n_patients, rate_trt = p_trt, rate_ctrl = p_ctrl)
-wealth <- compute_eRT(trial\$treatment, trial\$outcome, burn_in = burn_in, ramp = ramp)
+wealth <- compute_eRT(trial\$treatment, trial\$outcome,
+burn_in = burn_in, ramp = ramp)
 final_evalues[sim] <- wealth[n_patients]
@@
 se <- sqrt(rejection_rate * (1 - rejection_rate) / n_sims)
-median_stop <- median(first_crossing, na.rm = TRUE)
-[…]
-[…]
-[…]
-if (sim_type == "Power") {
-[…]
-[…]
-} else {
-[…]
-return(list(
-sim_type = sim_type,
+list(
 rejection_rate = rejection_rate,
@@
 true_ARR = true_ARR,
-target_power = target_power,
-alpha = alpha,
-median_crossing = median_stop,
+median_crossing = median(first_crossing, na.rm = TRUE),
 first_crossing = first_crossing,
 final_evalues = final_evalues
-# Plot trajectories
-plot_trajectories <- function(n_trials = 30,
-p_ctrl,
-p_trt = NULL,
-hypothesized_ARR = NULL,
-target_power = 0.80,
-alpha = 0.05,
-burn_in = 50,
-ramp = 100,
+plot_trajectories <- function(n_trials = 30, p_ctrl, p_trt = NULL,
+hypothesized_ARR = NULL, target_power = 0.80,
+alpha = 0.05, burn_in = 50, ramp = 100,
 title = NULL) {
@@
 stop("Mustspecifyeitherp_trtorhypothesized_ARR")
-if (is.null(hypothesized_ARR)) {
-hypothesized_ARR <- p_ctrl - p_trt
-if (is.null(p_trt)) {
-p_trt <- p_ctrl # Null
-# Calculate sample size
+if (is.null(hypothesized_ARR)) hypothesized_ARR <- p_ctrl - p_trt
+if (is.null(p_trt)) p_trt <- p_ctrl
 $s$ <- power.prop.test(p1 = p_ctrl, p2 = p_ctrl - hypothesized_ARR,
@@
 if (is.null(title)) {
-if (true_ARR == 0) {
-title <- sprintf("e-RTUnderNull(n=%d,designedfor%.0f%%ARR)",
+title <- if (true_ARR == 0) {
+sprintf("e-RTUnderNull(n=%d,designedfor%.0f%%ARR)",
 n_patients, hypothesized_ARR * 100)
 } else {
-title <- sprintf("e-RTUnderAlternative(n=%d,%.0f%%power)",
+sprintf("e-RTUnderAlternative(n=%d,%.0f%%power)",
 n_patients, target_power * 100)
-trajectories <- list()
-for (i in 1:n_trials) {
+trajectories <- lapply(1:n_trials, function(i) {
 trial <- simulate_trial(n_patients, rate_trt = p_trt, rate_ctrl = p_ctrl)
-wealth <- compute_eRT(trial\$treatment, trial\$outcome, burn_in = burn_in, ramp = ramp)
-trajectories[[i]] <- data.frame(
-patient = 1:n_patients,
-wealth = wealth,
-trial = i
+wealth <- compute_eRT(trial\$treatment, trial\$outcome,
+burn_in = burn_in, ramp = ramp)
+data.frame(patient = 1:n_patients, wealth = wealth, trial = i)
 df <- bind_rows(trajectories)
-p <- ggplot(df, aes(x = patient, y = wealth, group = trial)) +
+ggplot(df, aes(x = patient, y = wealth, group = trial)) +
 geom_line(alpha = 0.4, color = "steelblue") +
@@
 scale_y_log10() +
-labs(
-title = title,
+labs(title = title,
 subtitle = sprintf("Control=%.0f%%,Treatment=%.0f%%,TrueARR=%.0f%%",
 p_ctrl * 100, p_trt * 100, true_ARR * 100),
-x = "Patient",
-y = "e-value(logscale)"
+x = "Patient", y = "e-value(logscale)") +
 annotate("text", x = n_patients * 0.95, y = 1/alpha * 1.5,
@@
 theme_minimal() +
-theme(
-plot.title = element_text(face = "bold"),
-panel.grid.minor = element_blank()
-return(p)
-# Run all simulations
-run_all <- function(n_sims = 5000,
-p_ctrl = 0.40,
-ARRs = c(0.05, 0.10),
-target_powers = c(0.80, 0.90),
-alpha = 0.05,
-seed = 42) {
-set.seed(seed)
-results <- list()
-# Create grid of scenarios
-scenarios <- expand.grid(
-hypothesized_ARR = ARRs,
-target_power = target_powers,
-stringsAsFactors = FALSE
-# For each scenario, run Type I error and Power
-all_results <- data.frame()
-for (i in 1:nrow(scenarios)) {
-hyp_ARR <- scenarios\$hypothesized_ARR[i]
-tgt_pow <- scenarios\$target_power[i]
-[…]
-[…]
-hyp_ARR * 100, tgt_pow * 100))
-[…]
-# Type I error (null true, same sample size)
-[…]
-t1 <- simulate_eRT(
-n_sims = n_sims,
-p_ctrl = p_ctrl,
-p_trt = NULL, # Null is true
-hypothesized_ARR = hyp_ARR,
-target_power = tgt_pow,
-alpha = alpha
-# Power (alternative true)
-[…]
-pow <- simulate_eRT(
-n_sims = n_sims,
-p_ctrl = p_ctrl,
-p_trt = p_ctrl - hyp_ARR, # True effect = hypothesized
-target_power = tgt_pow,
-alpha = alpha
-all_results <- rbind(all_results, data.frame(
-hypothesized_ARR = hyp_ARR,
-target_power = tgt_pow,
-n_patients = t1\$n_patients,
-type1_error = t1\$rejection_rate,
-type1_se = t1\$se,
-eRT_power = pow\$rejection_rate,
-power_se = pow\$se,
-median_crossing = pow\$median_crossing
-results\$summary <- all_results
-# Print summary
-[…]
-[…]
-[…]
-print(all_results %>%
-mutate(
-hypothesized_ARR = sprintf("%.0f%%", hypothesized_ARR * 100),
-target_power = sprintf("%.0f%%", target_power * 100),
-type1_error = sprintf("%.3f", type1_error),
-eRT_power = sprintf("%.1f%%", eRT_power * 100)
-select(hypothesized_ARR, target_power, n_patients,
-type1_error, eRT_power, median_crossing))
-return(results)
-# Run
-if (interactive()) {
-results <- run_all(
-n_sims = 5000,
-p_ctrl = 0.40,
-ARRs = c(0.05, 0.10),
-target_powers = c(0.80, 0.90),
-alpha = 0.05
-# Trajectory plots - Alternative
-p1 <- plot_trajectories(p_ctrl = 0.40, p_trt = 0.30, target_power = 0.80)
-ggsave("traj_alt_10pct_80pow.pdf", p1, width = 8, height = 5)
-p2 <- plot_trajectories(p_ctrl = 0.40, p_trt = 0.30, target_power = 0.90)
-ggsave("traj_alt_10pct_90pow.pdf", p2, width = 8, height = 5)
-# Trajectory plots - Null
-p3 <- plot_trajectories(p_ctrl = 0.40, p_trt = NULL, hypothesized_ARR = 0.10, target_power = 0.80)
-ggsave("traj_null_10pct_80pow.pdf", p3, width = 8, height = 5)
-p4 <- plot_trajectories(p_ctrl = 0.40, p_trt = NULL, hypothesized_ARR = 0.10, target_power = 0.90)
-ggsave("traj_null_10pct_90pow.pdf", p4, width = 8, height = 5)
-# e-RTC (Randomization E-Process for Continuous Outcomes)
-# Type I Error and Power Simulations (Unified)
-setwd("")
-suppressPackageStartupMessages({
-library(tidyverse)
-# Core e-RTC function
-# treatment: 0/1 vector (0 = control, 1 = intervention)
-# outcome : continuous outcome (numeric)
-# p : allocation prob for treatment (default 0.5)
-# burn_in : number of patients before bets start ramping up
-# ramp : additional patients over which betting strength ramps to max
-# c_max : maximum betting strength (0 < c_max <= 0.5 recommended)
-compute_eRTC <- function(treatment, outcome,
-p = 0.5,
-burn_in = 20,
-ramp = 50,
-c_max = 0.6) {
+theme(plot.title = element_text(face = "bold"), panel.grid.minor = element_blank())
+# --- e-RTC: Continuous Outcomes ---
+compute_eRTC <- function(treatment, outcome, p = 0.5, burn_in = 20,
+ramp = 50, c_max = 0.6) {
 n <- length(treatment)
@@
 for (i in 2:n) {
-# Past outcomes
 out_prev <- outcome[1:(i-1)]
-# Require enough past data before betting
 if ((i - 1) < burn_in) {
@@
 next
-# Robust center and scale: median + MAD
 med_prev <- median(out_prev, na.rm = TRUE)
@@
 if (!is.finite(mad_prev) $|$ mad_prev <= 0) mad_prev <- 1
-# Current outcome
 y_i <- outcome[i]
-# Standardized residual, squashed to (-1, 1)
 r_i <- y_i - med_prev
 s_i <- r_i / mad_prev
-g_i <- s_i / (1 + abs(s_i)) # in (-1,1), robust to outliers
-# Learn direction from past data (NEW)
+g_i <- s_i / (1 + abs(s_i))
 trt_prev <- treatment[1:(i-1)]
@@
 mean_ctrl <- mean(out_prev[trt_prev == 0])
-# Handle edge cases early in trial
-if (is.nan(mean_trt) $|$ is.nan(mean_ctrl)) {
-direction <- 0
-} else {
-direction <- sign(mean_trt - mean_ctrl)
-# Ramping betting strength
+direction <- if (is.nan(mean_trt) $|$ is.nan(mean_ctrl)) 0 else sign(mean_trt - mean_ctrl)
 ramp_frac <- max(0, min(1, (i - burn_in) / ramp))
@@
 lambda <- max(0.001, min(0.999, lambda))
-# Wealth update based on actual assignment
-if (treatment[i] == 1) {
-multiplier <- lambda / p
-} else {
-multiplier <- (1 - lambda) / (1 - p)
+multiplier <- if (treatment[i] == 1) lambda / p else (1 - lambda) / (1 - p)
 wealth[i] <- wealth[i-1] * multiplier
 return(wealth)
-# Simulate a single trial with continuous outcomes
-# n : total number of patients
-# p_trt : probability of being assigned to treatment (default 0.5)
-# mu_ctrl : mean in control arm
-# mu_trt : mean in treatment arm
-# sd : common standard deviation (default 1)
-simulate_trial_continuous <- function(n,
-p_trt = 0.5,
-mu_ctrl = 0,
-mu_trt = 0,
-sd = 1) {
+simulate_trial_continuous <- function(n, p_trt = 0.5, mu_ctrl = 0, mu_trt = 0, sd = 1) {
 treatment <- rbinom(n, 1, p_trt)
@@
 outcome[treatment == 1] <- rnorm(sum(treatment == 1), mean = mu_trt, sd = sd)
-return(data.frame(treatment = treatment, outcome = outcome))
-# Unified simulation function for e-RTC
-# n_sims : number of simulated trials
-# mu_ctrl : control mean (default 0)
-# mu_trt : treatment mean (NULL => null, mu_trt = mu_ctrl)
-# hypothesized_d: Cohen’s d used for sample size calc (delta/sd)
-# target_power : design power (0.80, 0.90)
-# alpha : significance level
-# burn_in, ramp, c_max: e-RTC tuning parameters
-simulate_eRTC <- function(n_sims = 5000,
-mu_ctrl = 0,
-mu_trt = NULL,
-hypothesized_d,
-target_power = 0.80,
-alpha = 0.05,
-burn_in = 20,
-ramp = 50,
-c_max = 0.6) {
-if (missing(hypothesized_d)) {
-stop("Mustspecifyhypothesized_d(Cohen’sdfordesign).")
-# Treat SD = 1 for design (Cohen’s d = delta / sd)
+data.frame(treatment = treatment, outcome = outcome)
+simulate_eRTC <- function(n_sims = 5000, mu_ctrl = 0, mu_trt = NULL,
+hypothesized_d, target_power = 0.80, alpha = 0.05,
+burn_in = 20, ramp = 50, c_max = 0.6) {
+if (missing(hypothesized_d)) stop("Mustspecifyhypothesized_d")
 sd_true <- 1
 delta_design <- hypothesized_d * sd_true
-# Sample size via t-test (two-sample, equal sizes)
-$s$ <- power.t.test(delta = delta_design,
-sd = sd_true,
-power = target_power,
-sig.level = alpha,
-type = "two.sample",
-alternative = "two.sided")
+$s$ <- power.t.test(delta = delta_design, sd = sd_true, power = target_power,
+sig.level = alpha, type = "two.sample", alternative = "two.sided")
 n_per_arm <- ceiling(ss\$n)
 n_patients <- 2 * n_per_arm
-# Determine true mu_trt (null if not specified)
 if (is.null(mu_trt)) {
-mu_trt <- mu_ctrl # Null is true
+mu_trt <- mu_ctrl
 true_d <- 0
-sim_type <- "TypeIError"
 } else {
 true_d <- (mu_trt - mu_ctrl) / sd_true
-sim_type <- "Power"
-[…]
-[…]
-mu_ctrl, hypothesized_d, target_power * 100))
-[…]
-[…]
-mu_ctrl, mu_trt, true_d))
-[…]
 rejections <- 0
@@
 for (sim in 1:n_sims) {
-trial <- simulate_trial_continuous(n_patients,
-mu_ctrl = mu_ctrl,
-mu_trt = mu_trt,
-sd = sd_true)
-wealth <- compute_eRTC(trial\$treatment,
-trial\$outcome,
-p = 0.5,
-burn_in = burn_in,
-ramp = ramp,
-c_max = c_max)
+trial <- simulate_trial_continuous(n_patients, mu_ctrl = mu_ctrl,
+mu_trt = mu_trt, sd = sd_true)
+wealth <- compute_eRTC(trial\$treatment, trial\$outcome, p = 0.5,
+burn_in = burn_in, ramp = ramp, c_max = c_max)
 final_evalues[sim] <- wealth[n_patients]
@@
 se <- sqrt(rejection_rate * (1 - rejection_rate) / n_sims)
-median_stop <- median(first_crossing, na.rm = TRUE)
-[…]
-[…]
-[…]
-rejection_rate - 1.96 * se,
-rejection_rate + 1.96 * se))
-if (sim_type == "Power") {
-[…]
-[…]
-} else {
-[…]
-return(list(
-sim_type = sim_type,
+list(
 rejection_rate = rejection_rate,
@@
 true_d = true_d,
-target_power = target_power,
-alpha = alpha,
-median_crossing = median_stop,
+median_crossing = median(first_crossing, na.rm = TRUE),
 first_crossing = first_crossing,
 final_evalues = final_evalues
-# Plot trajectories under null / alternative
-# n_trials : number of trajectories to overlay
-# mu_ctrl, mu_trt, hypothesized_d, target_power, alpha, burn_in, ramp, c_max
-plot_trajectories_eRTC <- function(n_trials = 30,
-mu_ctrl = 0,
-mu_trt = NULL,
-hypothesized_d,
-target_power = 0.80,
-alpha = 0.05,
-burn_in = 20,
-ramp = 50,
-c_max = 0.6,
-title = NULL) {
-if (missing(hypothesized_d)) {
-stop("Mustspecifyhypothesized_d(Cohen’sd).")
+plot_trajectories_eRTC <- function(n_trials = 30, mu_ctrl = 0, mu_trt = NULL,
+hypothesized_d, target_power = 0.80,
+alpha = 0.05, burn_in = 20, ramp = 50,
+c_max = 0.6, title = NULL) {
+if (missing(hypothesized_d)) stop("Mustspecifyhypothesized_d")
 sd_true <- 1
-delta_design <- hypothesized_d * sd_true
-$s$ <- power.t.test(delta = delta_design,
-sd = sd_true,
-power = target_power,
-sig.level = alpha,
-type = "two.sample",
-alternative = "two.sided")
+$s$ <- power.t.test(delta = hypothesized_d, sd = sd_true, power = target_power,
+sig.level = alpha, type = "two.sample", alternative = "two.sided")
 n_patients <- 2 * ceiling(ss\$n)
 if (is.null(mu_trt)) {
-mu_trt <- mu_ctrl # Null
+mu_trt <- mu_ctrl
 true_d <- 0
@@
 if (is.null(title)) {
-if (true_d == 0) {
-title <- sprintf(
-"e-RTCUnderNull(n=%d,designedford=%.2f)",
-n_patients, hypothesized_d
+title <- if (true_d == 0) {
+sprintf("e-RTCUnderNull(n=%d,designedford=%.2f)", n_patients, hypothesized_d)
 } else {
-title <- sprintf(
-"e-RTCUnderAlternative(n=%d,trued=%.2f,%.0f%%powerdesign)",
-n_patients, true_d, target_power * 100
-trajectories <- vector("list", n_trials)
-for (i in 1:n_trials) {
-trial <- simulate_trial_continuous(n_patients,
-mu_ctrl = mu_ctrl,
-mu_trt = mu_trt,
-sd = sd_true)
-wealth <- compute_eRTC(trial\$treatment,
-trial\$outcome,
-p = 0.5,
-burn_in = burn_in,
-ramp = ramp,
-c_max = c_max)
-trajectories[[i]] <- data.frame(
-patient = 1:n_patients,
-wealth = wealth,
-trial = i
+sprintf("e-RTCUnderAlternative(n=%d,trued=%.2f)", n_patients, true_d)
+trajectories <- lapply(1:n_trials, function(i) {
+trial <- simulate_trial_continuous(n_patients, mu_ctrl = mu_ctrl,
+mu_trt = mu_trt, sd = sd_true)
+wealth <- compute_eRTC(trial\$treatment, trial\$outcome, p = 0.5,
+burn_in = burn_in, ramp = ramp, c_max = c_max)
+data.frame(patient = 1:n_patients, wealth = wealth, trial = i)
 df <- bind_rows(trajectories)
-p <- ggplot(df, aes(x = patient, y = wealth, group = trial)) +
+ggplot(df, aes(x = patient, y = wealth, group = trial)) +
 geom_line(alpha = 0.4, color = "steelblue") +
@@
 scale_y_log10() +
-labs(
-title = title,
+labs(title = title,
 subtitle = sprintf("mu_ctrl=%.2f,mu_trt=%.2f,trued=%.2f",
 mu_ctrl, mu_trt, true_d),
-x = "Patient",
-y = "e-value(logscale)"
-annotate("text",
-x = n_patients * 0.95,
-y = (1 / alpha) * 1.5,
-label = sprintf("1/alpha=%.0f", 1 / alpha),
-color = "red", hjust = 1) +
+x = "Patient", y = "e-value(logscale)") +
+annotate("text", x = n_patients * 0.95, y = (1/alpha) * 1.5,
+label = sprintf("1/alpha=%.0f", 1/alpha), color = "red", hjust = 1) +
 theme_minimal() +
-theme(
-plot.title = element_text(face = "bold"),
-panel.grid.minor = element_blank()
-return(p)
-# Run all scenarios (grid over d and power)
-# ds : vector of Cohen’s d (e.g. c(0.2, 0.4, 0.6))
-# target_powers: vector of target powers (e.g. c(0.80, 0.90))
-run_all_eRTC <- function(n_sims = 5000,
-mu_ctrl = 0,
-ds = c(0.2, 0.4, 0.6),
-target_powers = c(0.80, 0.90),
-alpha = 0.05,
-burn_in = 20,
-ramp = 50,
-c_max = 0.6,
-seed = 42) {
-set.seed(seed)
-results <- list()
-scenarios <- expand.grid(
-hypothesized_d = ds,
-target_power = target_powers,
-stringsAsFactors = FALSE
-all_results <- data.frame()
-for (i in 1:nrow(scenarios)) {
-hyp_d <- scenarios\$hypothesized_d[i]
-tgt_pow <- scenarios\$target_power[i]
-[…]
-[…]
-hyp_d, tgt_pow * 100))
-[…]
-# Type I error
-[…]
-t1 <- simulate_eRTC(
-n_sims = n_sims,
-mu_ctrl = mu_ctrl,
-mu_trt = NULL, # Null true
-hypothesized_d = hyp_d,
-target_power = tgt_pow,
-alpha = alpha,
-burn_in = burn_in,
-ramp = ramp,
-c_max = c_max
-# Power
-[…]
-pow <- simulate_eRTC(
-n_sims = n_sims,
-mu_ctrl = mu_ctrl,
-mu_trt = mu_ctrl + hyp_d, # sd = 1, so delta = d
-hypothesized_d = hyp_d,
-target_power = tgt_pow,
-alpha = alpha,
-burn_in = burn_in,
-ramp = ramp,
-c_max = c_max
-all_results <- rbind(all_results, data.frame(
-hypothesized_d = hyp_d,
-target_power = tgt_pow,
-n_patients = t1\$n_patients,
-type1_error = t1\$rejection_rate,
-type1_se = t1\$se,
-eRTC_power = pow\$rejection_rate,
-power_se = pow\$se,
-median_crossing = pow\$median_crossing
-results\$summary <- all_results
-[…]
-[…]
-[…]
-print(all_results %>%
-mutate(
-hypothesized_d = sprintf("%.2f", hypothesized_d),
-target_power = sprintf("%.0f%%", target_power * 100),
-type1_error = sprintf("%.3f", type1_error),
-eRTC_power = sprintf("%.1f%%", eRTC_power * 100)
-select(hypothesized_d, target_power, n_patients,
-type1_error, eRTC_power, median_crossing))
-return(results)
-# Example run (interactive)
-if (interactive()) {
-# Grid of d and power
-results_eRTC <- run_all_eRTC(
-n_sims = 2000, # lower for quick tests; bump to 5000+ for final
-mu_ctrl = 0,
-ds = c(0.2, 0.4, 0.6),
-target_powers = c(0.80, 0.90),
-alpha = 0.05,
-burn_in = 20,
-ramp = 50,
-c_max = 0.6
-# Trajectory plots - Alternative examples
-p_alt_04_80 <- plot_trajectories_eRTC(
-mu_ctrl = 0,
-mu_trt = 0.4, # true d = 0.4
-hypothesized_d = 0.4,
-target_power = 0.80,
-alpha = 0.05
-ggsave("traj_eRTC_alt_d0.4_80pow.pdf", p_alt_04_80, width = 8, height = 5)
-# Trajectory plots - Null
-p_null_04_80 <- plot_trajectories_eRTC(
-mu_ctrl = 0,
-mu_trt = NULL, # null
-hypothesized_d = 0.4,
-target_power = 0.80,
-alpha = 0.05
-ggsave("traj_eRTC_null_d0.4_80pow.pdf", p_null_04_80, width = 8, height = 5)
-# e-Survival: Sequential Log-Rank Test (Unified)
-library(tidyverse)
-library(survival) # Used for standard verification if needed
-# Core e-Survival function
-# time: vector of event times
-# status: vector of status (1=event, 0=censored)
-# treatment: vector of assignments (1=trt, 0=ctrl)
-# burn_in: events to observe before betting
-# ramp: number of events to ramp up betting intensity
-# lambda: maximum betting scalar (0 < lambda < 1)
+theme(plot.title = element_text(face = "bold"), panel.grid.minor = element_blank())
+# --- e-Survival: Time-to-Event Outcomes ---
+library(survival)
 compute_eSurvival <- function(time, status, treatment,
-burn_in = 20,
-ramp = 50,
-lambda_max = 0.2) {
+burn_in = 20, ramp = 50, lambda_max = 0.2) {
 n <- length(time)
-# 1. Create a dataframe and sort by time (events happen sequentially)
-df <- data.frame(time, status, treatment) %>%
-arrange(time)
-# Vectors for processing
+df <- data.frame(time, status, treatment) %>% arrange(time)
 T_sorted <- df\$time
@@
 wealth[1] <- 1
-# Running Log-Rank Score (Observed - Expected)
 cumulative_Z <- 0
-# Track active Risk Set manually for speed/simplicity
-# Initially, everyone is at risk (assuming time 0 start)
 risk_trt <- sum(treatment == 1)
 risk_ctrl <- sum(treatment == 0)
-# We need to process events one by one.
-# Note: This simple loop assumes no tied times for clarity.
-# (Ties are rare in continuous simulation).
 for (i in 1:n) {
-# --- A. Betting Step (Before observing outcome i) ---
-# Determine bet direction based on history
 if (i > burn_in) {
-# Ramping factor
 c_i <- max(0, min(1, (i - burn_in) / ramp))
-# Direction:
-# If cumulative_Z is negative, Treatment has FEWER events than expected (Benefit).
-# We bet that this trend continues.
-# If Observed (1) - Expected (p) is likely negative, we want a NEGATIVE bet
-# to make the wealth multiplier (1 + bet * U) > 1.
-# Simple strategy: proportional to the sign of the current trend
 bet_direction <- sign(cumulative_Z)
-if (bet_direction == 0) bet_direction <- 0
-# The bet:
 b_i <- c_i * lambda_max * bet_direction
@@
 b_i <- 0
-# --- B. Observation Step ---
 is_event <- S_sorted[i] == 1
 tr_is_event <- (A_sorted[i] == 1)
-# Probability that *if* an event happens, it is in Treatment arm
-# p_null = (Risk Treatment) / (Total Risk)
 total_risk <- risk_trt + risk_ctrl
-if (total_risk > 0) {
-p_null <- risk_trt / total_risk
-} else {
-p_null <- 0.5
-# --- C. Update Wealth (Only on events, not censorings) ---
+p_null <- if (total_risk > 0) risk_trt / total_risk else 0.5
 if (is_event) {
-# U_i: The "score" contribution (Observed - Expected)
-# If Treatment event: 1 - p
-# If Control event: 0 - p
 obs <- ifelse(tr_is_event, 1, 0)
 U_i <- obs - p_null
-# Martingale update
-# Constraint: ensure (1 + b_i * U_i) is non-negative.
-# Since |U_i| < 1 and lambda_max < 1, this is generally safe.
 multiplier <- 1 + b_i * U_i
-# Update history
 cumulative_Z <- cumulative_Z + U_i
-if (i > 1) {
-wealth[i] <- wealth[i-1] * multiplier
+wealth[i] <- if (i > 1) wealth[i-1] * multiplier else multiplier
 } else {
-wealth[i] <- multiplier
-} else {
-# Censored event: No bet made, wealth carries forward
 if (i > 1) wealth[i] <- wealth[i-1]
-# --- D. Update Risk Set for next step ---
-# The person at index i is removed from risk set (event or censored)
 if (tr_is_event) {
@@
 risk_ctrl <- max(0, risk_ctrl - 1)
-# Map wealth back to original unsorted indices is tricky in sequential plots,
-# so we return the Sequential Wealth (wealth over event time).
-return(data.frame(
-event_num = 1:n,
-time = T_sorted,
-wealth = wealth
-# Simulate a single survival trial (Weibull)
-# HR < 1 implies Treatment is better (longer survival)
-simulate_trial_survival <- function(n,
-HR = 1,
-shape = 1.2,
-scale = 10,
-cens_prop = 0.0) {
-# 1. Treatments
+data.frame(event_num = 1:n, time = T_sorted, wealth = wealth)
+simulate_trial_survival <- function(n, HR = 1, shape = 1.2, scale = 10, cens_prop = 0.0) {
 treatment <- rbinom(n, 1, 0.5)
-# 2. True Times (Weibull AFT)
-# HR in Cox model corresponds roughly to scale change in Weibull
-# Scale_trt = Scale_ctrl / (HR^(1/shape))
 scale_trt <- scale / (HR^(1/shape))
@@
 true_time[treatment == 1] <- scale_trt * (-log(U[treatment == 1]))^(1/shape)
-# 3. Censoring
-# Simple independent uniform censoring to achieve approx proportion
 if (cens_prop > 0) {
-C <- runif(n, 0, 2 * scale) # broad censoring distribution
-# Adjust range to hit target prop roughly if needed, keeping simple here
+C <- runif(n, 0, 2 * scale)
 time <- pmin(true_time, C)
@@
 status <- rep(1, n)
-return(list(time = time, status = status, treatment = treatment))
-# Unified simulation function
-simulate_eSurvival <- function(n_sims = 1000,
-n_patients = NULL, # Change to NULL to trigger auto-calc
-HR_true = 1,
-target_HR = 0.7,
-target_power = 0.80, # Added for flexibility
-alpha = 0.05,
-burn_in = 30,
-ramp = 50,
-lambda_max = 0.25) {
-# Determine simulation type
-if (HR_true == 1) {
-sim_type <- "TypeIError(Null)"
-} else {
-sim_type <- "Power(Alternative)"
-# Calculate Sample Size if not provided
+list(time = time, status = status, treatment = treatment)
+simulate_eSurvival <- function(n_sims = 1000, n_patients = NULL, HR_true = 1,
+target_HR = 0.7, target_power = 0.80, alpha = 0.05,
+burn_in = 30, ramp = 50, lambda_max = 0.25) {
 if (is.null(n_patients)) {
-# Schoenfeld formula for events
 z_alpha <- qnorm(1 - alpha/2)
 z_beta <- qnorm(target_power)
-# Total events needed
-req_events <- ceiling( 4 * ((z_alpha + z_beta) / log(target_HR))^2 )
-# Assuming no censoring, Patients = Events
-n_patients <- req_events
-[…]
-target_power*100, target_HR, n_patients))
-[…]
-[…]
-[…]
+n_patients <- ceiling(4 * ((z_alpha + z_beta) / log(target_HR))^2)
 rejections <- 0
@@
 burn_in = burn_in, ramp = ramp, lambda_max = lambda_max)
-# Check stopping
 crossing <- which(res\$wealth >= 1/alpha)
@@
 se <- sqrt(rate * (1 - rate) / n_sims)
-median_stop <- median(first_crossing, na.rm = TRUE)
-[…]
-[…]
-if (sim_type == "Power(Alternative)") {
-[…]
-return(list(rate = rate, median_stop = median_stop, n_patients = n_patients))
-# Plot Trajectories
-plot_trajectories_surv <- function(n_trials = 20,
-n_patients = 400,
-HR_true = 0.7,
-alpha = 0.05,
-title = "e-SurvivalTrajectories") {
-trajectories <- list()
-for (i in 1:n_trials) {
+list(rate = rate, se = se, median_stop = median(first_crossing, na.rm = TRUE),
+n_patients = n_patients)
+plot_trajectories_surv <- function(n_trials = 20, n_patients = 400, HR_true = 0.7,
+alpha = 0.05, title = "e-SurvivalTrajectories") {
+trajectories <- lapply(1:n_trials, function(i) {
 trial <- simulate_trial_survival(n_patients, HR = HR_true)
@@
 res\$trial <- i
-trajectories[[i]] <- res
+res
 df <- bind_rows(trajectories)
@@
 scale_y_log10() +
-labs(
-title = title,
-subtitle = sprintf("TrueHR=%.2f", HR_true),
-x = "NumberofEvents",
-y = "e-value(logscale)"
+labs(title = title, subtitle = sprintf("TrueHR=%.2f", HR_true),
+x = "NumberofEvents", y = "e-value(logscale)") +
 annotate("text", x = n_patients, y = 1/alpha,
@@
 theme_minimal()
-# Run Scenarios
-if (interactive()) {
-# Define design parameters
-target_hr <- 0.80
-# 1. Check Type I Error (Null: HR = 1.0)
-# Passing n_patients = NULL triggers the auto-calculation based on target_HR
-simulate_eSurvival(
-n_sims = 2000,
-n_patients = NULL, # Auto-calculate N for 80% power
-HR_true = 1.0, # Reality: Null
-target_HR = target_hr,
-alpha = 0.05
-# 2. Check Power (Alternative: HR = 0.7)
-simulate_eSurvival(
-n_sims = 2000,
-n_patients = NULL, # Auto-calculate N for 80% power
-HR_true = 0.8, # Reality: Alternative works
-target_HR = target_hr,
-alpha = 0.05
-# 3. Plot Trajectories
-# We calculate N explicitly here just to pass it to the plot function
-req_events <- ceiling(4 * ((qnorm(0.975) + qnorm(0.80)) / log(target_hr))^2)
-p <- plot_trajectories_surv(
-n_trials = 30,
-n_patients = req_events,
-HR_true = 0.8,
-alpha = 0.05,
-title = sprintf("e-Survival(TargetHR=%.2f,N=%d)", target_hr, req_events)
-print(p)
-set.seed(2)
-p1<-plot_trajectories_surv(
-n_patients = ceiling(4 * ((qnorm(0.975) + qnorm(0.80)) / log(0.8))^2),
-HR_true = 0.8)
-ggsave("traj_survivaleRT_08.pdf",p1,
-width = 8, height = 5)
-p2<-plot_trajectories_surv(
-n_patients = ceiling(4 * ((qnorm(0.975) + qnorm(0.80)) / log(0.8))^2),
-HR_true = 1)
-p2
-ggsave("traj_survivaleRT_null.pdf",p2,
-width = 8, height = 5)
-# APPENDIX: Verification of Staggered vs. Simultaneous Entry
-# This script demonstrates that the "Simultaneous Entry" assumption used in the
-# main simulation yields statistically identical results to a realistic
-# "Staggered Entry" design, provided the analysis is based on "Time on Study".
-verify_staggered_equivalence <- function(n_sims = 1000,
-n_patients = 631, # Updated to match HR=0.8 scenario
-HR_true = 0.8, # Updated to match HR=0.8 scenario
-recruit_period = 12, # Recruitment over 12 months
+# --- Staggered vs Simultaneous Entry Verification ---
+verify_staggered_equivalence <- function(n_sims = 1000, n_patients = 631,
+HR_true = 0.8, recruit_period = 12,
 alpha = 0.05) {
-[…]
-# Storage for final e-values
 final_e_simultaneous <- numeric(n_sims)
@@
 for (i in 1:n_sims) {
-# 1. Generate the underlying biology (Survival Times)
-# -- Method A: Simultaneous (The Simple Way) --
+# Simultaneous
 trial_sim <- simulate_trial_survival(n_patients, HR = HR_true)
@@
 final_e_simultaneous[i] <- tail(res_sim\$wealth, 1)
-# -- Method B: Staggered (The "Real World" Way) --
-# We generate fresh biology for the staggered arm to compare distributions
+# Staggered
 trial_stag <- simulate_trial_survival(n_patients, HR = HR_true)
-# ... but we add "Recruitment Time" (Uniform over 12 months)
 entry_time <- runif(n_patients, 0, recruit_period)
-# The event happens at this Calendar Time:
-calendar_event_time <- entry_time + trial_stag\$time
-# The analyst calculates "Time on Study" (Analysis Time):
-# Analysis Time = Calendar Event Time - Entry Time
-calculated_study_time <- calendar_event_time - entry_time
-# Run e-process on the calculated study time
-# Note: compute_eSurvival handles the sorting of events internally
-res_stag <- compute_eSurvival(calculated_study_time,
-trial_stag\$status,
+calculated_study_time <- entry_time + trial_stag\$time - entry_time
+res_stag <- compute_eSurvival(calculated_study_time, trial_stag\$status,
 trial_stag\$treatment)
@@
 close(pb)
-# 2. Compare Results
 df_res <- data.frame(
 e_value = c(final_e_simultaneous, final_e_staggered),
-Method = rep(c("Simultaneous(Simplified)", "Staggered(Realistic)"), each = n_sims)
-# Summary Stats
-print(df_res %>%
-group_by(Method) %>%
-summarise(
-Mean_E = mean(e_value),
-Median_E = median(e_value),
-Power_Estimated = mean(e_value >= 1/alpha)
-# 3. Visualization
-p <- ggplot(df_res, aes(x = e_value, fill = Method)) +
+Method = rep(c("Simultaneous", "Staggered"), each = n_sims)
+ggplot(df_res, aes(x = e_value, fill = Method)) +
 geom_density(alpha = 0.5, color = NA) +
 scale_x_log10() +
-labs(
-title = "EquivalenceofSimultaneousvs.StaggeredEntry",
-subtitle = sprintf("Distributionoffinale-values(N=%d,HR=%.2f)", n_patients, HR_true),
-x = "Finale-value",
-y = "Density"
+labs(title = "EquivalenceofSimultaneousvs.StaggeredEntry",
+subtitle = sprintf("N=%d,HR=%.2f", n_patients, HR_true),
+x = "Finale-value", y = "Density") +
 theme_minimal() +
 theme(legend.position = "bottom")
-return(p)
-# Run the verification
-if (interactive()) {
-# Matching the N=631 / HR=0.8 scenario from the paper
-p_verify <- verify_staggered_equivalence(
-n_sims = 1000,
-n_patients = 631,
-HR_true = 0.8
-print(p_verify)
-# Optional: Save for supplement
-ggsave("supp_staggered_equivalence.pdf", p_verify, width = 8, height = 5)
+# --- e-RTms: Multi-State Models ---
+# States: 1=Ward, 2=ICU, 3=Home (absorbing), 4=Dead (absorbing)
+P_ctrl <- matrix(c(
+0.880, 0.070, 0.030, 0.020,
+0.070, 0.915, 0.000, 0.015,
+0.000, 0.000, 1.000, 0.000,
+0.000, 0.000, 0.000, 1.000
+), nrow = 4, byrow = TRUE)
+P_trt <- matrix(c(
+0.870, 0.050, 0.050, 0.030,
+0.090, 0.900, 0.000, 0.010,
+0.000, 0.000, 1.000, 0.000,
+0.000, 0.000, 0.000, 1.000
+), nrow = 4, byrow = TRUE)
+BURN_IN <- 30
+RAMP <- 50
+MAX_DAYS <- 28
+START_STATE <- 2
+THRESHOLD <- 20
+GOOD_TRANS <- c("2_1", "1_3")
+simulate_patient_ms <- function(P, max_days = MAX_DAYS, start_state = START_STATE) {
+state <- start_state
+transitions <- NULL
+for (day in 1:max_days) {
+if (state %in% c(3, 4)) break
+new_state <- sample(1:4, 1, prob = P[state, ])
+if (new_state != state) {
+transitions <- rbind(transitions, c(from = state, to = new_state, day = day))
+state <- new_state
+list(final_state = state, transitions = transitions)
+simulate_trial_ms <- function(n_patients, P_trt, P_ctrl) {
+all_trans <- NULL
+all_arms <- c()
+final_states <- data.frame(patient = 1:n_patients, arm = NA, final_state = NA)
+for (i in 1:n_patients) {
+arm <- sample(0:1, 1)
+P <- if (arm == 1) P_trt else P_ctrl
+pat <- simulate_patient_ms(P)
+final_states\$arm[i] <- arm
+final_states\$final_state[i] <- pat\$final_state
+if (!is.null(pat\$transitions)) {
+for (j in 1:nrow(pat\$transitions)) {
+all_trans <- rbind(all_trans, pat\$transitions[j, ])
+all_arms <- c(all_arms, arm)
+list(transitions = all_trans, arms = all_arms, final_states = final_states)
+compute_eRTms <- function(all_transitions, all_arms,
+burn_in = BURN_IN, ramp = RAMP,
+good_trans = GOOD_TRANS) {
+n <- nrow(all_transitions)
+if (is.null(n) $|$ n == 0) return(1)
+wealth <- numeric(n)
+wealth[1] <- 1
+n_good_trt <- 0; n_total_trt <- 0
+n_good_ctrl <- 0; n_total_ctrl <- 0
+for (i in 1:n) {
+trans <- all_transitions[i, ]
+arm <- all_arms[i]
+trans_key <- paste0(trans["from"], "_", trans["to"])
+is_good <- trans_key %in% good_trans
+if (i > burn_in && n_total_trt > 0 && n_total_ctrl > 0) {
+c_i <- min(1, max(0, (i - burn_in) / ramp))
+rate_trt <- n_good_trt / n_total_trt
+rate_ctrl <- n_good_ctrl / n_total_ctrl
+delta <- rate_trt - rate_ctrl
+lambda <- if (is_good) 0.5 + 0.5 * c_i * delta else 0.5 - 0.5 * c_i * delta
+} else {
+lambda <- 0.5
+lambda <- max(0.01, min(0.99, lambda))
+mult <- if (arm == 1) lambda / 0.5 else (1 - lambda) / 0.5
+wealth[i] <- if (i > 1) wealth[i - 1] * mult else mult
+if (arm == 1) {
+n_total_trt <- n_total_trt + 1
+if (is_good) n_good_trt <- n_good_trt + 1
+} else {
+n_total_ctrl <- n_total_ctrl + 1
+if (is_good) n_good_ctrl <- n_good_ctrl + 1
+return(wealth)
+run_simulation_ms <- function(n_sims, n_patients, P_trt, P_ctrl,
+threshold = THRESHOLD, verbose = TRUE) {
+rejections <- 0
+first_crossing <- numeric(n_sims)
+n_transitions <- numeric(n_sims)
+if (verbose) pb <- txtProgressBar(min = 0, max = n_sims, style = 3)
+for (s in 1:n_sims) {
+trial <- simulate_trial_ms(n_patients, P_trt, P_ctrl)
+if (is.null(trial\$transitions) $|$ nrow(trial\$transitions) < BURN_IN) {
+first_crossing[s] <- NA
+n_transitions[s] <- 0
+next
+wealth <- compute_eRTms(trial\$transitions, trial\$arms)
+n_transitions[s] <- length(wealth)
+crossing <- which(wealth >= threshold)
+if (length(crossing) > 0) {
+rejections <- rejections + 1
+first_crossing[s] <- crossing[1]
+} else {
+first_crossing[s] <- NA
+if (verbose) setTxtProgressBar(pb, s)
+if (verbose) close(pb)
+rate <- rejections / n_sims
+se <- sqrt(rate * (1 - rate) / n_sims)
+list(
+rate = rate,
+se = se,
+median_crossing = median(first_crossing, na.rm = TRUE),
+median_transitions = median(n_transitions)
+plot_trajectories_ms <- function(n_trials = 30, n_patients, P_trt, P_ctrl,
+threshold = THRESHOLD, title = "") {
+trajectories <- list()
+for (i in 1:n_trials) {
+trial <- simulate_trial_ms(n_patients, P_trt, P_ctrl)
+if (!is.null(trial\$transitions) && nrow(trial\$transitions) > BURN_IN) {
+wealth <- compute_eRTms(trial\$transitions, trial\$arms)
+trajectories[[length(trajectories) + 1]] <- data.frame(
+transition = 1:length(wealth), wealth = wealth, trial = i
+df <- bind_rows(trajectories)
+ggplot(df, aes(x = transition, y = wealth, group = trial)) +
+geom_line(alpha = 0.4, color = "steelblue") +
+geom_hline(yintercept = threshold, linetype = "dashed", color = "red") +
+geom_hline(yintercept = 1, linetype = "dotted", color = "gray50") +
+scale_y_log10() +
+labs(title = title, x = "TransitionNumber", y = "e-value(logscale)") +
+annotate("text", x = Inf, y = threshold * 1.5,
+label = sprintf("1/alpha=%.0f", threshold),
+color = "red", hjust = 1.1) +
+theme_minimal() +
+theme(plot.title = element_text(face = "bold"))

### v6

@@
 Sequential monitoring of randomized trials traditionally relies on parametric assumptions or asymptotic approximations.
-We discuss a nonparametric sequential test and its application to continuous and time-to-event endpoints that derives validity solely from the randomization mechanism.
-Using a betting framework, these tests constructs a test martingale by sequentially wagering on treatment assignments given observed outcomes.
+We discuss a family of nonparametric sequential tests—collectively called e-RT—for binary, deaths-only, continuous, time-to-event, and multi-state endpoints.
+All variants derive validity solely from the randomization mechanism.
+Using a betting framework, each test constructs a test martingale by sequentially wagering on treatment assignments given observed outcomes.
 Under the null hypothesis of no treatment effect, the expected wealth cannot grow, guaranteeing anytime-valid Type I error control regardless of stopping rule.
-We prove validity and present simulation studies demonstrating calibration and power.
+We prove validity for each variant, present simulation studies demonstrating calibration and power, and discuss the principled asymmetry in betting strategies across outcome types.
 These methods provide a conservative, assumption-free complement to model-based sequential analyses.
@@
 This may both delay the implementation of potentially useful strategies or prolong trials when safety signals arise early.
-E-values and e-processes offer an alternative framework (shafer2021; vovk2021; ramdas2025).
+E-values and e-processes offer an alternative framework (Shafer, 2021; Vovk and Wang, 2021; Ramdas and Wang, 2025).
 An e-value is a measure of evidence against a null hypothesis with a specific property: its expected value under the null is at most 1.
 This simple constraint yields anytime-valid inference: the Type I error guarantee holds at any stopping time, regardless of the stopping rule.
-pmlr-v177-duan22a introduced interactive rank testing by betting (i-bet), which tests treatment effects by wagering on treatment assignments given observed outcomes.
-The intuition is discussed by ramdas2021game.
+Duan et al.
+(2022) introduced interactive rank testing by betting (i-bet), which tests treatment effects by wagering on treatment assignments given observed outcomes.
+The intuition is discussed by Ramdas (2021).
 Under the null hypothesis, randomization ensures that assignments are independent of outcomes, so no betting strategy can systematically accumulate wealth.
-We propose e-RT (e-value Randomized Trial), a method for prospective sequential monitoring of randomized trials.
-Like i-bet (pmlr-v177-duan22a), e-RT uses betting martingales for inference, but differs in key respects: e-RT monitors sequentially as patients enroll rather than analyzing completed trial data; it requires no covariates or working models; and it uses an adaptive betting fraction derived from running outcome estimates rather than fixed wagers.
-This yields a method with minimal assumptions suitable for real-time trial monitoring.
+We propose e-RT (e-value Randomized Trial), a family of methods for prospective sequential monitoring of randomized trials.
+Like i-bet (Duan et al., 2022), e-RT uses betting martingales for inference, but differs in key respects: e-RT monitors sequentially as patients enroll rather than analyzing completed trial data; it requires no covariates or working models; and it uses adaptive betting fractions derived from running outcome estimates rather than fixed wagers.
+This yields methods with minimal assumptions suitable for real-time trial monitoring.
+We present five variants: the binary e-RT for event/no-event outcomes; e-RTd for deaths-only monitoring (requiring no survivor tracking); e-RTc for continuous endpoints; e-RTs for time-to-event data; and e-RTms for multi-state trajectory data.
+All share the same validity proof—the expected wealth multiplier is exactly 1 under the null—but differ in how they translate outcome data into wagers.
+We also discuss the principled asymmetry in betting strategy design: why sparse-update methods (survival, deaths-only) can afford aggressive fixed wagers while dense-update methods (binary, continuous) require conservative adaptive strategies.
 2 Overall construction
@@
 2.2 Wealth Process
-Following pmlr-v177-duan22a, we construct a wealth process by wagering on treatment assignments.
+Following Duan et al.
+(2022), we construct a wealth process by wagering on treatment assignments.
 After observing outcome $Y_{i}$ but before learning treatment assignment $T_{i}$, we choose $\lambda_{i}\in[0,1]$: the fraction wagered on intervention.
@@
 Corollary 1.
-By Ville’s inequality (ville1939), for any nonnegative martingale starting at 1:
+By Ville’s inequality (Ville, 1939), for any nonnegative martingale starting at 1:
 $\Pr_{H_{0}}\left(\sup_{n\geq 1}W_{n}\geq\frac{1}{\alpha}\right)\leq\alpha$ (9)
@@
 Under the alternative, approximately half of trajectories cross the threshold, typically around the midpoint of enrollment.
+4 Deaths-Only Monitoring (e-RTd)
+4.1 Motivation
+The binary e-RT requires knowing both the treatment arm and the outcome for every enrolled patient.
+In practice, this means a coordinator must ascertain whether each patient experienced the event or not—requiring follow-up, data entry, and outcome adjudication for all patients, including the majority who survive.
+In some settings, especially those embedded in electronic medical records (EMR) or when studying mortality in acutely ill patients, deaths are reliably captured but survivors may not be systematically followed.
+A death is an unmissable event; a survivor at day 28 requires active confirmation.
+This motivates a simpler variant: e-RTd monitors only the stream of deaths, ignoring survivors entirely.
+The trade-off is explicit: by discarding information about survivors, we lose statistical efficiency (approximately $2.5\times$ sample size inflation compared with frequentist methods).
+What we gain is operational simplicity—the only data required are the arm labels on deaths, in the order they occur.
+No denominators, no follow-up windows, no survivor tracking.
+4.2 Null Hypothesis
+Consider a trial with 1:1 randomization.
+Under the null hypothesis of no treatment effect on mortality, both arms have the same mortality rate.
+If both arms have equal mortality and equal enrollment, then each death is equally likely to come from either arm:
+$H_{0}:P(\text{death from treatment}\mid\text{a death occurred})=0.5$ (10)
+In plain language: if the treatment does not affect mortality, then looking at the arm label on a death certificate is like flipping a fair coin.
+Treatment deaths and control deaths should arrive in roughly equal numbers.
+This reduces trial monitoring to a sequential test of a Bernoulli coin: is the fraction of treatment deaths equal to $0.5$? We call this the “death coin.”
+4.3 Algorithm
+The algorithm maintains two counters: $d_{\text{trt}}$ and $d_{\text{ctrl}}$, both initialized to zero. For each death $i=1,2,\ldots$:
+Step 1: Estimate the coin bias from past data. Using all deaths before the current one, compute the plug-in estimator:
+$\hat{p}_{i-1}=\frac{d_{\text{trt}}}{d_{\text{trt}}+d_{\text{ctrl}}}$ (11)
+If no deaths have been observed yet, set $\hat{p}=0.5$.
+In words: $\hat{p}$ is the running fraction of deaths that came from the treatment arm.
+Under the null, this should hover around $0.5$.
+Under the alternative (if treatment helps), this should drift below $0.5$.
+Step 2: Compute the wager. The betting fraction is:
+$\lambda_{i}=0.5+c_{i}\cdot(\hat{p}_{i-1}-0.5)$ (12)
+clamped to $[0.001,0.999]$, where $c_{i}$ is the same ramp function as before:
+$c_{i}=\min\left(1,\max\left(0,\frac{i-n_{0}}{n_{r}}\right)\right)$ (13)
+with burn-in $n_{0}=30$ deaths and ramp $n_{r}=50$ deaths.
+During the burn-in, $c_{i}=0$ and $\lambda_{i}=0.5$ (a neutral bet—no information yet).
+After the ramp completes, $c_{i}=1$ and $\lambda_{i}=\hat{p}_{i-1}$: the wager fully reflects the observed death proportion.
+In plain language: we bet proportionally to what we have learned so far.
+If 40% of past deaths were from treatment ($\hat{p}=0.40$), we set $\lambda=0.40$.
+This means we place more of our wager on the “control death” side, because past data suggest control deaths are more common.
+At full ramp ($c_{i}=1$), this is the full Kelly bet—the wager that maximizes the expected logarithmic growth rate of wealth.
+This contrasts with the half-Kelly strategy used in the binary e-RT.
+Full Kelly is viable here because events are sparse: a trial with 2,000 patients and 15% mortality produces approximately 300 deaths.
+Over-betting compounds across only a few hundred multiplications, causing gradual wealth erosion rather than the catastrophic collapse seen with dense (every-patient) updates.
+Step 3: Update wealth.
+[…]
+starting from $W_{0}=1$.
+In plain language: we place $\lambda_{i}$ on “treatment death” and $1-\lambda_{i}$ on “control death.” The null hypothesis pays $2:1$ on the correct side (a fair payout for a $0.5$-probability event).
+If we guessed correctly, wealth grows; if we guessed wrong, wealth shrinks.
+Step 4: Update counters after betting.
+Increment $d_{\text{trt}}$ or $d_{\text{ctrl}}$ depending on which arm the death came from.
+This ordering—bet first, then update—ensures that the estimate $\hat{p}_{i-1}$ uses only past information, maintaining the martingale property.
+4.4 Worked Example
+Consider a trial comparing a new treatment against standard of care in the ICU.
+Suppose 80 deaths have been observed so far.
+Burn-in (30) and ramp (50) are complete, so $c_{i}=1$.
+Current counts: $d_{\text{trt}}=33$, $d_{\text{ctrl}}=47$.
+Death 81 arrives. Let us walk through the update.
+Step 1: $\hat{p}_{80}=33/(33+47)=33/80=0.4125$. Interpretation: so far, 41.25% of deaths came from the treatment arm. This is below 50%, suggesting treatment may be protective.
+Step 2: $\lambda_{81}=0.5+1.0\times(0.4125-0.5)=0.4125$.
+We place 41.25% of our bet on “treatment death” and 58.75% on “control death.” We are leaning toward control because past evidence suggests treatment deaths are less frequent.
+Step 3: The arm is revealed: it is a control death.
+$\text{Multiplier}=\frac{1-0.4125}{0.5}=\frac{0.5875}{0.5}=1.175$ (15)
+Wealth grows by 17.5%. Our bet was correct—we leaned toward control, and indeed it was a control death.
+Step 4: Update $d_{\text{ctrl}}=48$. Now $\hat{p}_{81}=33/81=0.407$.
+Had the death been from the treatment arm instead:
+$\text{Multiplier}=\frac{0.4125}{0.5}=0.825$ (16)
+Wealth would have shrunk by 17.5%. Our lean toward control would have been wrong.
+Under the null, treatment and control deaths arrive with equal probability, so wins and losses balance on average.
+Under the alternative ($\hat{p}<0.5$), control deaths are genuinely more frequent, and wealth grows systematically.
+4.5 Validity
+Theorem 2.
+[…]
+Proof.
+Condition on all past information and the current wager $\lambda_{i}$.
+Under the null, the probability that death $i$ is from the treatment arm is exactly $0.5$, independently of all past data.
+The expected wealth multiplier is:
+$\mathbb{E}\left[\frac{W_{i}}{W_{i-1}}\,\middle|\,\lambda_{i}\right]$ $=0.5\times\frac{\lambda_{i}}{0.5}+0.5\times\frac{1-\lambda_{i}}{0.5}$ (17)
+$=\lambda_{i}+(1-\lambda_{i})=1$ (18)
+This is the same identity as in the binary case.
+Since the expected multiplier is exactly 1, wealth cannot systematically grow under the null, regardless of how $\lambda_{i}$ was chosen (as long as it depends only on past deaths).
+∎
+By Ville’s inequality, $\Pr_{H_{0}}(\sup_{i\geq 1}W_{i}\geq 1/\alpha)\leq\alpha$. Rejecting when wealth crosses $1/\alpha$ controls Type I error at any stopping time.
+4.6 Bidirectionality
+The method automatically detects both benefit and harm without pre-specifying a direction:
+  • If $\hat{p}<0.5$ (fewer treatment deaths than expected): the method bets on control deaths being more common, and wealth grows when treatment is indeed protective.
+  • If $\hat{p}>0.5$ (more treatment deaths than expected): the method bets on treatment deaths being more common, and wealth grows when treatment is harmful.
+No investigator input about the expected direction is needed. The adaptive wager discovers the direction from the data.
+4.7 Signal Concentration
+A key property of e-RTd is that it can outperform the full-sample binary e-RT when baseline mortality is low. The intuition is that deaths concentrate the treatment signal.
+Consider a trial where the control mortality is 25% and treatment mortality is 20%, yielding a 5 percentage-point absolute risk reduction (ARR).
+In the full-sample binary e-RT, the signal per patient is diluted: most patients survive, and only the 20–25% who die carry information about differential mortality.
+The observed risk difference across all patients is 5 percentage points.
+In e-RTd, only deaths are observed. The death-coin probability is:
+$p_{\text{alt}}=\frac{p_{\text{trt}}}{p_{\text{trt}}+p_{\text{ctrl}}}=\frac{0.20}{0.20+0.25}=0.444$ (19)
+This is an 11.2-point tilt from 0.5—more than double the 5-point ARR. The signal is concentrated because deaths filter out the uninformative survivors.
+This advantage diminishes as baseline mortality increases.
+At higher event rates, the binary method sees more informative events per patient, and the death-coin tilt shrinks because both numerator and denominator grow:
+Table 2: Signal Concentration: Death-Coin Tilt vs. ARR for a 5pp Risk Reduction
+Baseline Mortality Treatment Mortality Death Coin $p_{\text{alt}}$ Tilt from 0.5 Tilt / ARR
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+The crossover with the binary e-RT occurs at approximately 25% baseline event rate (Table 4).
+Below that, the concentrated death-coin signal more than compensates for the smaller number of observations; above it, the binary method’s access to all patients provides the advantage.
+4.8 Simulation Studies
+We evaluated e-RTd operating characteristics using the same simulation framework as for binary e-RT.
+For each scenario, we calculated the frequentist sample size (two-proportion $z$-test) and applied a $2.5\times$ inflation factor to account for the reduced efficiency of deaths-only monitoring.
+We used burn-in =30=30 deaths, ramp =50=50 deaths, threshold $1/\alpha=20$, and 2,000 simulations per scenario.
+Table 3: Operating Characteristics for e-RTd (Deaths-Only)
+Baseline ARR Death Coin N (freq) N (e-RTd) Type I Power Median Crossing
+15% 5pp 0.400 1,372 3,430 2.20% 90.3% 179 deaths (42%)
+15% 10pp 0.250 282 705 0.35% 59.6% 63 deaths (89%)
+25% 5pp 0.444 2,188 5,470 2.90% 86.0% 490 deaths (40%)
+25% 10pp 0.375 500 1,250 2.15% 87.1% 123 deaths (49%)
+35% 5pp 0.462 2,754 6,885 3.05% 76.1% 1,036 deaths (46%)
+35% 10pp 0.417 658 1,645 3.15% 81.1% 232 deaths (47%)
+Type I error was well controlled across all scenarios (0.35–3.15%), consistently below the nominal $\alpha=0.05$.
+Power ranged from 59.6% to 90.3%, with the strongest performance at low baseline mortality rates where the signal concentration effect is most pronounced: at 15% baseline mortality with a 5pp ARR, the death-coin tilt of 10 points from 0.5 yielded 90.3% power.
+At higher baseline rates (35%), the death-coin tilt shrinks and power decreases, consistent with the signal concentration analysis above.
+When the threshold was crossed, it occurred at approximately 40–50% of total expected deaths for most scenarios.
+The exception is 15% baseline with 10pp ARR, where the very strong death-coin tilt (0.250, a 25-point deviation from 0.5) drives rapid crossing at the cost of requiring very few deaths overall (median 63), so the 89% figure reflects that the total expected death count (71) is small and crossing happens close to the end.
+4.9 Trajectory Examples
+Figure 3 shows example wealth trajectories for e-RTd.
+Under the null (left), the death coin is fair and wealth fluctuates around 1.
+Under the alternative (right), treatment deaths arrive less frequently than control deaths, the adaptive wager learns this imbalance, and wealth grows.
+Figure 3: Trajectories of the e-RTd process (25% baseline mortality, 5pp ARR, 500 deaths).
+Left: under the null hypothesis (death coin =0.50=0.50), wealth fluctuates randomly.
+Right: under the alternative hypothesis, wealth grows as the adaptive wager learns the death-coin imbalance.
+Dashed red line: rejection threshold ($1/\alpha=20$).
+4.10 Head-to-Head Comparison with Binary e-RT
+To quantify the signal concentration crossover, we ran both e-RTd and binary e-RT on the same simulated trials across a range of baseline mortality rates with a fixed 5pp ARR.
+For each baseline rate, we computed the frequentist sample size (two-proportion $z$-test, 80% power), enrolled that many patients, and analyzed the data with both methods: binary e-RT processed all patients; e-RTd processed only the death stream.
+We used 2,000 simulations per scenario.
+Table 4: Head-to-Head Comparison: e-RTd vs. Binary e-RT (5pp ARR, same trial data)
+Baseline Death Coin NN Deaths Binary Power e-RTd Power $\Delta$ Winner
+10% 0.333 870 66 12.2% 8.4% 3.7-3.7pp Binary
+15% 0.400 1,372 172 26.7% 44.6% +17.9+17.9pp e-RTd
+20% 0.429 1,812 318 32.4% 44.0% +11.6+11.6pp e-RTd
+[…]
+30% 0.455 2,502 689 42.3% 38.4% 3.9-3.9pp Binary
+35% 0.462 2,754 896 45.3% 36.1% 9.2-9.2pp Binary
+40% 0.467 2,942 1,104 48.8% 33.1% 15.6-15.6pp Binary
+The crossover occurs at exactly 25% baseline mortality, confirming the analytical prediction from Table 2.
+At 15–20% baseline, e-RTd outperforms binary e-RT by 12–18 percentage points despite seeing fewer observations, because the death-coin tilt (10–7 points from 0.5) more than compensates for the smaller sample.
+Above 30%, binary e-RT’s access to all patients provides an increasingly large advantage.
+The 10% exception.
+At 10% baseline, the frequentist $N=870$ produces only $\sim$66 expected deaths—fewer than the burn-in (30) plus ramp (50) = 80 deaths required for e-RTd to reach full betting strength.
+The e-RTd process literally never completes its learning phase, so it cannot leverage the strong 16.7-point death-coin tilt.
+This highlights a practical constraint: e-RTd requires sufficient deaths (>80>80) to be effective.
+The 10pp ARR case.
+With a larger effect (10pp ARR), the frequentist $N$ shrinks dramatically (282 at 15% baseline), producing even fewer deaths.
+Binary e-RT dominates across all tested baseline rates because the larger per-patient effect can be detected with fewer observations, while e-RTd cannot accumulate enough deaths during its burn-in/ramp phase.
+Practical guidance.
+Use e-RTd when (i) baseline mortality is 15–25%, (ii) the expected ARR is modest ($\leq$5pp), and (iii) outcome ascertainment beyond death is impractical.
+Above 25% baseline or with larger expected effects, binary e-RT is more powerful.
 5 Continuous Outcomes
@@
 Finally, we cap the maximum betting strength at $c_{\max}\in(0,0.5]$ to avoid pathological bets.
+Additionally, we need a direction estimate: which arm has better outcomes? Using all previous data, we compute the running Cohen’s $d$:
+$\hat{d}_{i-1}=\frac{\bar{Y}_{\text{trt}}-\bar{Y}_{\text{ctrl}}}{s_{\text{pooled}}}$ (24)
+clamped to [1,1][-1,1], where $\bar{Y}_{\text{trt}}$ and $\bar{Y}_{\text{ctrl}}$ are the arm-specific means and $s_{\text{pooled}}$ is the pooled standard deviation.
+This is the “doubly adaptive” structure: $g_{i}$ captures how informative the current observation is, while $\hat{d}_{i-1}$ captures the direction and magnitude of the treatment effect estimated from all past data.
 The betting fraction $\lambda_{i}$ is then
-$\lambda_{i}=0.5+c_{i}c_{\max}g_{i}.$ (14)
+$\lambda_{i}=0.5+c_{i}\cdot c_{\max}\cdot g_{i}\cdot\hat{d}_{i-1}.$ (25)
 By construction, $\lambda_{i}\in(0,1)$ and is predictable: it depends only on past outcomes and the new $Y_{i}$, not on $T_{i}$.
+In words: if the current observation $Y_{i}$ is extreme in the direction that past data associate with the treatment arm, $\lambda_{i}$ deviates substantially from 0.5, placing a confident bet on treatment.
+If $Y_{i}$ is typical, or if past data show no treatment effect ($\hat{d}\approx 0$), $\lambda_{i}$ stays near 0.5 and we barely bet.
 Intuitively:
-  • If $Y_{i}$ is close to the historical median, $g_{i}\approx 0$ and $\lambda_{i}\approx 0.5$: we essentially do not bet.
-  • If $Y_{i}$ is unusually low (for example, very high ventilator-free days in a trial where we expect intervention to improve outcomes), then $g_{i}<0$ or $g_{i}>0$ depending on how we code the outcome and the expected direction of benefit.
-The sign of $g_{i}$ determines which arm we lean towards.
+  • If $Y_{i}$ is close to the historical median, $g_{i}\approx 0$ and $\lambda_{i}\approx 0.5$: we essentially do not bet, regardless of how strong the estimated treatment effect is.
+[…]
+  • If $Y_{i}$ is extreme but $\hat{d}\approx 0$ (no estimated effect yet), $\lambda_{i}$ stays near 0.5—an unusual observation is not informative if we do not yet know which direction to bet.
   • Early in the trial, $c_{i}$ is small, so even unusual observations lead to mild bets. As data accumulates, $c_{i}$ approaches 1 and the bets become more confident.
@@
 Imagine a trial where the outcome is ventilator-free days, and higher is better.
-Suppose that after 100 patients, the median and MAD of $Y$ are roughly stable.
+Suppose that after 100 patients, the median and MAD of $Y$ are roughly stable, and the running Cohen’s $d$ estimate is approximately 1.0 (a large effect).
 Patient 101 has an unusually high number of ventilator-free days compared with this distribution.
 The standardized residual $r_{101}$ is positive and large, so $g_{101}\approx 0.8$ and, after burn-in, $c_{101}\approx 1$.
-With $c_{\max}=0.6$, we might get $\lambda_{101}\approx 0.5+0.6\times 0.8\approx 0.98$: we strongly bet that this patient was in the intervention arm.
+With $c_{\max}=0.6$ and $\hat{d}\approx 1.0$, we get $\lambda_{101}\approx 0.5+0.6\times 0.8\times 1.0\approx 0.98$: we strongly bet that this patient was in the intervention arm.
 If they indeed were, wealth increases by roughly a factor of $0.98/0.5\approx 2$ for this one patient.
@@
 We then simulated trials with that sample size, assigning patients 1:1 to intervention or control, with outcomes drawn from normal distributions of equal variance and means differing by $d$ under the alternative.
-The test was run sequentially with a burn-in period and ramp (for example, burn-in = 20 patients, ramp = 50 patients, $c_{\max}=0.6$).
+The test was run sequentially with a burn-in period and ramp (burn-in = 50 patients, ramp = 100 patients, $c_{\max}=0.6$).
 Under the null (no mean difference between arms), Type I error was close to or below the nominal level.
@@
 One can extend the randomization e-process to survival data, constructing a sequential Log-Rank test that updates wealth at every observed event.
-pmlr-v146-grunwald21a developed a safe logrank test using e-values under a proportional hazards model with a prior on the hazard ratio.
+Grünwald et al.
+(2021) developed a safe logrank test using e-values under a proportional hazards model with a prior on the hazard ratio.
 We attempted to construct a nonparametric approach where validity derives solely from randomization, not from a correctly specified hazard model.
@@
 Once evidence favors one direction, the bet size is fixed at $\lambda_{\max}$ regardless of how strong the accumulated evidence is.
-This mirrors the Kelly criterion in betting theory: the optimal wager size depends on the expected edge, and $\lambda_{\max}=0.25$ is calibrated for moderate effects (HR $\approx 0.80$(kelly1956).
+This mirrors the Kelly criterion in betting theory: the optimal wager size depends on the expected edge, and $\lambda_{\max}=0.25$ is calibrated for moderate effects (HR $\approx 0.80$(Kelly, 1956).
 The connection between the wager, $\lambda$, and Kelly’s ideas on fraction of betting needs to be further explored.
@@
 For trials where the clinical question is “does treatment help patients recover faster?”, e-RTms provides a natural framework.
+8 Betting Strategy Design: The Wage Asymmetry
+The five e-RT variants use different betting strategies.
+Binary and continuous methods use adaptive wagers that learn from data; survival uses a fixed wager $\lambda=0.25$; deaths-only uses adaptive full Kelly.
+This section explains why these differences are principled, not accidental.
+8.1 The Core Asymmetry: Update Density and Over-Betting
+The wealth process is a product of multipliers: $W_{n}=\prod_{i=1}^{n}(1+b_{i}u_{i})$.
+When the bet magnitude $|b_{i}|$ is too large relative to the true effect size (“over-betting”), some multipliers fall well below 1.
+The critical question is: how fast does wealth recover?
+The answer depends on how frequently the product updates:
+Survival (e-RTs): Wealth updates only at events (deaths or failures).
+Between events, wealth is unchanged.
+A trial with $N=500$ events produces 500 multiplications.
+Over-betting makes individual multipliers more volatile, but the limited number of updates means wealth bleeds slowly.
+In simulation, setting $\lambda=0.25$ when the true hazard ratio is 0.90 (where the optimal $\lambda$ is approximately 0.15) yields a median final e-value of approximately 0.04—low, but not zero.
+The wealth process is damaged but not destroyed.
+[…]
+Continuous (e-RTc): The problem is even worse.
+Beyond the dense updates (every patient), the $g$-score introduces a second source of variability in each wager.
+Fixed-wage betting is catastrophic at all tested levels for effects of $d\leq 0.20$.
+In plain language: imagine a gambler who bets too aggressively.
+If they play once a week (survival), a bad streak hurts but they can recover.
+If they play every hour (binary), the same over-bet compounds into ruin.
+If they play every hour and each bet has extra noise on top of the sizing error (continuous), the ruin is faster still.
+8.2 Survival: Why Fixed $\lambda=0.25$ Works
+For survival data, the bet has fixed magnitude but adaptive direction:
+$b_{i}=c_{i}\times\lambda_{\max}\times\text{sign}(Z_{i-1})$ (38)
+where $\lambda_{\max}=0.25$ is fixed and $Z_{i-1}$ is the cumulative log-rank score.
+Simulation studies across the clinically relevant range of hazard ratios show that $\lambda=0.25$ is not universally optimal but is robustly adequate:
+Table 10: Survival: Fixed $\lambda=0.25$ vs. Adaptive Half-Kelly
+True HR $|\log(\text{HR})|$ Fixed $\lambda=0.25$ Adaptive $\frac{1}{2}$-Kelly Difference
+0.70 0.357 43.1% 26.6% ++16.5pp
+0.75 0.288 58.9% 35.0% ++23.9pp
+0.80 0.223 62.6% 35.9% ++26.7pp
+0.85 0.163 55.2% 37.8% ++17.4pp
+0.90 0.105 40.5% 41.6% -1.1pp
+Fixed $\lambda=0.25$ outperforms adaptive half-Kelly by 17–27 percentage points across HR =0.70=0.700.850.85.
+At HR =0.90=0.90 (a small effect), the strategies are approximately equivalent.
+The adaptive approach converges correctly to the true $\log(\text{HR})$ but produces bets that are systematically too conservative: for HR =0.80=0.80, half-Kelly yields $\lambda\approx 0.11$, which is well below $0.25$.
+The fixed strategy succeeds because events are sparse.
+A trial with approximately 500 events tolerates moderate over-betting without catastrophic wealth destruction.
+Pre-specifying $\lambda$ is analogous to pre-specifying the alternative hypothesis for sample size calculation—every frequentist trial makes a similar design choice.
+8.3 Binary: Why Adaptive Half-Kelly Is Necessary
+For binary data, the wager scales with the running risk difference:
+$\lambda_{i}=0.5\pm 0.5\times c_{i}\times\hat{\delta}_{i-1}$ (39)
+The factor of $0.5$ before $c_{i}\hat{\delta}$ implements half-Kelly betting. Why not use the full estimate, or a fixed wager?
+Simulation demonstrates the catastrophic cost of over-betting in binary trials. For a true ARR of 5% (Kelly-optimal $\lambda\approx 0.05$):
+Table 11: Binary: Cost of Over-Betting (True ARR = 5%, N =2,942=2{,}942)
+Fixed $\lambda$ Relative to Kelly Power Median Final E-value
+0.05 $1\times$ 56.3% 1.15
+0.10 $2\times$ 21.9% $\approx 0$
+0.15 $3\times$ 16.5% $\approx 0$
+0.20 $4\times$ 10.0% $\approx 0$
+At $2\times$ Kelly, the median e-value is already zero—wealth is irreversibly destroyed. Every patient updates the product, and thousands of slightly-wrong multiplications compound into oblivion.
+Moreover, no single fixed $\lambda$ works universally.
+The optimal fixed $\lambda$ ranges from 0.05 for small effects to 0.15 for moderate effects.
+In practice, trialists routinely power for optimistic effects (e.g., ARR $10\%$) when the truth may be 3–5%.
+If $\lambda$ inherits this optimism, wealth is destroyed.
+The adaptive half-Kelly wager is not a power sacrifice—it is insurance against the trialist’s optimism.
+The 7-percentage-point power loss relative to a perfectly calibrated fixed wager ($1\times$ Kelly: 56.3% vs.
+adaptive: 49.4%) is the cost of being agnostic to the design alternative.
+That cost is worth paying because the alternative (inheriting an inflated design effect) risks irreversible wealth destruction.
+8.4 Continuous: The Strongest Case for Adaptive Betting
+The continuous e-RTc has a “doubly adaptive” structure:
+$\lambda_{i}=0.5+c_{i}\times c_{\max}\times g_{i}\times\hat{d}_{\text{past}}$ (40)
+Both $g_{i}$ (observation-level informativeness) and $\hat{d}_{\text{past}}$ (running Cohen’s $d$) are adaptive.
+Removing the magnitude of $\hat{d}$ and using only its sign (making $\lambda$ “fixed” in the direction sense) is catastrophic:
+Table 12: Continuous: Adaptive vs. Fixed Direction Magnitude
+Cohen’s $d$ NN Adaptive Fixed $c=0.3$ Fixed $c=0.6$
+0.50 128 16.8% 15.5% 45.3%
+0.30 352 50.1% 50.1% 25.2%
+0.20 788 53.9% 29.3% 13.6%
+At Cohen’s $d=0.20$ (a typical clinical trial effect size), fixed $c=0.6$ retains only 13.6% power (vs.
+53.9% adaptive), and fixed $c=0.3$ drops to 29.3%.
+The median final e-value is $\approx 0$ in both cases—wealth is irreversibly destroyed.
+The $g$-score adds a second noise source that, combined with dense updates, makes even moderate fixed betting lethal.
+8.5 The Design Principle
+Table 13: Summary: Betting Strategy by e-RT Variant
+Property e-RTs e-RTd e-RT e-RTc
+Update frequency Events only Events only Every patient Every patient
+Over-bet cost Low (slow bleed) Moderate Catastrophic Catastrophic
+Universal $\lambda$ possible? Yes Marginal No No
+Kelly fraction Fixed ($\lambda=0.25$) Full Half Doubly adaptive
+The hierarchy of over-betting risk is: continuous >> binary $\gg$ survival.
+This maps directly to how frequently the wealth product is updated.
+More frequent updates mean more opportunities for over-betting to compound, requiring increasingly conservative or adaptive strategies.
+The e-RTd falls between survival and binary.
+Like survival, it updates only at events (deaths), making the product length moderate.
+This permits full Kelly rather than half-Kelly.
+However, unlike survival where $\lambda_{\max}=0.25$ is fixed, e-RTd uses the plug-in $\hat{p}$ directly because the death-coin estimator converges quickly and the natural scale of the problem ($p\in[0,1]$) bounds the wager without external calibration.
 9 Discussion
-The e-RT is a nonparametric sequential test for randomized trials based on the betting framework for e-values (i-bet pmlr-v177-duan22a).
-The method requires only that treatment assignment is randomized — no distributional assumptions about outcomes are needed.
-This makes it a robust complement to model-based analyses.
-It can be extended to continuous endpoints and time-to-event analyses.
+The e-RT family comprises five nonparametric sequential tests for randomized trials based on the betting framework for e-values (i-bet Duan et al.
+(2022)).
+All variants require only that treatment assignment is randomized—no distributional assumptions about outcomes are needed.
+This makes them robust complements to model-based analyses.
+The variants cover binary outcomes, deaths-only monitoring, continuous endpoints, time-to-event analyses, and multi-state trajectory data.
 9.1 Operating characteristics
-Simulations demonstrate that the method properly controlled for Type I error (2–3% vs nominal 5%) while providing approximately 50% power for early stopping in trials designed with 80% power, and 63–66% power in trials designed with 90% power.
-When early stopping occurs, it happens at roughly half the planned sample size.
-Both the continuous (e-RTC) and the survival (e-survival) implementations had similar characteristics: Power was traded for continuous Type 1 error control.
-All simulations showed that Type 1 error was exceptionally well controlled but power was nominally low.
+Across all five variants, simulations demonstrate proper Type I error control (consistently 2–4% vs.
+nominal 5%), confirming the theoretical guarantee from the martingale property.
+Power varies by variant and scenario: for binary outcomes, approximately 50% power for early stopping in trials designed with 80% frequentist power, and 63–66% in trials designed with 90% power.
+Continuous and survival variants show similar patterns.
+The multi-state variant achieved 89% power at N=1, [^0] N=1{,}000 for the specific transition matrix tested.
+The deaths-only variant (e-RTd) requires approximately $2.5\times$ sample size inflation compared with the frequentist benchmark, the cost of discarding survivor information.
+When early stopping occurs across all variants, it typically happens at approximately 45–56% of the planned sample size or event count.
 These results should be interpreted carefully.
-From a clinical trial perspective, it is uncertain whether those methods can replace traditional frequentist or Bayesian paradigms, but they may provide a continuous monitoring option that requires no alpha spending and no prespecified interim analysis schedule.
-If the e-process crosses its threshold, one could consider stopping the trial, for example.
+From a clinical trial perspective, it is uncertain whether these methods can replace traditional frequentist or Bayesian paradigms, but they may provide a continuous monitoring option that requires no alpha spending and no prespecified interim analysis schedule.
+If the e-process crosses its threshold, one could consider stopping the trial.
 If it does not cross, the trial may proceed to its planned conclusion and primary analysis.
+9.2 Traditional statistics at crossing
+When an e-RT crosses its threshold ($\geq 1/\alpha$), one may ask what a traditional frequentist analysis would report at that exact moment.
+In practice, the standard test will agree: the p-value will be small and the confidence interval will exclude the null.
+The e-RT is not detecting phantoms.
+However, there is a critical distinction.
+The traditional analysis at crossing is formally invalid because the analyst observed the data at multiple time points and stopped at a favorable one.
+The frequentist p-value is anti-conservative (too small) and the standard confidence interval has less than 95% coverage at this data-driven stopping time.
+The e-value, by contrast, is valid regardless of the stopping rule—this is the entire point of the anytime-valid guarantee.
+The e-value itself carries the inferential weight; the traditional statistics at crossing serve only as a sanity check confirming that the e-RT is detecting a real signal, not an artifact.
 9.3 What is the null hypothesis being tested?
@@
 Rejecting the null means we have found a profitable strategy—outcomes predict assignments better than chance.
-The user is also referenced to the pioneer lessons by ramdas2021game.
+The user is also referenced to the pioneer lessons by Ramdas (2021).
 This framing clarifies both the method’s strength and its limitation.
@@
 9.4 Relationship to existing work
-The betting framework for hypothesis testing was developed by shafer2021.
-E-values and e-processes have been extensively studied (vovk2021; ramdas2022; ramdas2025).
-pmlr-v177-duan22a introduced interactive rank testing by betting (i-bet), which applies the betting framework directly to randomized experiments: an analyst sequentially bets on treatment assignments based on observed outcomes, with wealth forming a test martingale under the null.
+The betting framework for hypothesis testing was developed by Shafer (2021).
+E-values and e-processes have been extensively studied (Vovk and Wang, 2021; Ramdas et al., 2022; Ramdas and Wang, 2025).
+Duan et al.
+(2022) introduced interactive rank testing by betting (i-bet), which applies the betting framework directly to randomized experiments: an analyst sequentially bets on treatment assignments based on observed outcomes, with wealth forming a test martingale under the null.
 The binary approach implements this framework with a specific adaptive betting strategy tied to outcome values.
-Betting approaches have been established for estimating means of bounded random variables (waudbysmith2023estimating).
+Betting approaches have been established for estimating means of bounded random variables (Waudby-Smith and Ramdas, 2023).
 The continuous extension adapts these principles to the two-sample randomization setting using a standardization strategy.
-koning2025 develops e-values for group invariance, including permutation tests, using batch-based likelihood ratio statistics normalized by permutation expectations.
-pmlr-v146-grunwald21a developed the ‘Safe Log-rank Test’ based on evaluating likelihood ratios with specific priors on the hazard ratio to ensure growth rate optimality.
+Koning (2025) develops e-values for group invariance, including permutation tests, using batch-based likelihood ratio statistics normalized by permutation expectations.
+Grünwald et al.
+(2021) developed the ‘Safe Log-rank Test’ based on evaluating likelihood ratios with specific priors on the hazard ratio to ensure growth rate optimality.
 In contrast, the survival approach constructs a linear test martingale directly from the score function using an adaptive betting strategy.
@@
 This is an experimental method under development.
-Simulations are not exhaustive and were done for binary, continuous endpoints, and time to event.
-It is uncertain how this method would behave in more complex models, including competing risk models.
-The methods proposed test only whether there are differences between arms.
-They do not directly estimate treatment effects or provide confidence intervals.
+Simulations are not exhaustive, and the operating characteristics reported are specific to the scenarios tested.
+It is uncertain how these methods would behave in more complex models, including competing risk models.
+The methods test only whether there are differences between arms; they do not provide point estimates or confidence intervals.
 The adaptive learning of $\hat{\delta}$ requires a burn-in period during which little evidence accumulates.
-Third, for trials where parametric assumptions are plausible, model-based sequential methods will generally have better power.
-Fourth, our simulations used specific betting strategies; other choices may yield different operating characteristics.
-Hypothetically, it could be possible to increase betting aggressiveness over time as evidence accumulates; this may result in better power.
-Finally, it is unclear how this method will behave in situations where heterogeneity in treatment effects exist or there are temporal instabilities in effect size.
-The method could be extended to bet according to relative effect size approaches, such as the odds ratio.
-This is also under development.
+For trials where parametric assumptions are plausible, model-based sequential methods will generally have better power.
+Our simulations used specific betting strategies; other choices may yield different operating characteristics.
+The betting strategy design section provides guidance on strategy selection, but optimal calibration for specific clinical scenarios remains an open question.
+Finally, it is unclear how these methods will behave when heterogeneity in treatment effects exists or there are temporal instabilities in effect size.
+Extensions to relative effect size approaches (e.g., odds ratio) are under development.
 9.6 Future Directions
@@
 Gemini 3.0 Pro aided with coding for survival approach.
-Claude 4.5 Opus provided final spellchecking and organization of this draft, which remains suboptimal despite the efforts of several LLMs.
+Claude Opus 4.6 aided with the deaths-only extension and the wage asymmetry analysis in V6.
 10.3 Acknowledgments
@@
   4. 4.
-    Fourth Version (Dec 17, 2025): Correction on signal direction on e-RTC. Text adjustments
+    Fourth Version (Dec 17, 2025): Correction on signal direction on e-RTC. Text adjustments.
   5. 5.
     Fifth Version (Dec 31, 2025): Added multi-state extension (e-RTms).
+  6. 6.
+    Sixth Version (Feb 15, 2026): Added deaths-only monitoring (e-RTd).
+Added betting strategy design section explaining wage asymmetry across variants.
+Reconciled e-RTc parameters (burn-in 50, ramp 100) and direction formula (full Cohen’s $d$ clamped to [1,1][-1,1]).
+Added traditional statistics at crossing discussion.
+Updated abstract and introduction to cover all five variants.
+Reordered sections.
 Appendix A R Code
@@
 theme(plot.title = element_text(face = "bold"), panel.grid.minor = element_blank())
+# --- e-RTd: Deaths-Only Monitoring ---
+compute_eRTd <- function(arms, burn_in = 30, ramp = 50, threshold = 20) {
+n <- length(arms)
+if (is.factor(arms)) arms <- as.numeric(arms) - 1
+arms <- as.numeric(arms)
+wealth <- numeric(n)
+wealth[1] <- 1
+d_trt <- 0; d_ctrl <- 0
+crossed <- FALSE; crossed_at <- NA
+for (i in 1:n) {
+total <- d_trt + d_ctrl
+p_obs <- if (total > 0) d_trt / total else 0.5
+if (i > burn_in && total > 0) {
+c_i <- min(1, max(0, (i - burn_in) / ramp))
+lambda <- 0.5 + c_i * (p_obs - 0.5)
+lambda <- max(0.001, min(0.999, lambda))
+} else {
+lambda <- 0.5
+mult <- if (arms[i] == 1) lambda / 0.5 else (1 - lambda) / 0.5
+wealth[i] <- if (i == 1) mult else wealth[i - 1] * mult
+if (arms[i] == 1) d_trt <- d_trt + 1 else d_ctrl <- d_ctrl + 1
+if (!crossed && wealth[i] >= threshold) {
+crossed <- TRUE; crossed_at <- i
+total <- d_trt + d_ctrl
+final_p <- if (total > 0) d_trt / total else 0.5
+final_rr <- if (final_p > 0.001 && final_p < 0.999) {
+final_p / (1 - final_p)
+} else if (final_p <= 0.001) { 0 } else { Inf }
+list(wealth = wealth, crossed = crossed, crossed_at = crossed_at,
+final_p = final_p, final_rr = final_rr, d_trt = d_trt, d_ctrl = d_ctrl)
+simulate_deaths <- function(n_deaths, p_trt_given_death) {
+rbinom(n_deaths, 1, p_trt_given_death)
+death_coin <- function(p_ctrl, p_trt) {
+p_trt / (p_trt + p_ctrl)
+expected_deaths <- function(n_patients, p_ctrl, p_trt) {
+ceiling(n_patients / 2 * (p_ctrl + p_trt))
+simulate_eRTd <- function(n_sims = 1000, p_ctrl, p_trt,
+n_patients = NULL, target_power = 0.80,
+alpha = 0.05, burn_in = 30, ramp = 50) {
+threshold <- 1 / alpha
+if (is.null(n_patients)) {
+$s$ <- power.prop.test(p1 = p_ctrl, p2 = p_trt,
+power = target_power, sig.level = alpha)
+n_freq <- 2 * ceiling(ss\$n)
+n_patients <- ceiling(n_freq * 2.5)
+n_deaths_null <- expected_deaths(n_patients, p_ctrl, p_ctrl)
+n_deaths_alt <- expected_deaths(n_patients, p_ctrl, p_trt)
+p_coin_alt <- death_coin(p_ctrl, p_trt)
+null_rej <- 0; alt_rej <- 0
+alt_crossing <- numeric(n_sims)
+for (sim in 1:n_sims) {
+# Null
+deaths <- simulate_deaths(n_deaths_null, 0.5)
+res <- compute_eRTd(deaths, burn_in, ramp, threshold)
+if (res\$crossed) null_rej <- null_rej + 1
+# Alternative
+deaths <- simulate_deaths(n_deaths_alt, p_coin_alt)
+res <- compute_eRTd(deaths, burn_in, ramp, threshold)
+if (res\$crossed) alt_rej <- alt_rej + 1
+alt_crossing[sim] <- ifelse(res\$crossed, res\$crossed_at, NA)
+list(type1 = null_rej / n_sims, power = alt_rej / n_sims,
+median_crossing = median(alt_crossing, na.rm = TRUE),
+n_patients = n_patients, n_deaths_alt = n_deaths_alt)
+plot_trajectories_eRTd <- function(n_trials = 30, n_deaths,
+p_trt_given_death,
+burn_in = 30, ramp = 50, alpha = 0.05,
+title = NULL) {
+threshold <- 1 / alpha
+if (is.null(title)) {
+title <- if (p_trt_given_death == 0.5) {
+sprintf("e-RTdUnderNull(p=0.50,%ddeaths)", n_deaths)
+} else {
+sprintf("e-RTdUnderAlternative(p=%.3f,%ddeaths)",
+p_trt_given_death, n_deaths)
+trajectories <- lapply(1:n_trials, function(i) {
+deaths <- simulate_deaths(n_deaths, p_trt_given_death)
+res <- compute_eRTd(deaths, burn_in, ramp, threshold)
+data.frame(death = 1:n_deaths, wealth = res\$wealth, trial = i)
+df <- bind_rows(trajectories)
+ggplot(df, aes(x = death, y = wealth, group = trial)) +
+geom_line(alpha = 0.4, color = "steelblue") +
+geom_hline(yintercept = threshold, linetype = "dashed", color = "red") +
+geom_hline(yintercept = 1, linetype = "dotted", color = "gray50") +
+scale_y_log10() +
+labs(title = title, x = "DeathNumber", y = "e-value(logscale)") +
+annotate("text", x = n_deaths * 0.95, y = threshold * 1.5,
+label = sprintf("1/alpha=%.0f", threshold),
+color = "red", hjust = 1) +
+theme_minimal() +
+theme(plot.title = element_text(face = "bold"),
+panel.grid.minor = element_blank())
+signal_concentration_table <- function(
+baseline_rates = c(0.10, 0.15, 0.20, 0.25, 0.30, 0.35, 0.40),
+arr = 0.05) {
+data.frame(
+baseline = baseline_rates,
+trt_rate = baseline_rates - arr,
+death_coin = sapply(baseline_rates, function(p) death_coin(p, p - arr)),
+tilt_from_half = sapply(baseline_rates,
+function(p) abs(death_coin(p, p - arr) - 0.5)),
+arr = arr
 # --- e-RTC: Continuous Outcomes ---
-compute_eRTC <- function(treatment, outcome, p = 0.5, burn_in = 20,
-ramp = 50, c_max = 0.6) {
+compute_eRTC <- function(treatment, outcome, p = 0.5, burn_in = 50,
+ramp = 100, c_max = 0.6) {
 n <- length(treatment)
@@
 mean_ctrl <- mean(out_prev[trt_prev == 0])
-direction <- if (is.nan(mean_trt) $|$ is.nan(mean_ctrl)) 0 else sign(mean_trt - mean_ctrl)
+# Cohen’s d (clamped to [-1, 1]) -- not just sign
+if (is.nan(mean_trt) $|$ is.nan(mean_ctrl)) {
+d_hat <- 0
+} else {
+sd_trt <- sd(out_prev[trt_prev == 1])
+sd_ctrl <- sd(out_prev[trt_prev == 0])
+if (is.na(sd_trt) $|$ sd_trt == 0) sd_trt <- 1
+if (is.na(sd_ctrl) $|$ sd_ctrl == 0) sd_ctrl <- 1
+s_pooled <- sqrt((sd_trt^2 + sd_ctrl^2) / 2)
+d_hat <- max(-1, min(1, (mean_trt - mean_ctrl) / s_pooled))
 ramp_frac <- max(0, min(1, (i - burn_in) / ramp))
 c_i <- ramp_frac * c_max
-lambda <- 0.5 + c_i * g_i * direction
+lambda <- 0.5 + c_i * g_i * d_hat
 lambda <- max(0.001, min(0.999, lambda))
@@
 hypothesized_d, target_power = 0.80, alpha = 0.05,
-burn_in = 20, ramp = 50, c_max = 0.6) {
+burn_in = 50, ramp = 100, c_max = 0.6) {
 if (missing(hypothesized_d)) stop("Mustspecifyhypothesized_d")
@@
 hypothesized_d, target_power = 0.80,
-alpha = 0.05, burn_in = 20, ramp = 50,
+alpha = 0.05, burn_in = 50, ramp = 100,
 c_max = 0.6, title = NULL) {
@@
 compute_eSurvival <- function(time, status, treatment,
-burn_in = 20, ramp = 50, lambda_max = 0.2) {
+burn_in = 30, ramp = 50, lambda_max = 0.25) {
 n <- length(time)